### v1

# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

Abstract
In conventional treatments of electron transport, momentum relaxation within a perfect, defect-free crystal is commonly assumed to require phonon creation or annihilation. Here we treat the crystal as finite and isolated, retaining the lattice center-of-mass (recoil) degree of freedom and enforcing conservation of total mechanical momentum alongside discrete crystal pseudomomentum. Starting from the density-density form of the electron-lattice interaction, we find that an electron in the interior of a perfect crystal admits strong, and in some regimes dominant, elastic momentum-relaxing scattering channels, in which momentum is conserved by recoil of the lattice background without phonon excitation. In addition, we identify mixed quasielastic and superelastic channels in which phonon occupations change but do not account fully for the electron’s momentum transfer. These results provide a microscopic basis for momentum relaxation that does not rely on local energy dissipation. They naturally reconcile a wide range of experimental observations—including weak localization, quantum oscillations, ultrasonic attenuation, and the separation of momentum and energy relaxation times—with predominantly elastic scattering in clean crystals. The framework clarifies how diffusive transport, including Planckian-scale diffusion, can emerge from elastic dynamics in a time-dependent lattice background.
preprint: APS/123-QED
I Introduction
Electron-lattice interactions are among the most important processes in condensed matter physics. They govern momentum relaxation, energy dissipation, and the loss of quantum coherence that underlies electrical resistance. In conventional transport theory, electron momentum relaxation in a perfect, defect-free crystal is commonly identified with inelastic scattering processes in which phonons are created or annihilated. Within this view, deflection of an electron by thermal lattice motion is inseparable from energy transfer to internal vibrational degrees of freedom.
This perspective is codified in the Bloch-Grüneisen and Boltzmann approaches to resistivity, which treat electronic transport as an incoherent sequence of scattering events and assume no electronic phase coherence beyond that retained in lowest-order scattering matrix elements. It is further reinforced by the Fröhlich interaction [10], which couples electrons explicitly to phonon creation and annihilation operators and therefore enforces inelasticity at first order in perturbation theory. Within this canonical framework, thermal deflection of electrons in a perfect crystal is taken to require phonon excitation, while elastic scattering is associated primarily with defects or impurities [39, 4].
Here we present evidence that this identification is overly restrictive. We show that thermally induced electron scattering in a pure crystal is often largely elastic when viewed on an event-by-event basis. Elastic and quasi-elastic momentum-relaxing channels exist even in a perfectly periodic lattice, without the necessity of phonon creation or annihilation. Previously, elastic scattering was almost exclusively associated with explicit disorder; we demonstrate that this association is not required by symmetry, conservation laws, or kinematics.
Early discussions of electronic transport already contained hints of this distinction. Peierls emphasized that electrical resistance requires degradation of electronic (crystal) momentum rather than dissipation of electronic energy, and noted explicitly that the lattice, as a massive collective object, can absorb momentum without requiring significant energy transfer [27]. Pippard similarly argued that defects and impurities are important but not indispensable, and that momentum relaxation could, in principle, proceed through elastic or quasi-elastic mechanisms even in very pure crystals [29]. These insights acknowledged the logical separation between momentum relaxation and energy dissipation, although this separation was not consistently maintained in subsequent formulations of transport theory.
In contrast, the canonical Fröhlich Hamiltonian and Bloch-Grüneisen framework [7] implicitly exclude rigid translation of the crystal as a whole from the dynamical description. The lattice center-of-mass (recoil) degree of freedom is effectively clamped, and momentum transfer is assigned exclusively to internal phonon excitations or to Umklapp processes. Once these assumptions are imposed, momentum loss is forced to coincide with energy dissipation, and elastic deflection by a perfect lattice is excluded by construction.
We revisit this assumption by treating the crystal as finite, isolated, and freely floating, and by explicitly retaining its total mechanical momentum. Enforcing conservation of total momentum more completely cannot be formally incorrect; at worst it may prove quantitatively unimportant. As we show below, however, restoring the lattice center-of-mass degree of freedom qualitatively alters the available scattering channels for an electron deep inside a pure conducting crystal. In particular, it reveals an elastic recoil channel that is absent by design in the conventional formulation.
The physical content of this channel is elementary. If a particle embedded in a composite object receives a sudden impulse, the momentum of the whole must change immediately, even if the internal degrees of freedom remain largely unexcited. A familiar example is provided by Mössbauer scattering, where momentum is absorbed predominantly by recoil of the entire lattice while phonon excitation is strongly suppressed. As we shall show, an analogous mechanism operates for electrons interacting with a thermally fluctuating but otherwise perfect crystal.
The ambiguity of elastic scattering
Elastic scattering is most familiar in the context of impurity or defect scattering. In that setting, momentum conservation is usually justified in one of two ways. First, the presence of an impurity breaks discrete translational symmetry, so crystal pseudomomentum need not be conserved. Second, within a particle picture, one argues that if an electron enters with one momentum and exits with another, the atomic lattice must recoil so as to conserve total mechanical momentum. Both statements are essentially correct. However, neither addresses the internal bookkeeping problem that arises for an electron scattering deep inside a perfectly periodic crystal: how does a change in the electron’s pseudomomentum communicate with the mechanical momentum of the lattice as a whole?
This unresolved point lies at the heart of the transport problem in a translationally invariant system. As we show below, it is resolved once three elements are made explicit: (i) the lattice is treated as a fully dynamical object rather than a fixed reference, (ii) pseudomomentum is assigned not only to the electron but also to the crystal background in the absence of the electron density (see the next section), and (iii) the electron-lattice interaction is formulated in density-density form [24, 10], which makes the exchange of pseudomomentum between electron and lattice explicit even in a perfectly periodic crystal.
In any untethered collection of atoms, an impulse imparted to even a single constituent instantaneously changes the total momentum of the assembly. While the subsequent redistribution of that momentum among internal degrees of freedom proceeds causally and at finite velocity, the change in the collective momentum itself is immediate. This total momentum is not an abstract bookkeeping device but a genuine dynamical variable that appears explicitly in the many-body wavefunction.
In the present formulation, this momentum conservation is enforced transparently. As shown in Eq. 12, a momentum transfer $\mathbf{q}$ to the foreground electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background. This statement holds independently of whether phonons are created or destroyed in the process.
Finally, we clarify our use of the term elastic. In a finite-temperature many-body system, strictly elastic processes do not exist in an absolute sense. Even the dominant “elastic” line in Mössbauer scattering exhibits slight thermal broadening, reflecting the thermal motion of the emitting nucleus and its finite velocity autocorrelation time. The term elastic is nevertheless standard in this context, and we adopt it here with the same understood margin. When we use the term quasielastic, we instead refer to processes involving energy exchange that is appreciably larger than the thermal broadening scale, but still too small to account for the full momentum change of the electron.
The landscape seen by an electron
At temperatures of order $T\sim 100\,$K, thermally excited acoustic distortions in a typical metal generate substantial deformation-potential landscapes on nanometer length scales. Using standard values for the deformation potential ($D\sim 5$$15\,\mathrm{eV}$), sound velocity ($v_{s}\sim 3$$5\times 10^{3}\,\mathrm{m/s}$), and the characteristic thermal phonon wavevector $q_{T}\sim k_{B}T/\hbar v_{s}$, one finds order-of-magnitude estimates for the resulting deformation-potential field gradients of
$|\nabla V_{\rm def}|\sim 10^{4}\text{–}10^{6}\ \mathrm{V/cm}\qquad(T\sim 100\,\mathrm{K}).$ (1)
Over a representative lateral scale of $L\sim 10\,$nm, this corresponds to peak-to-peak variations of the screened scalar potential of order
$\Delta V_{\rm pp}\sim|\nabla V_{\rm def}|\,L\sim 10\text{–}300\ \mathrm{mV}.$ (2)
[…]
II Foreground, Background, and Pseudomomentum
The conceptual distinction between momentum, crystal momentum, and electrical current has been recognized since the early development of band theory, but its implications for microscopic transport have remained subtle. In a particularly prescient analysis, Kohn emphasized that crystal momentum is a conserved quantity associated with translational symmetry, while the electrical current need not be proportional to it, and that momentum may be exchanged with the lattice without the necessity of phonon excitation or energy dissipation [17]. This observation already points to the possibility that momentum conservation can be satisfied through coherent motion of the lattice background rather than through inelastic scattering processes. What has remained largely implicit, however, is how this distinction should be implemented at the level of a local scattering event occurring deep in the interior of a crystal. In particular, the possibility that total crystal may be explicitly partitioned between a tagged electron and the remaining lattice, both with pseudomomenta appearing as dynamical variables in the many-body wavefunction, has not been systematically developed.
The present work makes this partition explicit and operational, thereby extending Kohn’s insight to a microscopic framework in which elastic deflection, coherent lattice recoil, and diffusive transport can coexist naturally in a perfect crystal. In what follows, we formalize a symmetric foreground-background structure, derive the elastic scattering amplitude within first-order perturbation theory, and compute the elastic fraction explicitly. We show that this fraction is close to unity for copper at room temperature. We also clarify how the present formulation extends the Fröhlich framework without contradicting its successful predictions, and discuss the implications for the interpretation of transport phenomena in crystalline solids.
We distinguish explicitly between foreground and background degrees of freedom. The foreground is the object of interest, here a single electron or a collection of them. The background consists of all remaining degrees of freedom, including the lattice atoms and any other electrons not treated explicitly. This distinction is essential for identifying the relevant symmetries and conserved quantities.
In a crystalline environment, the electronic Hamiltonian is invariant under discrete translations of the electron by lattice vectors, with the background held fixed. The associated conserved quantity is crystal momentum (pseudomomentum); for an electron in a periodic potential this is the familiar Bloch momentum $\hbar\mathbf{k}$.
Crucially, the same discrete translational symmetry applies to the background (see Fig. 1). If the electron (foreground A) coordinate is held fixed and the entire background B is translated by a lattice vector, the Hamiltonian is again unchanged. A consistent application of translational symmetry therefore assigns pseudomomentum not only to the electron but also to the background, even though only their sum corresponds to a mechanical momentum.
One might object that the problem can always be formulated in the rest frame of the background, with only the electron moving. However, this choice of frame obscures the conservation of total momentum of the combined system and suppresses the dynamical role of background recoil, which is essential for understanding local scattering events in the interior of a perfect crystal.
Figure 1: Sliding Blocks. Floating in space, two blocks attract one another and slide frictionlessly along their common long axis, $x$. At the top, both blocks carry conserved mechanical momenta, and clearly $P_{\rm tot}=p_{A}+p_{B}$. At the bottom, the same relation holds, but $p_{A}$ and $p_{B}$ are now pseudomomenta; their sum nevertheless equals the conserved mechanical momentum. The two blocks serve as surrogates for a foreground electron (A) in a perfect background crystal (B), illustrating that both the foreground and the background acquire pseudomomentum in each other’s presence.
Let $\hat{U}(\mathbf{h})$ translate all coordinates, foreground and background, by an arbitrary vector $\mathbf{h}\in\mathbb{R}^{3}$. For an isolated electron+background system,
[…]
so the corresponding Noether generator is the conserved total mechanical momentum, denoted by $\hat{\mathbf{Q}}$.
In the foreground/background formulation $\hat{\mathbf{Q}}$ is obtained as the sum of two subsystem pseudomomenta. Because $\hat{U}(\mathbf{h})$ factorizes into a translation of the foreground coordinate and an equal translation of the background configuration,
$\hat{U}(\mathbf{h})=\hat{U}_{\mathrm{f}}(\mathbf{h})\,\hat{U}_{\mathrm{b}}(\mathbf{h}),$ (4)
we have
$\hat{\mathbf{Q}}=\hat{\mathbf{P}}+\hat{\mathbf{p}}$ (5)
Here $\hat{\mathbf{p}}$ is the pseudomomentum carried by the foreground electron (the generator of its translation), and $\hat{\mathbf{P}}$ is the pseudomomentum carried by the background. (the generator of translating the background as a whole). Their sum $\hat{\mathbf{Q}}=\hat{\mathbf{P}}+\hat{\mathbf{p}}$ is the conserved mechanical momentum of the isolated system. For convenience, we can set it to 0; it never changes and we can now ignore it.
Earlier we remarked that a pseudomomentum cannot be combined with a mechanical momentum. We have now seen, however, that two pseudomomenta can combine to form a mechanical momentum. These rules are analogous to those governing the collision of two objects in free space, except that here the interaction is periodic in their relative coordinate, apart from defects and lattice vibrations.
The total momentum $\hat{\mathbf{Q}}$ is conserved, but the foreground and background momenta, $\hat{\mathbf{p}}$ and $\hat{\mathbf{P}}$, undergo equal and opposite changes, in exact analogy with collisions in free space. This exchange does not require the creation or annihilation of internal lattice excitations and occurs immediately at the collision event (see Fig. 2).
When a crystal containing an internal electron is treated as a closed many-body system, the total background momentum $\mathbf{P}=\sum_{i}\mathbf{p}_{i}$ is defined at each instant as the generator of translations of the background degrees of freedom. Since the foreground is held fixed in this construction, $\mathbf{P}$ is a pseudomomentum. An impulse imparted to a single atom in the background by a foreground object changes the total background momentum immediately by the same amount, resulting in a shift of the center-of-mass velocity $\Delta\mathbf{V}_{\mathrm{CM}}=\Delta\mathbf{p}/M$, independent of any subsequent internal redistribution of momentum.
The finite-speed propagation of stress, phonons, or other internal excitations governs how momentum is shared among individual degrees of freedom, but it does not affect the instantaneous value of the conserved quantity itself. This kinematic distinction follows directly from translational invariance and does not imply nonlocal dynamics or superluminal signaling; rather, it reflects the separation between global conservation laws and the causal evolution of local observables.
Figure 2: The total momentum (black) remains constant as the foreground (electron, blue) and background (lattice, red) momenta jump abruptly at the collision time $t^{*}$. There is no delay in the total lattice momentum change, even though only a single atom initially carries the momentum. While this is elementary, we emphasize it because there is a common tendency to associate momentum flow directly with energy flow in the medium. Indeed there is momentum flow among particles, but the total is fixed immediately. The total appears in the wavefunctions.
Finite-window symmetry
The interaction is periodic only over a finite distance $L$, as required for any physically realizable system. This situation is directly analogous to the classic particle in a box. Momentum is not strictly conserved in such a system; nevertheless, traveling wave packets carrying well-defined momentum can be constructed as linear combinations of standing-wave eigenstates, which individually carry no momentum. Over finite times these wave packets propagate as if momentum were conserved, until a collision with a boundary reverses their motion.
This type of approximate symmetry, valid only over finite spatial and temporal windows, is sometimes implicit in discussions of wave-packet dynamics in finite systems. For clarity, we refer to it here as finite-window symmetry. If one insists on working with finite crystals, this concept must be adopted: translational symmetry holds only within a spatial and temporal window set by the system size.
In a large crystal this construction is an excellent approximation. Localized corpuscles can be formed that propagate nearly as Bloch waves with well-defined pseudomomentum and experience long travel times before encountering a boundary. We emphasize that this corpuscle construction is a calculational convenience that permits the use of the notation and results associated with the formally unphysical Born-von Kármán boundary conditions, without assuming their literal validity.
III Lattice degrees of freedom and zero modes
We consider a finite crystal consisting of atoms with coordinates $\hat{\mathbf{R}}_{j}$ and masses $m_{j}$, with total lattice mass $M=\sum_{j}m_{j}$, minus the foreground electron. We define the center-of-mass coordinate of the lattice background as
$\hat{\mathbf{R}}\equiv\frac{1}{M}\sum_{j}m_{j}\,\hat{\mathbf{R}}_{j}.$ (6)
It is convenient to decompose the atomic coordinates as
$\hat{\mathbf{R}}_{j}$ $=\hat{\mathbf{R}}+\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j},$ (7)
$\sum_{j}m_{j}\bm{\xi}^{(0)}_{j}$ $=\mathbf{0},\qquad\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$
where $\hat{\mathbf{R}}$ is the exact lattice center-of-mass coordinate, $\bm{\xi}^{(0)}_{j}$ denote fixed reference positions measured relative to the center of mass, and $\hat{\mathbf{u}}_{j}$ represent internal, dynamical lattice displacements. The mass-weighted constraints ensure that all internal degrees of freedom carry zero net momentum, so that $\hat{\mathbf{R}}$ describes the unique translational zero mode.
The classical and quantum theory of lattice vibrations explicitly separates rigid motion from internal vibrations. Born and Huang showed that the equations of motion of a finite crystal admit three zero-frequency normal modes corresponding to rigid translation of the crystal as a whole, while the internal normal modes satisfy the mass-weighted constraint $\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$ so that internal vibrations carry no net mechanical momentum [8]. All mechanical momentum of the crystal resides in the translational zero mode.
The total center-of-mass degree of freedom of the combined foreground-background system is always conserved and need not be carried explicitly. For an interior electron, it is crucial to track the pseudomomentum of the foreground electron and separately the background,
IV Electron-Lattice interaction
IV.1 Foreground and background Hamiltonians
Given our aim of formulating transport without assuming pre-existing quasiparticles, care is required in defining the Hamiltonian. We write the total Hamiltonian as
$H=H_{\rm fg}+H_{\rm bg}+H_{\rm int},$ (8)
where $H_{\rm fg}$ describes the foreground electron, $H_{\rm bg}$ the background degrees of freedom, and $H_{\rm int}$ their coupling.
For a single foreground electron we take
$H_{\rm fg}=\frac{\hat{\mathbf{p}}^{2}}{2m},$ (9)
where $\hat{\mathbf{p}}$ denotes the generator of translations of the foreground coordinate $\hat{\mathbf{r}}$. The use of a bare kinetic-energy operator reflects the fact that the electron has not been pre-diagonalized into Bloch states. All effects of the periodic crystal potential, including band structure, pseudomomentum exchange, and scattering, are instead generated dynamically through the interaction Hamiltonian $H_{\rm int}$. This is the same starting point adopted by Fröhlich [10].
The remaining electrons of the metal are treated as part of the background. Together with the ions they establish the effective lattice potential, provide screening, and ensure mechanical stability. Their degrees of freedom are therefore implicit in $H_{\rm bg}$.
Density operator and exact factorization
The lattice density operator admits an exact factorization into a center-of-mass translation and an internal part:
$\hat{\rho}_{\mathbf{q}}$ $=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{\mathbf{q}},$ (10)
$\hat{\rho}^{\rm int}_{\mathbf{q}}$ $=\sum_{j}e^{-i\mathbf{q}\cdot(\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j})}.$
This identity expresses the fact that a uniform translation of the lattice multiplies all density Fourier components by the phase $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, while all information about internal lattice dynamics is contained in $\hat{\rho}^{\rm int}_{\mathbf{q}}$.
Density–density interaction and pseudomomentum conservation
With this factorization, the electron–lattice interaction may be written in density–density form as
$H_{\rm int}=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}\,\hat{\rho}_{-\mathbf{q}},$ (11)
which follows the standard Born-approximation scattering formalism introduced by Van Hove [34]. Using Eq. (10), this becomes
$H_{\rm int}$ $=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}\,e^{i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}$ (12)
$=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}.$
[…]
Crucially, this exchange of pseudomomentum applies whether or not there is a change a change in phonon occupation numbers. Elastic matrix elements of $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ are generically nonzero. The background zero mode is therefore capable of absorbing finite pseudomomentum with negligible energy cost, providing the kinematic basis for elastic but momentum-relaxing scattering. Any combination of phonon involment and lattice background recoil is allowed. It seems this should matter to measureable properties.
V Elastic internal scattering and the Debye-Waller fraction
In this section we show that, even within a strictly perturbative treatment based on Bloch electrons and phonon number states, robust elastic scattering channels exist already at first order in the electron–lattice coupling. These channels become explicit once the full lattice density operator is retained and the translational zero mode of the lattice is treated as a dynamical degree of freedom. The resulting processes are elastic with respect to internal lattice excitations, despite involving momentum exchange with the lattice as a whole.
Density-density coupling
We consider a single tagged (foreground) electron with coordinate $\hat{\mathbf{r}}$. The interaction between the foreground electron and the background lattice is taken to be of density-density form,
$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}).$ (13)
Separating the background density into its static equilibrium expectation value and fluctuations, $\rho_{\rm bg}=\langle\rho_{\rm bg}\rangle+\delta\rho_{\rm bg}$, the static component generates a periodic potential acting on the foreground electron,
$U_{\rm Bloch}(\hat{\mathbf{r}})=\sum_{\mathbf{G}}V(\mathbf{G})\,e^{i\mathbf{G}\cdot\hat{\mathbf{r}}}\,\langle\rho_{\rm bg}(-\mathbf{G})\rangle,$ (14)
where $\mathbf{G}$ are reciprocal lattice vectors. The resulting foreground Hamiltonian is therefore
[…]
while the remaining part of $H_{\rm int}$ describes the dynamical coupling to background fluctuations. In this way the familiar Bloch Hamiltonian emerges directly from the same density-density interaction, without being assumed a priori.
The lattice degrees of freedom (the background) are described by a Hamiltonian $H_{\rm bg}$. The coupling between foreground and background is retained in the form
$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}),$ (16)
where the tagged-electron density operator is
$\rho_{\rm fg}(\mathbf{q})=e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}$, (17)
and the full background density operator is
$\rho_{\rm bg}(-\mathbf{q})=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (18)
with $\hat{\mathbf{R}}_{j}=\hat{\mathbf{R}}+\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j}$ as defined previously.
Initial and final states
For a perturbative treatment we consider product states of foreground and background,
$\Psi_{i}=\Psi_{n\mathbf{k}}(\mathbf{r})\,\Psi^{\rm bg}_{i}(\{\mathbf{u}_{j}\}),\qquad\Psi_{f}=\Psi_{n^{\prime}\mathbf{k}^{\prime}}(\mathbf{r})\,\Psi^{\rm bg}_{f}(\{\mathbf{u}_{j}\}),$ (19)
where $\Psi^{\rm bg}_{i,f}$ are phonon number states or superpositions thereof. The electronic states are Bloch eigenstates of the static lattice potential.
Using the factorized form of the interaction, any nonvanishing matrix element $\langle\Psi_{f}|H_{\rm int}|\Psi_{i}\rangle$ implies the pseudomomentum transfers
$\Delta\mathbf{p}=+\hbar\mathbf{q},\qquad\Delta\mathbf{P}=-\hbar\mathbf{q},$ (20)
so that the total mechanical momentum of the combined system is conserved. The recoil pseudomomentum is carried by the lattice center-of-mass degree of freedom.
While much stronger coupling regimes can be treated nonperturbatively—for example within wave-on-wave (WoW) simulations—the purpose of the present analysis is to emphasize that even within standard perturbation theory there exist elastic scattering processes that are usually excluded by more restrictive formulations.
Elastic internal channel and the Debye-Waller factor
For a harmonic lattice, the internal displacements admit the normal-mode expansion
$\hat{\mathbf{u}}_{j}=\frac{1}{\sqrt{N}}\sum_{\mathbf{k},\lambda}\sqrt{\frac{\hbar}{2M\omega_{\mathbf{k}\lambda}}}\,\mathbf{e}_{\mathbf{k}\lambda}\left(a_{\mathbf{k}\lambda}e^{i\mathbf{k}\cdot\bm{\xi}^{(0)}_{j}}+a^{\dagger}_{\mathbf{k}\lambda}e^{-i\mathbf{k}\cdot\bm{\xi}^{(0)}_{j}}\right)$. (21)
The operator $e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}$ therefore contains all orders in the phonon operators and automatically generates elastic, quasielastic, and inelastic processes.
In particular, in a phonon number state basis the diagonal matrix elements are generically nonzero,
$\left\langle\{n\}\left|e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\right|\{n\}\right\rangle\neq 0,$ (22)
so that a scattering channel strictly elastic with respect to internal lattice excitations is present already at first order in $H_{\rm int}$. Upon thermal averaging (or in the ground state), this diagonal factor reduces to the Debye-Waller form
$\left\langle e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\right\rangle_{T}=\exp[-W(\mathbf{q},T)],$ (23)
which is the direct analog of the recoil-free (zero-phonon-line) fraction in Mössbauer spectroscopy and in neutron or x-ray scattering.
Relation to the Fröhlich truncation
In the original Fröhlich formulation, the exact factorization Eq. (10) is not employed. The lattice center-of-mass coordinate $\hat{\mathbf{R}}$ is not retained explicitly, which is equivalent to setting the recoil factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ to unity and treating the lattice as an infinitely massive reference frame. At the same time, the internal displacement exponential is truncated to linear order,
$e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\rightarrow 1-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}.$ (24)
The zeroth-order term reproduces the static periodic potential responsible for Bloch band formation, while the linear term generates single-phonon processes.
Although this procedure is internally consistent and highly successful in the weak-coupling regime, it removes the translational zero mode by construction. As a result, even if the full exponential in $\hat{\mathbf{u}}_{j}$ were retained, the truncated theory would still lack an explicit degree of freedom capable of absorbing the total recoil momentum of the lattice. The standard Fröhlich Hamiltonian therefore describes momentum exchange with internal lattice modes, but not with the lattice as a whole.
Elastic, quasielastic, and inelastic processes
Retaining the full operator $e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}$ in the background density ensures that elastic, quasielastic, and inelastic scattering processes are treated on equal footing already at first order in the density-density interaction. At this level, the interaction transfers a definite total momentum $\hbar\mathbf{q}$ from the foreground electron to the background, independent of how that momentum is subsequently absorbed. It is therefore useful to decompose the momentum transfer as $\mathbf{q}=\mathbf{q}_{\rm el}+\mathbf{q}_{\rm in}$, where $\mathbf{q}_{\rm el}$ corresponds to center-of-mass recoil of the background and $\mathbf{q}_{\rm in}$ is taken up by its internal degrees of freedom. The latter satisfies the exact kinematic relation
$\hbar\mathbf{q}_{\rm in}=-\sum_{\mathbf{k},\lambda}\hbar\mathbf{k}\,\Delta n_{\mathbf{k}\lambda},$ (25)
where $\Delta n_{\mathbf{k}\lambda}$ denotes the change in occupation of the background normal modes. In this formulation, the distinction between elastic and inelastic scattering is determined by the background state rather than by an expansion in ionic displacements.
Thermal averaging and the Debye-Waller fraction
In thermal equilibrium the internal lattice degrees of freedom are described by the density matrix
$\hat{\varrho}=\frac{e^{-\beta\hat{H}_{\mathrm{lat}}}}{Z}=\sum_{n}p_{n}\,|n\rangle\langle n|,\qquad p_{n}=\frac{e^{-\beta E_{n}}}{Z},$ (26)
where $|n\rangle$ denote exact many-body eigenstates of the lattice Hamiltonian $\hat{H}_{\mathrm{lat}}$ with energies $E_{n}$. For elastic internal scattering the relevant quantity is the thermal average of the phonon-diagonal matrix element of the internal density operator,
$\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}=\sum_{n}p_{n}\,\langle n|\rho^{\rm int}_{\mathbf{q}}|n\rangle.$ (27)
The corresponding elastic (coherent) scattering intensity is then given by
$I_{\mathrm{el}}(\mathbf{q})=\big|\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}\big|^{2}.$ (28)
By contrast, the total scattering probability summed over all final internal states is
$I(\mathbf{q})=\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}.$ (29)
The elastic internal branching fraction is therefore
$f_{\mathrm{el}}(\mathbf{q},T)=\frac{\sum_{\alpha}p_{n}\,\big|\langle n|\rho^{\rm int}_{\mathbf{q}}|n\rangle\big|^{2}}{\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}}.$ (30)
For a harmonic lattice one finds[18, 23]
$f_{\mathrm{el}}(\mathbf{q},T)\simeq e^{-2W(\mathbf{q},T)},\qquad W(\mathbf{q},T)=\frac{1}{2}\left\langle(\mathbf{q}\cdot\hat{\mathbf{u}})^{2}\right\rangle_{T},$ (31)
showing that a substantial fraction of momentum-relaxing scattering events can be strictly elastic internally even at finite temperature.
Estimate of the elastic internal fraction at representative momentum transfers
For momentum relaxation in a metal it is natural to consider momentum transfers of order the Fermi momentum, since large-angle scattering dominates the transport relaxation rate through the usual $(1-\cos\theta)$ weighting. For copper, a free-electron estimate yields
$k_{F}\simeq 1.36\ \text{\AA }^{-1},\qquad 2k_{F}\simeq 2.7\ \text{\AA }^{-1}.$ (32)
We therefore evaluate the elastic internal (phonon-diagonal) fraction $f_{\mathrm{el}}(\mathbf{q},T)$ at representative values $q\ll k_{F}$, $q\sim k_{F}$, and $q\sim 2k_{F}$.
Within the isotropic Debye-Waller estimate used above,
$2W(\mathbf{q},T)=\frac{q^{2}\langle u^{2}\rangle_{T}}{3},\qquad f_{\mathrm{el}}(\mathbf{q},T)\simeq e^{-2W(\mathbf{q},T)}.$ (33)
Using the room-temperature mean-square displacement for copper, $\langle u^{2}\rangle_{300\mathrm{K}}\simeq 1.9\times 10^{-2}\ \text{\AA }^{2}$, one finds the values summarized in Table 1.
qq-1) $2W$ $f_{\mathrm{el}}$
$0.3$ $5.7\times 10^{-4}$ $0.9994$
$k_{F}\simeq 1.36$ $1.2\times 10^{-2}$ $0.988$
$2k_{F}\simeq 2.7$ $4.6\times 10^{-2}$ $0.955$
Table 1: Estimated elastic internal fraction $f_{\mathrm{el}}$ for copper at room temperature ($T=300$ K) for representative momentum transfers. Even for large-angle scattering with $q\sim 2k_{F}$, the majority of momentum-relaxing events remain phonon-diagonal, with momentum absorbed predominantly by the lattice background rather than by internal phonon excitation.
These estimates show that internal phonon excitation constitutes only a minor branching channel for momentum relaxation in copper, even at room temperature and for momentum transfers relevant to transport. The dominant mechanism for absorbing electron momentum is thus elastic or quasi-elastic coupling to the lattice background, consistent with the discussion above.
VI Experiments suggesting large elastic contributions
Experimental evidence for predominantly elastic momentum relaxation in metals at low temperature has accumulated for decades, particularly in the mesoscopic-transport and quantum-interference literature. In this section we show that the framework developed here resolves several longstanding disparities between theory and experiment regarding electronic transport. We review a range of experiments that demonstrate a clear separation between momentum relaxation and energy relaxation, consistent with elastic or quasi-elastic scattering dominating momentum randomization.
Hierarchy of momentum and energy relaxation times
Electrical resistivity and optical conductivity probe momentum relaxation, quantum-interference phenomena such as weak localization probe phase coherence, and nonequilibrium measurements including shot noise, Johnson-noise thermometry, and hot-electron relaxation directly access energy exchange. Strikingly, these probes consistently reveal a strong hierarchy of timescales,
$\tau_{\mathrm{p}}\ll\tau_{E},$ (34)
established using independent experimental techniques and across a wide range of materials. Here $\tau_{\mathrm{p}}$ is the timescale on which electronic momentum (or pseudomomentum) is randomized by scattering, while $\tau_{E}$ is the timescale over which electrons exchange energy with the background and relax toward thermal equilibrium.
The consistent appearance of this hierarchy demonstrates that momentum relaxation and energy relaxation are distinct physical processes. While conventional language often associates momentum randomization with inelastic scattering, these experiments show that electronic momentum can be efficiently scrambled by a time-dependent environment without significant energy exchange. The lattice need not act as an energy sink in order to act as a momentum sink. This separation is precisely what is expected when electrons scatter elastically from a dynamically fluctuating lattice background.
Quantum interference and dephasing
Traditionally, the suppression of quantum-interference effects with increasing temperature has been attributed to inelastic scattering processes that transfer energy to the environment [22, 2]. However, it is important to distinguish carefully between dephasing, as operationally defined in transport experiments, and irreversible decoherence in the strict quantum-mechanical sense.
A wide range of experiments demonstrate that genuine energy relaxation remains slow over the same temperature range in which phase coherence is strongly suppressed. In particular, mesoscopic interference experiments show a rapid decrease—and in some cases saturation—of the phase coherence time $\tau_{\phi}$ at low temperatures, even as independent probes indicate that electron-phonon energy exchange rates continue to vanish [26, 16, 35]. This separation of timescales shows that momentum and phase randomization need not be accompanied by significant energy transfer.
As observed in weak-localization experiments on mesoscopic metals, the loss of quantum interference reflects the destruction of coherent backscattering between time-reversed trajectories. This coherence is far more sensitive than the transport relaxation time and can be suppressed by elastic, time-dependent scattering with negligible energy exchange and without irreversible decoherence [22, 26].
From this perspective, dephasing reflects the extreme sensitivity of quantum interference to weak, time-dependent fluctuations of the environment. Even in the absence of thermal excitations, lattice coordinates remain quantum operators with nonvanishing variance due to zero-point motion. As a result, elastic scattering amplitudes are not unity even at $T=0$, as expressed by the Debye-Waller factor. When such quantum fluctuations are slow on electronic timescales—as occurs for long-wavelength modes in finite samples—they act as an effectively time-dependent scattering potential, suppressing interference under ensemble and time averaging. These processes are elastic on electronic energy scales and do not constitute irreversible decoherence.
Hot-electron relaxation
A particularly direct separation of momentum relaxation from energy relaxation is provided by hot-electron experiments in metals [36, 30, 11, 12]. In these measurements, nonequilibrium electron distributions are injected and their subsequent momentum randomization and energy loss are probed independently. Over wide temperature ranges, the momentum-relaxation time inferred from transport or magnetotransport is found to be much shorter than the energy-relaxation time associated with thermalization. Electrons therefore undergo many momentum-deflecting collisions while remaining essentially isoenergetic. Such a pronounced separation of timescales cannot be explained by purely inelastic phonon scattering and instead points to an effectively elastic deflection channel intrinsic to the metallic state.
Johnson-noise thermometry
Johnson-noise thermometry provides a particularly stringent probe of energy relaxation in driven metallic microstructures [36, 30]. In these experiments, a dc bias drives a substantial current while the local electronic temperature is inferred directly from the Johnson-Nyquist noise. Because the noise power is proportional to the electronic energy distribution, this technique provides a quantitative and noninvasive measure of energy flow, independent of momentum transport.
A key observation is that under bias the electrical resistivity and momentum randomization rate increase strongly, while the electronic temperature remains comparatively low and spatially nonuniform. This demonstrates that electrons can undergo many momentum-deflecting scattering events without efficiently transferring energy to the lattice. Such behavior is difficult to reconcile with a picture in which inelastic phonon scattering controls momentum relaxation, but follows naturally if the dominant scattering mechanism is elastic deflection from a dynamically fluctuating lattice background.
VII Discussion
The purpose of this section is twofold. First, we examine the consistency of the present framework with well-established experimental facts, including the Wiedemann-Franz law, weak localization, and quantum oscillations. These topics are not speculative; they address potential objections that might otherwise arise when elastic scattering plays a dominant microscopic role. Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott-Ioffe-Regel crossover. These latter points are necessarily more exploratory and are presented as a coherent physical interpretation rather than as definitive conclusions.
VII.1 Implications for the Lorenz ratio
Because elastic scattering randomizes momentum just as effectively as inelastic scattering, no contradiction arises between predominantly elastic microscopic dynamics and the well-established success of conventional transport theory.
In a degenerate metal, the Wiedemann-Franz (WF) law states that the ratio of the electronic thermal conductivity $\kappa$ to the electrical conductivity $\sigma$ times temperature approaches the universal Sommerfeld value
$L\equiv\frac{\kappa}{\sigma T}\;\xrightarrow[T\to 0]{}\;L_{0}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}.$ (35)
While some fraction of inelastic processes must exist to relax the electronic energy current and establish global thermal equilibrium, it has long been understood that these processes need not control the transport coefficients themselves [39, 21, 31].
If the dominant momentum-relaxing channel is elastic and only weakly dependent on energy over the thermal window $\sim k_{B}T$ about the Fermi level, then both electrical and thermal currents are governed by the same relaxation time $\tau_{\mathrm{el}}$,
$\sigma\propto\tau_{\mathrm{el}},\qquad\kappa\propto\tau_{\mathrm{el}},$ (36)
and the Sommerfeld value $L_{0}$ is recovered. Rare inelastic events, with rate $1/\tau_{\mathrm{in}}\ll 1/\tau_{\mathrm{el}}$, serve to transfer energy from the electronic system to internal lattice degrees of freedom, but do not enter the leading expressions for $\sigma$ or $\kappa$. A transport regime dominated by elastic momentum randomization is therefore fully compatible with the WF law, even when true energy relaxation is parametrically slow.
VII.2 Weak localization
Weak localization (WL) provides a particularly clear illustration of the dominance of elastic backscattering in metals at low temperature. The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that return to the same point after multiple elastic scattering events [6, 22]. The very existence of a WL correction therefore implies that the dominant backscattering processes are coherent and elastic.
If backscattering were predominantly inelastic, interference between time-reversed paths would be washed out and no WL correction could develop. Instead, experiments show a well-defined suppression of the conductivity at low temperature together with a characteristic negative magnetoresistance when time-reversal symmetry is broken by an applied magnetic field. As temperature is increased, WL is gradually suppressed. Although this suppression is often attributed to “inelastic scattering,” experimentally it reflects the loss of phase coherence rather than rapid energy relaxation.
Independent probes such as hot-electron relaxation and Johnson-noise thermometry show that electron-phonon energy exchange remains slow over the same temperature range. The suppression of WL therefore signals the onset of a time-dependent scattering environment that randomizes phase while leaving electronic energy essentially unchanged. Within the present framework, WL fits naturally into the hierarchy of timescales
$\tau_{p}\ll\tau_{\phi}\ll\tau_{E},$
with elastic scattering producing diffusive transport and slower temporal evolution of the lattice background destroying phase coherence without necessitating phonon creation or significant energy dissipation.
VII.3 Shubnikov-de Haas oscillations
Quantum oscillation phenomena such as the Shubnikov-de Haas and de Haas-van Alphen effects provide an independent and equally stringent demonstration that electronic scattering in metals is predominantly elastic at low temperature [33, 3]. These oscillations arise from the quantization of electronic orbits in a magnetic field and reflect coherent motion around the Fermi surface over many cyclotron periods.
The oscillation amplitude is controlled by the Dingle factor [9],
$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right),$ (37)
where $\tau_{q}$ is the quantum lifetime associated with Landau-level broadening. Importantly, $\tau_{q}$ is primarily sensitive to elastic momentum-randomizing processes, while genuinely inelastic energy exchange rapidly suppresses oscillations. Any scattering event that transfers an energy comparable to $\hbar\omega_{c}$ destroys the coherence required to complete a cyclotron orbit.
The observation of quantum oscillations over broad temperature ranges and down to low magnetic fields therefore implies that electrons undergo many momentum-deflecting collisions without significant energy exchange. As in the case of weak localization, the temperature dependence of the oscillation amplitude reflects phase smearing due to a time-dependent scattering environment rather than rapid inelastic backscattering. Quantum oscillations thus provide direct experimental evidence that elastic scattering dominates electron dynamics in normal metals at low temperature.
VII.4 Sound attenuation
Ultrasonic attenuation in metals provides a complementary probe of electron-lattice dynamics from the lattice side. Experimentally, sound waves propagating through a metal are strongly attenuated even in very clean samples, and this attenuation is often interpreted as evidence for inelastic electron-phonon scattering [28, 25]. However, attenuation of a coherent acoustic wave does not, by itself, imply irreversible energy dissipation.
From a microscopic perspective, attenuation measures the loss of coherence of a propagating phonon mode. This loss can arise through elastic scattering processes that redistribute acoustic momentum among phonon modes or transfer momentum to the electronic subsystem, even when the total energy of the combined electron-lattice system is conserved. In this sense, sound attenuation is directly analogous to the suppression of quantum interference in electronic transport: coherent propagation is degraded without requiring strong energy relaxation [19, 1].
Standard theories of ultrasonic attenuation distinguish regimes based on the relative timescales of the sound frequency and electronic or phononic relaxation rates. In the collisionless (Landau-Rumer) regime, attenuation arises from scattering of sound by thermal excitations and mode conversion rather than from viscous energy dissipation [20, 25]. More generally, the acoustic wave loses forward coherence through elastic momentum exchange with a dynamically fluctuating background, while energy relaxation enters only at longer timescales.
Within the present framework, elastic momentum exchange between electrons and the lattice background provides a natural mechanism for ultrasonic attenuation. Time-dependent lattice distortions scatter sound waves, redirecting acoustic momentum and reducing the amplitude of the forward-propagating mode, while leaving electronic and lattice energies largely unchanged. Energy dissipation occurs subsequently through rare inelastic processes that establish thermal equilibrium.
This interpretation is consistent with experimental observations in normal metals, where substantial ultrasonic attenuation persists even in temperature regimes where independent measurements show that electron-phonon energy transfer is strongly suppressed [28, 25]. As in weak localization and quantum oscillation phenomena, sound attenuation therefore supports a picture in which elastic scattering dominates momentum transfer, while true energy relaxation plays a secondary role.
VII.5 Planckian diffusion and linear-in-$T$ resistivity
Much of the recent literature introduces Planckian behavior as a bound on a microscopic relaxation time $\tau_{\mathrm{Pl}}\sim\hbar/k_{B}T$, often interpreted as the fastest possible rate of inelastic scattering or energy dissipation [37, 14, 13, 32]. Within this narrative, transport coefficients are obtained by inserting $\tau_{\mathrm{Pl}}$ into kinetic or hydrodynamic expressions.
The framework developed in Ref. [38] and here suggests a different interpretation. The primary microscopic result is the emergence of real-space diffusion with a diffusion constant
$D\sim\frac{\hbar}{m^{\ast}},$ (38)
arising from elastic momentum scrambling in a time-dependent lattice background. This form of Planckian diffusion does not rely on quasiparticles, strong inelastic scattering, or rapid energy dissipation; it follows directly from quantum dynamics in a fluctuating potential and has been demonstrated explicitly in wave-on-wave simulations of the full electron-lattice problem [5, 38].
Once diffusion is established, Einstein relations connect $D$ to transport coefficients,
$\sigma=\chi D,\qquad D=\mu\frac{k_{B}T}{e},$ (39)
so that recasting $D$ in terms of an effective relaxation time yields
$\tau_{\mathrm{eff}}\sim\frac{\hbar}{k_{B}T},$ (40)
up to factors of order unity. From this perspective, the Planckian timescale is not a fundamental microscopic bound, but an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics. It constrains how rapidly charge spreads in space, not how rapidly energy must be irreversibly transferred to internal degrees of freedom.
This viewpoint also clarifies the emergence of linear-in-$T$ resistivity across and below the Debye scale. Within a conventional Bloch-Grüneisen picture, the Debye temperature sets the scale at which all phonon modes become available for inelastic scattering. In contrast, within the quantum acoustics framework there is no special anomaly associated with crossing the Debye temperature. Linear-in-$T$ resistivity arises smoothly as coherence is lost in a time-dependent elastic background, while energy relaxation remains comparatively slow. The Debye scale therefore marks neither the onset of momentum relaxation nor a transition to a new inelastic regime, but rather a coherence crossover.
VII.6 Mott-Ioffe-Regel crossover as loss of quasiparticle coherence
The Mott-Ioffe-Regel (MIR) condition $k_{F}\ell\sim 1$ is often interpreted as a boundary beyond which quasiparticles “cease to exist” and transport must become anomalous. From the present perspective, this interpretation is misleading. The MIR condition marks the point at which a semiclassical mean free path loses a literal geometric meaning, not a dynamical singularity in transport.
Within the quantum acoustics framework, transport evolves smoothly through $k_{F}\ell\sim 1$. Diffusion persists, but it is no longer usefully described in terms of well-defined trajectories between discrete scattering events. Crucially, this breakdown of a quasiparticle picture does not coincide with a loss of coherence. Were strongly inelastic processes truly dominant at the MIR scale, interference phenomena such as weak localization and quantum oscillations would be rapidly suppressed, contrary to experimental observation. Instead, experiments show that momentum relaxation can be strong while inelastic energy relaxation remains weak.
Wave-on-wave simulations sharpen this picture by explicitly demonstrating that momentum can be transferred elastically to a thermally distorted lattice background while preserving electronic energy. The MIR crossover therefore reflects a loss of quasiparticle coherence rather than the onset of strong dissipation, fully consistent with elastic-dominated momentum relaxation in a dynamical lattice environment.
VIII Summary
Summary
In the conventional clamped-lattice formulation of electron-phonon scattering, the factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ associated with rigid translation of the lattice background is removed by assumption. This elimination suppresses an otherwise allowed elastic momentum-transfer channel, forcing momentum relaxation to be identified exclusively with inelastic phonon creation or annihilation. By restoring the lattice translational zero mode, we show that this identification is not required by symmetry, conservation laws, or kinematics. In a finite crystal, elastic pseudomomentum transfer between an electron and the lattice background is generically present and can dominate momentum relaxation at low and intermediate temperatures.
Making this elastic channel explicit requires abandoning several silent but consequential restrictions embedded in the traditional formulation. First, the lattice must be treated as a finite, isolated object so that its center-of-mass degree of freedom remains dynamical. (If the crystal is attached to a laboratory or substrate, this simply enlarges the background without altering the underlying momentum bookkeeping.) Second, discrete translational symmetry must be implemented symmetrically, with foreground and background pseudomomenta treated as equally real and conserved quantities. Third, the electron-lattice interaction must be retained in its full density-density form, rather than truncated in a way that projects out elastic momentum-transfer processes by construction. When these elements are retained, elastic scattering of electrons from thermal lattice distortions is not merely allowed but unavoidable.
In this framework, pseudomomentum is exchanged between the electron and the lattice background, while total mechanical momentum is conserved through recoil of the lattice zero mode, without the necessity of phonon creation. The long-standing identification of lattice motion with inelasticity is thus revealed to be an artifact of an incomplete theoretical description.
Restoring elastic momentum-transfer processes to their proper role provides a more complete and internally consistent foundation for electronic transport in crystals. It offers a natural interpretation of experimental observations that indicate robust elastic scattering in clean materials and clarifies how momentum relaxation can be efficient even when energy relaxation is slow. More broadly, these results raise the possibility that in at least some strange-metal regimes, momentum relaxation is dominated by elastic scattering in a time-dependent background, with genuinely inelastic processes playing a secondary role in establishing energy equilibration rather than setting the transport rate.
IX Acknowledgements
E. Heller thanks B. Halperin for many discussions that provided a thorough understanding of the traditional theory. He also thanks Prof. Martina Hentschel for insightful comments that have improved the manuscript. This work was supported by the U.S. Department of Energy, Office of Science, under Grant No. DE-SC0025489.
Appendix A Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
We consider a finite one-dimensional tight-binding chain of $N$ identical atoms, isolated in space, with nearest-neighbor springs and an electron moving in a basis of localized $p$ orbitals. The chain possesses an exact translational zero mode, corresponding to its center-of-mass (CM) motion. This allows the electron to exchange momentum with the lattice as a whole without necessarily exciting internal phonon modes. As a result, at finite temperature there exists a strictly elastic backscattering channel for an electronic wavepacket.
Hamiltonian and separation of the CM mode
Let the atomic coordinates be written as
$x_{n}=R+na+u_{n},\qquad\sum_{n=1}^{N}u_{n}=0,$ (41)
where $R$ is the center-of-mass coordinate, $a$ the equilibrium lattice spacing, and $\{u_{n}\}$ the internal displacements. The lattice Hamiltonian separates exactly into CM and internal parts,
$H_{\rm lat}=\frac{P_{R}^{2}}{2M_{\rm tot}}+\sum_{n=1}^{N}\frac{p_{n}^{2}}{2M}+\frac{K}{2}\sum_{n=1}^{N-1}(u_{n+1}-u_{n})^{2},$ (42)
with total mass $M_{\rm tot}=NM$ and $\sum_{n}p_{n}=0$.
The electron is described by a nearest-neighbor tight-binding Hamiltonian,
$H_{e}=-\sum_{n=1}^{N-1}t_{n}\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right),$ (43)
where the hopping amplitudes depend linearly on the bond lengths,
$t_{n}\equiv t(x_{n+1}-x_{n})\simeq t_{0}-\alpha(u_{n+1}-u_{n}).$ (44)
The electron-lattice coupling depends only on internal coordinates; the CM coordinate $R$ drops out exactly.
Strictly elastic backscattering via CM recoil
Because the full electron-lattice system is translationally invariant, there exists an exactly conserved total momentum. The electronic band energy satisfies $\varepsilon(k)=\varepsilon(-k)$, so reversing the electronic momentum does not change the electronic energy.
Consider an incoming electronic wavepacket narrowly peaked at momentum +k+k, and an internal lattice eigenstate $|\nu\rangle$. A scattering process of the form
$|k\rangle\otimes|\nu\rangle\otimes|P_{R}\rangle\;\longrightarrow\;|-k\rangle\otimes|\nu\rangle\otimes|P_{R}+2\hbar k\rangle$ (45)
leaves the internal phonon state unchanged and transfers the momentum $2\hbar k$ entirely to the CM of the chain.
The associated change in CM kinetic energy is
$\Delta E_{\rm CM}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{4\hbar k\,P_{R}+4\hbar^{2}k^{2}}{2M_{\rm tot}}.$ (46)
Exact energy conservation requires $\Delta E_{\rm CM}=0$, which is satisfied when
$P_{R}=-\hbar k.$ (47)
For this value of the CM momentum, the backscattering process $k\rightarrow-k$ is strictly elastic, involving no excitation or absorption of internal phonon modes. Alternately, if $N$ is very large, the requirement $P_{R}=-\hbar k$ can be relaxed. For a finite, free chain of total mass $M$, electron backscattering $k\rightarrow-k$ changes the electron momentum by $\Delta p_{e}=-2\hbar k$. Momentum conservation then requires the chain to absorb a recoil $\Delta P_{\rm chain}=+2\hbar k$. If the chain initially has CM momentum $P$, energy conservation implies an associated recoil energy
$\Delta E_{\rm recoil}=\frac{(P+2\hbar k)^{2}-P^{2}}{2M}=\frac{2\hbar k\,(P+\hbar k)}{M}.$ (48)
Strictly elastic backscattering without internal excitation is therefore possible only for the special value $P=-\hbar k$; for generic initial CM momentum, an energy mismatch must be accommodated.
For a macroscopic chain, however, $M\propto N$ and the recoil energy scales as $\Delta E_{\rm recoil}\sim 1/N$, becoming parametrically small in the thermodynamic limit. In this regime the energy mismatch can be absorbed without resolving a discrete internal excitation. In particular, an electron wavepacket of spatial extent $L$ has an intrinsic energy spread
$\Delta E_{\rm pkt}\sim\hbar v_{g}\,\Delta k\sim\hbar v_{g}/L,$ (49)
which is independent of $N$ and typically exceeds $\Delta E_{\rm recoil}$ for sufficiently large chains. Equivalently, if the scattering occurs over a finite time $\tau_{\rm scatt}$, the associated time-energy uncertainty $\hbar/\tau_{\rm scatt}$ exceeds the recoil energy, rendering the process effectively elastic from the electron’s perspective.
In a phonon number-state description, the accommodation of this small energy mismatch corresponds to the creation or annihilation of one or more extremely low-frequency acoustic phonons. Importantly, however, such number-state changes do not account for the electron’s momentum reversal $k\rightarrow-k$: a single low-$q$ phonon carries momentum $\hbar q\ll 2\hbar k$. The dominant momentum transfer is instead absorbed by the collective motion of the chain (center-of-mass recoil), while the phonon sector merely provides an energetically soft channel required by exact energy conservation.
Thus, for large $N$, electron backscattering involves an exact momentum transfer of $2\hbar k$ to the background with a vanishing energy cost, and any phonon number-state changes reflect quasi-elastic energy bookkeeping rather than the mechanism of momentum reversal itself.
Role of finite temperature
At finite temperature $T$, the CM momentum is thermally distributed,
$f(P_{R})\propto\exp\!\left(-\frac{P_{R}^{2}}{2M_{\rm tot}k_{B}T}\right),$ (50)
with width $\sigma_{P}=\sqrt{M_{\rm tot}k_{B}T}$. The resonant value $P_{R}=-\hbar k$ therefore occurs with nonzero probability density for any $T>0$. Consequently, strictly elastic electronic backscattering events exist at finite temperature, mediated entirely by recoil of the lattice center of mass.
Bulk and boundary backscattering in a isolated chain
The elastic CM-recoil backscattering channel does not require the electron to encounter the ends of the chain and may occur locally in the bulk. The electron-lattice interaction in Eq. (44) depends only on nearest-neighbor bond distortions and is independent of absolute position.
For a wavepacket localized well away from the ends, with spatial extent $\ell\ll Na$, the matrix element connecting $|k\rangle$ and $|-k\rangle$ takes the form
$\langle-k,P_{R}+2\hbar k|H_{\rm int}|k,P_{R}\rangle\propto\sum_{n}e^{2ikna}\,\langle P_{R}+2\hbar k|P_{R}\rangle,$ (51)
where the sum runs over bulk sites only. The momentum transfer is absorbed entirely by the CM degree of freedom, while the internal lattice state remains unchanged.
Free boundaries of the finite chain can also reflect an electronic wavepacket once it reaches an end. Such boundary reflections are elastic and likewise involve a momentum reversal $k\rightarrow-k$ accompanied by CM recoil. Free boundaries therefore provide a spatial location for reflection but do not act as a momentum sink. Both bulk and boundary backscattering processes are governed by the same momentum conservation principle enforced by the CM zero mode.
Other scattering channels
In addition to the strictly elastic CM recoil channel, a variety of inelastic and hybrid scattering processes are allowed within a isolated chain.
Single-phonon emission and absorption
If momentum is transferred to internal phonon modes, the internal lattice state changes. A generic single-phonon process takes the form
$|k\rangle\otimes|\{n_{q}\}\rangle\longrightarrow|k^{\prime}\rangle\otimes|\{n_{q}^{\prime}\}\rangle,$
with $n_{q}^{\prime}=n_{q}\pm 1$. Energy conservation requires
$\varepsilon(k^{\prime})=\varepsilon(k)\pm\hbar\omega_{q}+\Delta E_{\rm CM}.$
The corresponding Golden-Rule rate is
$W_{k\to k^{\prime};q}=\frac{2\pi}{\hbar}\,|M_{k,k^{\prime};q}|^{2}\,\delta\!\left(\varepsilon(k^{\prime})-\varepsilon(k)\mp\hbar\omega_{q}-\Delta E_{\rm CM}\right),$
with matrix elements scaling as
$M_{k,k^{\prime};q}\sim\alpha\sqrt{\frac{\hbar}{2M\omega_{q}}}\,F_{k,k^{\prime};q}.$
Multi-phonon and hybrid CM-phonon processes
When single-phonon processes are kinematically suppressed, higher-order multi-phonon processes and hybrid processes involving both CM recoil and internal phonon exchange are allowed. These channels relax the strict resonance condition Eq. (47) at the cost of reduced matrix elements.
Kinematic summary and temperature dependence
  • Elastic channels: exact backscattering $k\rightarrow-k$ without phonon excitation occurs via CM recoil and is enabled by translational invariance.
  • Inelastic channels: internal phonon emission and absorption become increasingly important as temperature increases and as phonon spectra permit energy matching.
  • Finite size: the discreteness of the phonon spectrum in a finite chain strongly constrains available inelastic channels and can enhance the relative importance of CM-mediated elastic processes.
Practical consequences
  • In wavepacket simulations, elastic backscattering events are identified by momentum reversal without change in internal phonon occupation, accompanied by a compensating shift in CM momentum.
  • Both bulk and boundary reflections obey the same momentum conservation law and are distinguished only by their spatial location.
  • The CM zero mode plays a central dynamical role and cannot be neglected without qualitatively altering the scattering physics.