### v2

# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

Abstract
In standard treatments of electron transport, momentum relaxation in a perfect, defect-free crystal is linked with phonon creation or annihilation. In this work, we reconsider this problem for a finite, isolated crystal, retaining the lattice center-of-mass (recoil) degree of freedom and enforcing conservation of total mechanical momentum together with discrete crystal pseudomomentum. Starting from the density–density form of the electron–lattice interaction, we show that an electron in the interior of a perfect crystal admits elastic momentum-transfer channels in which total momentum is conserved by recoil of the lattice background without phonon excitation. These elastic channels can provide the leading contribution to momentum relaxation. We further identify mixed quasielastic and superelastic processes in which phonon occupations change but do not account entirely for the electron’s momentum transfer. The elastic channels arise within the standard microscopic Hamiltonian and do not require additional disorder or defects. The resulting framework provides a complementary microscopic perspective on momentum relaxation in clean crystals, and is consistent with experimental phenomena such as weak localization, quantum oscillations, ultrasonic attenuation, and the observed separation of momentum and energy relaxation times.
preprint: APS/123-QED
I Introduction
Electron–lattice interactions play a central role in condensed matter physics. They govern both momentum relaxation and energy exchange in crystalline solids, and therefore underlie electrical resistivity and the loss of electronic coherence. Experimentally, however, momentum and energy relaxation often occur on markedly different timescales. In clean metals over broad temperature ranges, electrical resistivity and magnetotransport measurements indicate rapid momentum randomization, while hot-electron experiments, Johnson-noise thermometry, and related nonequilibrium probes show that energy relaxation to the lattice can remain comparatively slow. The resulting hierarchy
$\tau_{p}\ll\tau_{E},$ (1)
where $\tau_{p}$ is the momentum-relaxation time and $\tau_{E}$ the energy-relaxation time, indicates that electronic momentum can be efficiently scrambled without substantial net energy transfer to internal lattice excitations.
In conventional transport theory, momentum relaxation in a perfect, defect-free crystal is typically described in terms of phonon-mediated scattering processes. Within the Bloch–Grüneisen and semiclassical Boltzmann frameworks, transport is treated as a sequence of transitions between electronic states, with rates evaluated using the electron–phonon interaction [35, 3]. In the standard Fröhlich Hamiltonian [9], the interaction is expressed explicitly in terms of phonon creation and annihilation operators, and momentum transfer is therefore associated, at lowest order, with changes in phonon occupation. In this commonly employed description, elastic momentum relaxation is primarily attributed to impurities or other forms of explicit disorder.
The observed separation between $\tau_{p}$ and $\tau_{E}$ motivates a careful examination of how momentum conservation is implemented microscopically in a perfectly periodic crystal. The conceptual distinction between momentum relaxation and energy dissipation has long been recognized. Peierls emphasized that electrical resistance requires degradation of electronic (crystal) momentum rather than dissipation of electronic energy, and noted that the lattice, as a massive collective object, can absorb momentum with minimal energy cost [24]. Pippard similarly stressed that while defects often play an important role, momentum relaxation need not be uniquely tied to strongly inelastic processes [25].
In this work we revisit the problem starting from the same microscopic electron–lattice Hamiltonian, but retaining explicitly the lattice center-of-mass degree of freedom. In conventional treatments the crystal is effectively considered as a fixed reference frame, and the translational zero mode does not appear among the dynamical variables. Here we adopt an expanded viewpoint in which the crystal total mechanical momentum is retained. This does not modify the underlying Hamiltonian; rather, it makes explicit a degree of freedom that is commonly suppressed.
When total momentum conservation is enforced in this manner, a transfer of momentum $\mathbf{q}$ to an electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background. This exchange can occur without excitation of internal phonon modes, leading to an elastic recoil channel even in a perfectly periodic crystal. We also identify mixed quasielastic and superelastic processes in which phonon occupations change but do not fully account for the electron’s momentum transfer. These channels arise within the standard density–density form of the electron–lattice interaction [22, 9] and require no additional disorder.
The physical content of the elastic recoil channel is elementary. In any untethered composite object, an impulse imparted to one constituent immediately changes the total momentum of the whole. While redistribution of that momentum among internal degrees of freedom proceeds causally and at finite velocity, the collective momentum of the backgroud crystal appears in the wavefunction and is not dependent of the details of redistribution of the momentum among the constitutuents of the whole. . An instructive analogue is provided by Mössbauer scattering, in which the dominant elastic line is recoil of the entire lattice without internal phonon excitation.
A conceptual roadblock has been that, in contrast to processes such as Mössbauer and neutron scattering, where elastic, non-Bragg diffuse deflection is widely acknowledged, conduction electrons do not enter and exit; they start and stay in the crystal. The roadblock is addressed by assigning pseudomomentum to both a “foreground” electron density and a “background” consisting of everything else. We hasten to point out that the question here involves real physics, and not mere representation: Must phonons be created or destroyed when electrons deflect in a pure crystal, or not?
On the meaning of elastic scattering
Elastic scattering of interior electrons is most familiar in the context of impurity or defect scattering, where translational symmetry is explicitly broken. Here, we restore pseudomomentum to the status of a conserved momentum, and as in free space what then is conserved in such an elastic deflection? The answer will not be controversial: the total momentum of the whole system is conserved. What we believe is new is that two pseudmomomenta have changed equally and oppositely in such a collision: one belonging to the electron, another to the rest of the crystal. The total momentum, the sum of the two, is unchanged. In a perfectly periodic crystal, the relationship between changes in electronic pseudomomentum and the mechanical momentum of the lattice background is less often made explicit. In the present formulation this bookkeeping becomes transparent once (i) the lattice is treated as fully dynamical, (ii) pseudomomentum is assigned both to the electron and to the background in the absence of the electron density, and (iii) the interaction is expressed in density–density form, making the exchange of pseudomomentum explicit.
Finally, we clarify our terminology. In a finite-temperature many-body system, strictly elastic processes do not exist in an absolute sense. Even the dominant line in Mössbauer scattering exhibits slight thermal broadening. We therefore use the term elastic to denote processes in which phonon occupation does not change, and quasielastic to denote processes involving energy exchange that is small compared with the momentum transfer involved.
The landscape seen by an electron
At temperatures of order $T\sim 100\,$K, thermally excited acoustic distortions in a typical metal generate substantial deformation-potential landscapes on nanometer length scales. Using standard values for the deformation potential ($D\sim 5$$15\,\mathrm{eV}$), sound velocity ($v_{s}\sim 3$$5\times 10^{3}\,\mathrm{m/s}$), and the characteristic thermal phonon wavevector $q_{T}\sim k_{B}T/\hbar v_{s}$, one finds order-of-magnitude estimates for the resulting deformation-potential field gradients of
$|\nabla V_{\rm def}|\sim 10^{4}\text{–}10^{6}\ \mathrm{V/cm}\qquad(T\sim 100\,\mathrm{K}).$ (2)
Over a representative lateral scale of $L\sim 10\,$nm, this corresponds to peak-to-peak variations of the screened scalar potential of order
$\Delta V_{\rm pp}\sim|\nabla V_{\rm def}|\,L\sim 10\text{–}300\ \mathrm{mV}.$ (3)
[…]
II Foreground, Background, and Pseudomomentum
The conceptual distinction between momentum, crystal momentum, and electrical current has been recognized since the early development of band theory, but its implications for microscopic transport have remained subtle. In a particularly prescient analysis, Kohn emphasized that crystal momentum is a conserved quantity associated with translational symmetry, while the electrical current need not be proportional to it, and that momentum may be exchanged with the lattice without the necessity of phonon excitation or energy dissipation [17]. This observation already points to the possibility that momentum conservation can be satisfied through recoil of the lattice background rather than through inelastic scattering processes. What has remained largely implicit, however, is how this distinction should be implemented at the level of a local scattering event occurring deep in the interior of a crystal. Here, we explicitly assign pseudomomentum to both a foreground electron and the background lattice, with both pseudomomenta appearing as dynamical variables in the many-body wavefunction.
Making this partition explicit is a realization of Kohn’s insight, within a framework in which elastic deflection, coherent lattice recoil, and diffusive transport can coexist naturally in a perfect crystal. In what follows, we formalize a symmetric foreground-background structure, derive the elastic scattering amplitude within first-order perturbation theory, and compute the elastic fraction explicitly. We show that this fraction is close to unity for copper at room temperature. We also clarify how the present formulation extends the Fröhlich framework without contradicting its successful predictions, and discuss the implications for the interpretation of transport phenomena in crystalline solids.
We distinguish explicitly between foreground and background degrees of freedom. The foreground is the object of interest, here a single electron or a collection of them. The background consists of all remaining degrees of freedom, including the lattice atoms and any other electrons not treated explicitly. This distinction is essential for identifying the relevant symmetries and conserved quantities.
In a crystalline environment, the electronic Hamiltonian is invariant under discrete translations of the electron by lattice vectors, with the background held fixed. The associated conserved quantity is pseudomomentum (crystal momentum); for an electron in a periodic potential this is the familiar Bloch momentum $\hbar\mathbf{k}$.
Crucially, the same discrete translational symmetry applies to the background (see Fig. 1). If the electron (foreground A) coordinate is held fixed and the entire background B is translated by a lattice vector, the Hamiltonian is again unchanged. A consistent application of translational symmetry therefore assigns pseudomomentum not only to the electron but also to the background, even though only their sum corresponds to a mechanical momentum.
One might object that the problem can always be formulated in the rest frame of the background, with only the electron moving. However, this choice of frame obscures the conservation of total momentum of the combined system and suppresses the dynamical role of background recoil, which is essential for understanding local scattering events in the interior of a perfect crystal.
Figure 1: Sliding Blocks. Floating in space, two blocks attract one another and slide frictionlessly along their common long axis. At the top, both blocks carry conserved mechanical momenta, and clearly $P_{\rm tot}=p_{A}+p_{B}$. At the bottom, the same relation holds, but $p_{A}$ and $p_{B}$ are now pseudomomenta; their sum nevertheless equals the conserved mechanical momentum. The two blocks serve as surrogates for a foreground electron (A) in a perfect background crystal (B), illustrating that both the foreground and the background acquire pseudomomentum in each other’s presence.
Let $\hat{U}(\mathbf{h})$ translate all coordinates, foreground and background, by an arbitrary vector $\mathbf{h}\in\mathbb{R}^{3}$. For an isolated electron+background system,
[…]
so the corresponding Noether generator is the conserved total mechanical momentum, denoted by $\hat{\mathbf{Q}}$.
In the foreground/background formulation $\hat{\mathbf{Q}}$ is obtained as the sum of two subsystem pseudomomenta. Because $\hat{U}(\mathbf{h})$ factorizes into a translation of the foreground coordinate and an equal translation of the background configuration,
$\hat{U}(\mathbf{h})=\hat{U}_{\mathrm{f}}(\mathbf{h})\,\hat{U}_{\mathrm{b}}(\mathbf{h}),$ (5)
we have
$\hat{\mathbf{Q}}=\hat{\mathbf{P}}+\hat{\mathbf{p}}$ (6)
Here $\hat{\mathbf{p}}$ is the pseudomomentum carried by the foreground electron (the generator of its translation), and $\hat{\mathbf{P}}$ is the pseudomomentum carried by the background. (the generator of translating the background as a whole). Their sum $\hat{\mathbf{Q}}=\hat{\mathbf{P}}+\hat{\mathbf{p}}$ is the conserved mechanical momentum of the isolated system. For convenience, we can set it to 0; it never changes and we can now ignore it.
Even under the usual rule stating that defects, such as a crystal edge or impurity, scatter electrons elastically, there is a subtlety: the electron changes its pseudomomentum, but the lattice cannot the respond or recoil in kind. A pseudomomentum cannot be combined with a mechanical momentum. We have now seen, however, that both foreground and background have pseudomomentum, and can exchage it freely. Total momentum is conserved for example in the very real deflection of an electron at an edge by equal and opposite pseudomomentum changes.
These rules are analogous to those governing the collision of two objects in free space, except that here the interaction is periodic in their relative coordinate, apart from defects and lattice vibrations.
The total momentum $\hat{\mathbf{Q}}$ is conserved, but the foreground and background momenta, $\hat{\mathbf{p}}$ and $\hat{\mathbf{P}}$, undergo equal and opposite changes, in exact analogy with collisions in free space. This exchange does not require the creation or annihilation of internal lattice excitations and occurs immediately at the collision event (see Fig. 2). The very real recoil of the atom labeled in red in Fig. 2 may very well prove to have excited no vibrational quanta when measured in a number state basis. This is just the elementary fact that a slightly displaced ground state of a harmonic oscillator is still mostly in the ground state, if measured.
Two measurable subsystem momenta and the total momentum
Figure 1 shows two variations on a mechanical system. At the top, two smooth blocks interact frictionlessly. At the bottom, both blocks are corrugated, leading to periodic interaction.
Figure 2: The total momentum (black) remains constant as the foreground (electron, blue) and background (lattice, red) momenta jump abruptly at the collision time $t^{*}$. There is no delay in the total lattice momentum change, even though only a single atom initially carries the momentum. While this is elementary, we emphasize it because there is a common tendency to associate momentum flow directly with energy flow in the medium. Indeed there is momentum flow among particles, but the total is fixed immediately. The total appears in the wavefunctions.
Finite-window symmetry
In any physically realizable crystal the interaction is periodic only over a finite spatial extent $L$. Strict translational invariance is therefore absent at the boundaries, and crystal momentum is not an exact conserved quantity for the full finite system. This situation is directly analogous to the textbook particle in a box: the exact eigenstates are standing waves that do not carry definite momentum, yet traveling wave packets with well-defined average momentum can be constructed as superpositions of these eigenstates. Over times short compared with the traversal time to the boundary, such wave packets propagate as if momentum were conserved.
The same reasoning applies to a large but finite crystal. Although global translational symmetry is broken by the boundaries, the Hamiltonian is locally periodic in the bulk. Wave packets localized far from the boundaries evolve, for long but finite times, according to the translationally invariant bulk Hamiltonian. In this regime one may assign a well-defined crystal pseudomomentum to the packet, up to corrections that vanish as $L\to\infty$.
This type of approximate symmetry is standard in discussions of the thermodynamic limit, bulk transport theory, and wave-packet dynamics in finite systems. For clarity, we refer to it here as finite-window symmetry: translational invariance holds within a spatial and temporal window set by the system size. No new symmetry is implied; rather, this terminology emphasizes that the use of bulk momentum quantum numbers in a finite crystal is an approximation controlled by the ratio of the observation time and length scales to the system size.
In a macroscopic crystal this approximation is exceedingly accurate. Localized wave packets (“corpuscles”) can be constructed that propagate nearly as Bloch waves with well-defined pseudomomentum and experience long travel times before interacting with boundaries. This construction permits the use of the notation and results associated with Born–von Kármán boundary conditions as a calculational device, without assuming literal periodicity of the finite sample.
In a finite crystal, the exact normal modes are standing waves. As exact eigenmodes they have zero net mechanical momentum. Any “phonon carrying momentum $\hbar q$” is then not an eigenstate; it’s a wavepacket (a superposition of standing modes) that only behaves like a traveling object transiently until it reflects, mode-mixes, etc.
III Lattice degrees of freedom and zero modes
We consider a finite crystal consisting of atoms with coordinates $\hat{\mathbf{R}}_{j}$ and masses $m_{j}$, with total lattice mass $M=\sum_{j}m_{j}$, minus the foreground electron. We define the center-of-mass coordinate of the lattice background as
$\hat{\mathbf{R}}\equiv\frac{1}{M}\sum_{j}m_{j}\,\hat{\mathbf{R}}_{j}.$ (7)
It is convenient to decompose the atomic coordinates as
$\hat{\mathbf{R}}_{j}$ $=\hat{\mathbf{R}}+\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j},$ (8)
$\sum_{j}m_{j}\bm{\xi}^{(0)}_{j}$ $=\mathbf{0},\qquad\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$
where $\hat{\mathbf{R}}$ is the exact lattice center-of-mass coordinate, $\bm{\xi}^{(0)}_{j}$ denote fixed reference positions measured relative to the center of mass, and $\hat{\mathbf{u}}_{j}$ represent internal, dynamical lattice displacements. The mass-weighted constraints ensure that all internal degrees of freedom carry zero net momentum, so that $\hat{\mathbf{R}}$ describes the unique translational zero mode.
The classical and quantum theory of lattice vibrations explicitly separates rigid translation from internal vibrations. Born and Huang showed that the equations of motion of a finite crystal admit three zero-frequency normal modes corresponding to rigid translation of the crystal as a whole, while the internal normal modes satisfy the mass-weighted constraint $\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$ so that internal vibrations carry no net mechanical momentum [7]. All mechanical momentum of the crystal resides in the translational zero mode.
IV Electron-Lattice interaction
IV.1 Foreground and background Hamiltonians
. We write the total Hamiltonian as
$H=H_{\rm fg}+H_{\rm bg}+H_{\rm int},$ (9)
where $H_{\rm fg}$ describes the foreground electron, $H_{\rm bg}$ the background degrees of freedom, and $H_{\rm int}$ their coupling.
For a single foreground electron we take
$H_{\rm fg}=\frac{\hat{\mathbf{p}}^{2}}{2m},$ (10)
where $\hat{\mathbf{p}}$ denotes the generator of translations of the foreground coordinate $\hat{\mathbf{r}}$. The use of a bare kinetic-energy operator reflects the fact that the electron has not been pre-diagonalized into Bloch states. All effects of the periodic crystal potential, including band structure, pseudomomentum exchange, and scattering, are instead generated dynamically through the interaction Hamiltonian $H_{\rm int}$. This is the same starting point adopted by Fröhlich [9].
The remaining electrons of the metal are treated as part of the background. Together with the ions they establish the effective lattice potential, provide screening, and ensure mechanical stability. Their degrees of freedom are therefore implicit in $H_{\rm bg}$.
Density operator and exact factorization
The lattice density operator admits an exact factorization into a center-of-mass translation and an internal part:
$\hat{\rho}_{\mathbf{q}}$ $=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{\mathbf{q}},$ (11)
$\hat{\rho}^{\rm int}_{\mathbf{q}}$ $=\sum_{j}e^{-i\mathbf{q}\cdot(\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j})}.$
Here $\hat{\mathbf{R}}$ is the lattice center-of-mass operator, $\bm{\xi}^{(0)}_{j}$ are equilibrium positions, and $\hat{\mathbf{u}}_{j}$ are internal displacements. Equation (11) is an exact operator identity, independent of any choice of basis (plane waves, Bloch waves, or standing waves).
This identity expresses a purely kinematic fact: a rigid translation of the lattice multiplies every density Fourier component by the phase $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, while all internal lattice dynamics are contained entirely in $\hat{\rho}^{\rm int}_{\mathbf{q}}$.
Density–density interaction and pseudomomentum conservation
The electron–lattice interaction may be written in density–density form as
$H_{\rm int}=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}\,\hat{\rho}_{-\mathbf{q}},$ (12)
following the formalism of Van Hove [30]. No assumption about plane-wave eigenstates is required; $\mathbf{q}$ labels Fourier components of the interaction kernel.
Using Eq. (11), this becomes
$H_{\rm int}$ $=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i{\mathbf{q}}\cdot\hat{\bm{\rho}}}\,\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (13)
$=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}.$
[…]
Crucially, this exchange of pseudomomentum applies whether or not there is a change a change in phonon occupation numbers.
Elastic matrix elements of $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ are generically nonzero. The background zero mode is therefore capable of absorbing finite pseudomomentum with negligible energy cost, providing the kinematic basis for elastic but momentum-relaxing scattering. Any combination of phonon involvement and lattice background recoil is allowed. It seems this should matter for measurable properties.
V Elastic internal scattering and the Debye-Waller fraction
In this section we show that, even within a strictly perturbative treatment based on Bloch electrons and phonon number states, robust elastic scattering channels exist already at first order in the electron–lattice coupling. These channels become explicit once the full lattice density operator is retained and the translational zero mode of the lattice is treated as a dynamical degree of freedom. The resulting processes are elastic with respect to internal lattice excitations, despite involving momentum exchange with the lattice as a whole.
Density–density coupling and emergence of the Bloch potential
We take the foreground–background interaction in density–density form,
$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}),$ (14)
with the tagged-electron density operator
$\rho_{\rm fg}(\mathbf{q})=e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}$. (15)
For the background we use the exact center-of-mass factorization
$\rho_{\rm bg}(-\mathbf{q})=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (16)
where $\hat{\mathbf{R}}_{j}=\hat{\mathbf{R}}+\bm{\xi}_{j}^{(0)}+\hat{\mathbf{u}}_{j}$.
Separating the background density operator into its equilibrium expectation value (taken in the background center-of-mass frame) and fluctuations,
$\rho_{\rm bg}(-\mathbf{q})=\langle\rho_{\rm bg}(-\mathbf{q})\rangle+\delta\rho_{\rm bg}(-\mathbf{q}),$ (17)
the static component produces an effective periodic potential acting on the foreground electron,
$U_{\rm Bloch}(\hat{\mathbf{r}})=\sum_{\mathbf{G}}V(\mathbf{G})\,e^{i\mathbf{G}\cdot\hat{\mathbf{r}}}\,\langle\rho_{\rm bg}(-\mathbf{G})\rangle,$ (18)
where $\mathbf{G}$ are reciprocal lattice vectors.
The corresponding single-particle foreground Hamiltonian becomes
[…]
which is the usual Bloch Hamiltonian. This periodic potential is not assumed a priori, but emerges as the static mean-field component of the underlying density–density coupling.
The remaining part of $H_{\rm int}$ couples the foreground electron to background density fluctuations $\delta\rho_{\rm bg}$, but with the crucial difference from the clamped-lattice formulation: the full operator $\rho_{\rm bg}(-\mathbf{q})$ retains the explicit center-of-mass translation factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, which enforces exact momentum bookkeeping in the finite, untethered crystal.
Initial and final states
For a perturbative treatment we consider product states of foreground and background,
$\Psi_{i}=\Psi_{n\mathbf{k}}(\mathbf{r})\,\Psi^{\rm bg}_{i}(\{\mathbf{u}_{j}\}),\qquad\Psi_{f}=\Psi_{n^{\prime}\mathbf{k}^{\prime}}(\mathbf{r})\,\Psi^{\rm bg}_{f}(\{\mathbf{u}_{j}\}),$ (20)
where $\Psi^{\rm bg}_{i,f}$ are phonon number states or superpositions thereof. The electronic states are Bloch eigenstates of the static lattice potential.
Using the factorized form of the interaction, any nonvanishing matrix element $\langle\Psi_{f}|H_{\rm int}|\Psi_{i}\rangle$ implies the pseudomomentum transfers
$\Delta\mathbf{p}=+\hbar\mathbf{q},\qquad\Delta\mathbf{P}=-\hbar\mathbf{q},$ (21)
so that the total mechanical momentum of the combined system is conserved. The recoil pseudomomentum is carried by the lattice center-of-mass degree of freedom, regardless of phonon number changes.
Elastic internal channel and the Debye-Waller factor
For a harmonic lattice, the internal displacements admit the normal-mode expansion
$\hat{\mathbf{u}}_{j}=\frac{1}{\sqrt{N}}\sum_{\mathbf{k},\lambda}\sqrt{\frac{\hbar}{2M\omega_{\mathbf{k}\lambda}}}\,\mathbf{e}_{\mathbf{k}\lambda}\left(a_{\mathbf{k}\lambda}e^{i\mathbf{k}\cdot\bm{\xi}^{(0)}_{j}}+a^{\dagger}_{\mathbf{k}\lambda}e^{-i\mathbf{k}\cdot\bm{\xi}^{(0)}_{j}}\right)$. (22)
The operator $e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}$ therefore contains all orders in the phonon operators and automatically generates elastic, quasielastic, and inelastic processes.
In particular, in a phonon number state basis the diagonal matrix elements are generically nonzero,
$\left\langle\{n\}\left|e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\right|\{n\}\right\rangle\neq 0,$ (23)
so that a scattering channel strictly elastic with respect to internal lattice excitations is present already at first order in $H_{\rm int}$. Upon thermal averaging (or in the ground state), this diagonal factor reduces to the Debye-Waller form
$\left\langle e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\right\rangle_{T}=\exp[-W(\mathbf{q},T)],$ (24)
which is the direct analog of the recoil-free (zero-phonon-line) fraction in Mössbauer spectroscopy and in neutron or x-ray scattering.
While much stronger coupling regimes can be treated nonperturbatively—for example within wave-on-wave (WoW) simulations [4] the present analysis shows that even within standard perturbation theory there exist elastic scattering processes that are absent when the center of mass is hidden or eliminated.
Relation to the Fröhlich truncation
In the original Fröhlich formulation, the exact factorization Eq. (11) is not employed. The lattice center-of-mass coordinate $\hat{\mathbf{R}}$ is not retained explicitly, which is equivalent to setting the recoil factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ to unity and treating the lattice as a clamped reference frame. At the same time, the internal displacement exponential is truncated to linear order,
$e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\rightarrow 1-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}.$ (25)
The zeroth-order term reproduces the static periodic potential responsible for Bloch band formation, while the linear term generates single-phonon processes.
Although this procedure is internally consistent and highly successful in the weak-coupling regime, it removes the background recoil zero mode by construction. As a result, even if the full exponential in $\hat{\mathbf{u}}_{j}$ were retained, the truncated theory would still lack an explicit degree of freedom capable of absorbing the total recoil momentum of the lattice. The standard Fröhlich Hamiltonian therefore describes momentum exchange with internal lattice modes, but the 0 mode, the lattice recoil as a whole, is absent.
Thermal averaging and the Debye-Waller fraction
In thermal equilibrium the internal lattice degrees of freedom are described by the density matrix
$\hat{\varrho}=\frac{e^{-\beta\hat{H}_{\mathrm{lat}}}}{Z}=\sum_{n}p_{n}\,|n\rangle\langle n|,\qquad p_{n}=\frac{e^{-\beta E_{n}}}{Z},$ (26)
where $|n\rangle$ denote exact many-body eigenstates of the lattice Hamiltonian $\hat{H}_{\mathrm{lat}}$ with energies $E_{n}$. For elastic internal scattering the relevant quantity is the thermal expectation value of the internal density operator,
$\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}=\sum_{n}p_{n}\,\langle n|\rho^{\rm int}_{\mathbf{q}}|n\rangle.$ (27)
The coherent elastic scattering intensity is
$I_{\mathrm{el}}(\mathbf{q})=\big|\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}\big|^{2}.$ (28)
The total scattering intensity summed over all final internal states is
$I(\mathbf{q})=\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}.$ (29)
The elastic (coherent) branching fraction is therefore
$f_{\mathrm{el}}(\mathbf{q},T)=\frac{\big|\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}\big|^{2}}{\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}}.$ (30)
For a harmonic lattice one finds[18, 21]
$f_{\mathrm{el}}(\mathbf{q},T)\simeq e^{-2W(\mathbf{q},T)},\qquad W(\mathbf{q},T)=\frac{1}{2}\left\langle(\mathbf{q}\cdot\hat{\mathbf{u}})^{2}\right\rangle_{T},$ (31)
the familiar Debye–Waller result. Even at finite temperature a substantial fraction of momentum-transfer events leave the lattice in the same internal state.
Estimate of the elastic internal fraction at representative momentum transfers
For momentum relaxation in a metal it is natural to consider momentum transfers of order the Fermi momentum, since large-angle scattering dominates the transport relaxation rate through the usual $(1-\cos\theta)$ weighting. For copper, a free-electron estimate yields
$k_{F}\simeq 1.36\ \text{\AA }^{-1},\qquad 2k_{F}\simeq 2.7\ \text{\AA }^{-1}.$ (32)
We therefore evaluate the elastic internal (phonon-diagonal) fraction $f_{\mathrm{el}}(\mathbf{q},T)$ at representative values $q\ll k_{F}$, $q\sim k_{F}$, and $q\sim 2k_{F}$.
Within the isotropic Debye-Waller estimate used above,
$2W(\mathbf{q},T)=\frac{q^{2}\langle u^{2}\rangle_{T}}{3},\qquad f_{\mathrm{el}}(\mathbf{q},T)\simeq e^{-2W(\mathbf{q},T)}.$ (33)
Using the room-temperature mean-square displacement for copper, $\langle u^{2}\rangle_{300\mathrm{K}}\simeq 1.9\times 10^{-2}\ \text{\AA }^{2}$, one finds the values summarized in Table 1.
qq-1) $2W$ $f_{\mathrm{el}}$
$0.3$ $5.7\times 10^{-4}$ $0.9994$
$k_{F}\simeq 1.36$ $1.2\times 10^{-2}$ $0.988$
$2k_{F}\simeq 2.7$ $4.6\times 10^{-2}$ $0.955$
Table 1: Estimated elastic internal fraction $f_{\mathrm{el}}$ for copper at room temperature ($T=300$ K) for representative momentum transfers. Even for large-angle scattering with $q\sim 2k_{F}$, the majority of momentum-relaxing events remain phonon-diagonal, with momentum absorbed predominantly by the lattice background rather than by internal phonon excitation.
These estimates show that, for momentum transfers relevant to transport, the overwhelming fraction of scattering events are internally elastic, in the sense that they leave the lattice in the same vibrational state. Phonon creation constitutes only a minor branching channel at room temperature, even for large-angle scattering. Momentum exchange with the lattice therefore occurs predominantly through elastic or quasi-elastic coupling to the background degrees of freedom, with internal phonon excitation representing a secondary correction.
VI Experiments suggesting large elastic contributions
Experimental evidence for predominantly elastic momentum relaxation in metals at low temperature has accumulated for decades, particularly in the mesoscopic-transport and quantum-interference literature. In this section we show that the framework developed here resolves several longstanding disparities between theory and experiment regarding electronic transport. We review a range of experiments that demonstrate a clear separation between momentum relaxation and energy relaxation, consistent with elastic or quasi-elastic scattering dominating momentum randomization.
Hierarchy of momentum and energy relaxation times
Electrical resistivity and optical conductivity probe momentum relaxation, quantum-interference phenomena such as weak localization probe phase coherence, and nonequilibrium measurements including shot noise, Johnson-noise thermometry, and hot-electron relaxation directly access energy exchange. Strikingly, these probes consistently reveal a strong hierarchy of timescales,
$\tau_{\mathrm{p}}\ll\tau_{E},$ (34)
established using independent experimental techniques and across a wide range of materials. Here $\tau_{\mathrm{p}}$ is the timescale on which electronic momentum (or pseudomomentum) is randomized by scattering, while $\tau_{E}$ is the timescale over which electrons exchange energy with the background and relax toward thermal equilibrium.
The consistent appearance of this hierarchy demonstrates that momentum relaxation and energy relaxation are distinct physical processes. While conventional language often associates momentum randomization with inelastic scattering, these experiments show that electronic momentum can be efficiently scrambled by a time-dependent environment without significant energy exchange. The lattice need not act as an energy sink in order to act as a momentum sink. This separation is precisely what is expected when electrons scatter elastically from a dynamically fluctuating lattice background.
Quantum interference and dephasing
The suppression of quantum-interference effects with increasing temperature is commonly described in terms of inelastic scattering processes that exchange energy with environmental degrees of freedom [20, 1]. It is useful, however, to distinguish between dephasing as operationally defined in transport experiments and irreversible decoherence in the strict quantum-mechanical sense.
A broad class of experiments indicates that substantial suppression of interference can occur even when direct energy relaxation to the lattice remains weak. Mesoscopic interference measurements, for example, show a pronounced reduction— and in some cases an apparent saturation—of the phase-coherence time $\tau_{\phi}$ at low temperatures, while independent probes suggest that electron–phonon energy exchange rates continue to decrease in the same regime [23, 15, 31]. This empirical separation of timescales illustrates that phase and momentum randomization need not scale directly with the rate of substantial energy transfer to internal lattice excitations.
Weak-localization is robust in experiments, implying that the backscattering is coherent and therefore elastic. [20, 23].Elastic scattering amplitudes, including diffuse non-Bragg amplitudes, are reduced by the well known Debye–Waller factor.
Hot-electron relaxation
A particularly direct separation of momentum and energy relaxation is provided by hot-electron experiments in metals [32, 26, 10, 11]. In these measurements, nonequilibrium electron distributions are injected, and the subsequent momentum randomization and energy loss are probed independently. Over broad temperature ranges, the momentum-relaxation time inferred from transport or magnetotransport is found to be substantially shorter than the energy-relaxation time associated with thermalization. Electrons therefore undergo many momentum-deflecting collisions while remaining close to isoenergetic on the same timescale.
This pronounced separation of timescales indicates that efficient momentum randomization does not require comparably rapid energy transfer to lattice excitations. While inelastic electron–phonon scattering ultimately governs thermalization, the observed hierarchy suggests that additional scattering channels, with weak net energy exchange on electronic scales, can play an important role in momentum relaxation within the metallic state.
VII Discussion
The purpose of this section is twofold. First, we examine the consistency of the present framework with well-established experimental results, including the Wiedemann–Franz law, weak localization, and quantum oscillations. These considerations address potential concerns regarding the role of predominantly elastic momentum-relaxing channels. Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott–Ioffe–Regel crossover. The latter discussion is exploratory and is presented as a coherent physical interpretation rather than as a definitive claim.
VII.1 Implications for the Lorenz ratio
Elastic scattering randomizes momentum as effectively as inelastic scattering within standard transport formalisms. Accordingly, a predominantly elastic microscopic channel does not conflict with the established success of conventional transport theory.
In a degenerate metal, the Wiedemann–Franz (WF) law states that the ratio of the electronic thermal conductivity $\kappa$ to the electrical conductivity $\sigma$ times temperature approaches the Sommerfeld value
$L\equiv\frac{\kappa}{\sigma T}\;\xrightarrow[T\to 0]{}\;L_{0}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}.$ (35)
Although some inelastic processes must ultimately relax the electronic energy current and establish global thermal equilibrium, it has long been understood that these processes need not determine the leading transport coefficients themselves [35, 19, 27].
If the dominant momentum-relaxing channel is elastic and only weakly energy dependent over the thermal window $\sim k_{B}T$ about the Fermi level, then both electrical and thermal currents are governed by the same relaxation time $\tau_{\mathrm{el}}$,
$\sigma\propto\tau_{\mathrm{el}},\qquad\kappa\propto\tau_{\mathrm{el}},$ (36)
and the Sommerfeld value $L_{0}$ follows in the usual way. Inelastic processes with rate $1/\tau_{\mathrm{in}}\ll 1/\tau_{\mathrm{el}}$ transfer energy between electrons and lattice degrees of freedom but do not enter the leading expressions for $\sigma$ or $\kappa$. A transport regime dominated by elastic momentum randomization is therefore fully compatible with the WF law, even when true energy relaxation is parametrically slower.
VII.2 Weak localization
Weak localization (WL) provides a clear illustration of the central role of elastic backscattering in low-temperature metallic transport. The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that revisit the same spatial region after multiple scattering events [6, 20]. The WL correction therefore requires that a substantial fraction of backscattering processes preserve phase coherence over the relevant timescale.
If scattering were strongly inelastic on the transport timescale, interference between time-reversed paths would be suppressed before a WL correction could develop. Experimentally, however, a well-defined suppression of conductivity and characteristic negative magnetoresistance are observed at low temperature. As temperature increases, the WL correction is gradually reduced. Although this reduction is commonly described in terms of increasing “inelastic scattering,” what is operationally measured is the loss of phase coherence rather than a rapid increase in electronic energy relaxation.
Independent probes, including hot-electron relaxation and Johnson-noise thermometry, indicate that electron–phonon energy exchange can remain comparatively slow in the same temperature range. The suppression of WL is therefore consistent with the onset of a time-dependent scattering environment that randomizes phase while involving only modest energy transfer on electronic scales.
Within the present framework this behavior fits naturally into the hierarchy of timescales
$\tau_{p}\ll\tau_{\phi}\ll\tau_{E},$
where elastic momentum randomization produces diffusive transport, while slower temporal evolution of the lattice background reduces phase coherence without requiring substantial phonon excitation or rapid energy dissipation.
VII.3 Shubnikov–de Haas oscillations
Quantum oscillation phenomena such as the Shubnikov–de Haas and de Haas–van Alphen effects provide an independent probe of electronic scattering in metals at low temperature [29, 2]. These oscillations arise from the quantization of electronic orbits in a magnetic field and reflect coherent motion around the Fermi surface over many cyclotron periods.
The oscillation amplitude is controlled by the Dingle factor [8],
$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right),$ (37)
where $\tau_{q}$ is the quantum lifetime associated with Landau-level broadening. This lifetime is sensitive to momentum-randomizing processes that broaden Landau levels, while sufficiently strong inelastic energy exchange suppresses the coherence required to complete cyclotron motion. Scattering events that transfer energies comparable to $\hbar\omega_{c}$ disrupt the phase coherence of the orbit.
The observation of well-defined quantum oscillations over broad temperature ranges and down to relatively low magnetic fields therefore indicates that electrons can undergo many momentum-deflecting collisions while maintaining phase coherence over multiple cyclotron periods. As in weak localization, the temperature dependence of the oscillation amplitude reflects phase smearing due to temporal fluctuations of the scattering environment. These considerations are consistent with a regime in which momentum randomization can occur without rapid or large energy exchange on electronic scales.
VII.4 Planckian diffusion and linear-in-$T$ resistivity
Much recent literature discusses Planckian behavior in terms of a microscopic relaxation time
$\tau_{\mathrm{Pl}}\sim\frac{\hbar}{k_{B}T},$
often interpreted as setting a characteristic scale for scattering or energy dissipation [33, 13, 12, 28]. Within this perspective, transport coefficients are commonly estimated by inserting $\tau_{\mathrm{Pl}}$ into kinetic or hydrodynamic expressions.
The framework developed in Ref. [5, 34] and extended here suggests a complementary interpretation. The primary microscopic result is the emergence of real-space diffusion with diffusion constant
$D\sim\frac{\hbar}{m^{\ast}},$ (38)
arising from elastic momentum scrambling in a time-dependent lattice background. In this formulation, diffusion emerges from quantum dynamics in a fluctuating potential and does not rely explicitly on strong inelastic scattering or rapid energy dissipation. Wave-on-wave simulations of the coupled electron–lattice problem demonstrate this mechanism explicitly [5, 16, 34].
Once diffusion is established, Einstein relations connect $D$ to transport coefficients,
$\sigma=\chi D,\qquad D=\mu\frac{k_{B}T}{e},$ (39)
so that expressing $D$ in terms of an effective relaxation time yields [5]
$\tau_{\mathrm{eff}}\sim\frac{\hbar}{k_{B}T},$ (40)
up to factors of order unity. From this viewpoint, the Planckian timescale may be understood not as a fundamental bound on inelastic scattering, but as an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics. It characterizes the rate at which charge spreads spatially rather than the rate at which energy must be irreversibly transferred to internal degrees of freedom.
VIII Summary
In the conventional clamped-lattice formulation of electron–phonon scattering, the lattice is treated as a fixed reference frame and its translational zero mode does not appear explicitly among the dynamical variables. Within that representation, momentum transfer is naturally described in terms of internal phonon excitations. When the lattice center-of-mass degree of freedom is retained explicitly, an additional elastic momentum-transfer channel becomes manifest. In a finite crystal, pseudomomentum can be exchanged between an electron and the lattice background while total mechanical momentum is conserved through recoil of the lattice zero mode. Such processes do not require phonon creation and may contribute significantly to momentum relaxation at low and intermediate temperatures.
Making this channel explicit requires adopting a formulation in which (i) the lattice is treated as a finite, mechanically isolated object so that its center-of-mass motion remains dynamical, (ii) discrete translational symmetry is implemented symmetrically, with foreground and background pseudomomenta treated on equal footing, and (iii) the electron–lattice interaction is retained in its full density–density form. These steps do not alter the microscopic Hamiltonian; rather, they retain degrees of freedom that are commonly suppressed by choice of reference frame. When treated in this manner, elastic momentum exchange between electron and lattice background appears naturally within the same theoretical structure.
In this framework, pseudomomentum is redistributed between the electron and the lattice background, while total mechanical momentum is conserved through recoil of the translational zero mode. Momentum relaxation need not coincide with phonon creation, and energy dissipation need not set the transport rate. This separation of roles clarifies how efficient momentum randomization can coexist with comparatively slow energy relaxation.
Restoring explicit center-of-mass dynamics provides a more complete description of momentum bookkeeping in finite crystals. It offers a unified interpretation of experimental observations that indicate robust elastic scattering in clean materials, including weak localization, quantum oscillations, and ultrasonic attenuation. More broadly, the results suggest that in certain regimes, including some materials exhibiting linear-in-$T$ resistivity, momentum relaxation may be governed primarily by elastic scattering in a time-dependent lattice background [5], while inelastic processes play a secondary role in establishing global thermal equilibrium.
In a perfectly clean, isolated, translation-invariant crystal, total momentum is conserved exactly. Such a system may exhibit internal equilibration and decay of electronic current, but it cannot exhibit finite DC resistivity in the strict steady-state sense. True resistivity requires a mechanism for removing total momentum from the system.
IX Acknowledgments
E. Heller thanks B. Halperin for many discussions that provided a thorough understanding of the traditional theory. He also thanks Prof. Martina Hentschel for insightful comments that have improved the manuscript. This work was supported by the U.S. Department of Energy, Office of Science, under Grant No. DE-SC0025489.
Figure 3: Isolated systems of mass $M+m$ containing a foreground particle of mass $m$ and a background of mass $M$. The particle is attracted to the background and moves frictionlessly, except in C and D it may excite vibrations of the chain. In A, the particle is free except for walls. In B, the particle is in a fixed periodic potential within its host, the box. In C we replace the fixed corrugation with a chain of masses connected by springs. In D, we dispense with the box, coming close to the situation that matters here. Of course the box could be sitting on a table, which could be fixed to a floating laboratory, which could be upon something so massive that it has significant gravitational attraction, yet that too is floating in space…
Appendix A Particle in a Free Box
We consider simple models that clarify the treatment of momentum. The usual particle-in-a-box problem fixes the walls and therefore violates momentum conservation. Here the box is given finite mass and allowed to move freely, so that total momentum is exactly conserved.
Consider a one-dimensional rigid box of mass $M$ and length $L$ with center-of-mass coordinate $R$, containing a particle of mass $m$ with coordinate $r$ (Fig. 3).
Introduce total and relative coordinates
$\mathcal{R}=\frac{mr+MR}{m+M},\qquad\rho=r-R,$
with total mass $M_{T}=m+M$ and reduced mass
$\mu=\frac{mM}{m+M}.$
The hard-wall constraint is $|\rho|<L/2$.
The Hamiltonian separates exactly:
$H=-\frac{\hbar^{2}}{2M_{T}}\frac{\partial^{2}}{\partial\mathcal{R}^{2}}+\left[-\frac{\hbar^{2}}{2\mu}\frac{\partial^{2}}{\partial\rho^{2}}+V_{\rm hw}(\rho)\right].$
Total momentum is conserved:
$\hat{P}_{\rm tot}=-i\hbar\frac{\partial}{\partial\mathcal{R}}.$
Case A: Flat interior
The relative Hamiltonian is
$H_{\rm rel}=-\frac{\hbar^{2}}{2\mu}\frac{d^{2}}{d\rho^{2}}+V_{\rm hw}(\rho).$
The normalized eigenfunctions are
$\phi_{n}(\rho)=\sqrt{\frac{2}{L}}\sin\!\left[\frac{n\pi}{L}\left(\rho+\frac{L}{2}\right)\right],\qquad n=1,2,\dots,$
with energies
$E_{n}=\frac{\hbar^{2}\pi^{2}n^{2}}{2\mu L^{2}}.$
Any internal state may be written
$\psi(\rho,t)=\sum_{n=1}^{\infty}c_{n}\,\phi_{n}(\rho)\,e^{-iE_{n}t/\hbar}.$
A Gaussian superposition centered at $\rho_{0}$ with mean wavenumber $k_{0}$,
$c_{n}=\mathcal{N}\exp\!\left[-\frac{(n-n_{0})^{2}}{4\sigma_{n}^{2}}\right]\exp\!\left[-ik_{n}\rho_{0}\right],\qquad k_{n}=\frac{n\pi}{L},$
produces a localized packet
$\psi(\rho,0)\approx A(\rho-\rho_{0})e^{ik_{0}(\rho-\rho_{0})}.$
Although each $\phi_{n}$ carries no relative momentum, the particle and box momenta may be nonzero individually, while $P_{\rm tot}$ remains fixed. At a wall collision both reverse sign, preserving total momentum.
Case B: Periodic interior
Let the interior potential be periodic relative to the box:
$V_{\rm per}(\rho)=V_{\rm per}(\rho+a).$
Then
$H_{\rm rel}=-\frac{\hbar^{2}}{2\mu}\frac{d^{2}}{d\rho^{2}}+V_{\rm per}(\rho)+V_{\rm hw}(\rho).$
Eigenfunctions take Bloch standing-wave form
$\varphi_{bn}(\rho)=\sqrt{\frac{2}{L}}u_{bn}(\rho)\sin\!\left(\frac{n\pi}{L}(\rho+L/2)\right),$
with
$u_{bn}(\rho+a)=u_{bn}(\rho).$
Wavepackets are superpositions
$\psi(\rho,0)=\sum_{n=1}^{\infty}A_{n}\varphi_{bn}(\rho).$
Shared pseudomomentum
Because $\rho=r-R$,
$e^{iq\rho}=e^{iqr}e^{-iqR}.$
Internal pseudomomentum $\hbar q$ is therefore shared between particle and box. For fixed total momentum $P_{0}$,
$p_{r}=\frac{m}{m+M}P_{0}+\hbar q,\qquad P_{R}=\frac{M}{m+M}P_{0}-\hbar q,$
so $p_{r}+P_{R}=P_{0}$.
Cases C and D: Dynamical periodic background
Let the periodic background be a dynamical elastic chain with atomic coordinates $R_{j}$:
$H=\frac{p_{r}^{2}}{2m}+\sum_{j}\frac{p_{j}^{2}}{2M}+\frac{K}{2}\sum_{j}(R_{j+1}-R_{j}-a)^{2}+H_{\rm int}.$
With density–density coupling
$H_{\rm int}=\int dr\,dr^{\prime}\,\rho_{p}(r)V(r-r^{\prime})\rho_{\rm lat}(r^{\prime}),$
where
$\rho_{p}(r)=\delta(r-\hat{r}),\qquad\rho_{\rm lat}(r)=\sum_{j}\delta(r-\hat{R}_{j}),$
the Hamiltonian depends only on coordinate differences and is invariant under rigid translation
$r\rightarrow r+a_{0},\qquad R_{j}\rightarrow R_{j}+a_{0}.$
Thus total momentum
$P_{\rm tot}=p_{r}+\sum_{j}p_{j}$
is exactly conserved. Diagonalizing the lattice,
$R_{j}=R_{\rm lat}+\sum_{q}u_{q}e^{iqja},$
reveals the zero mode $R_{\rm lat}$. The density operator contains the factor
$e^{-iqR_{\rm lat}},$
showing explicitly that momentum $\hbar q$ transferred to the particle is balanced by recoil of the lattice center of mass. Internal phonons redistribute momentum but are not required for momentum conservation.
Appendix B Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
We consider a finite one-dimensional tight-binding chain of $N$ identical atoms, isolated in space, with nearest-neighbor springs and an electron moving in a basis of localized orbitals. The chain possesses an exact translational zero mode (center-of-mass motion). An electron can therefore exchange momentum with the lattice as a whole without necessarily exciting internal phonon modes. This opens a phonon-diagonal (internally elastic) backscattering channel for an electronic wavepacket, with the required momentum reversal balanced by center-of-mass recoil.
Hamiltonian and separation of the CM mode
Write the atomic coordinates as
$x_{n}=R+na+u_{n},\qquad\sum_{n=1}^{N}u_{n}=0,$ (41)
where $R$ is the center-of-mass coordinate, $a$ the equilibrium lattice spacing, and $\{u_{n}\}$ the internal displacements. The lattice Hamiltonian separates into CM and internal parts,
$H_{\rm lat}=\frac{P_{R}^{2}}{2M_{\rm tot}}+\sum_{n=1}^{N}\frac{p_{n}^{2}}{2M}+\frac{K}{2}\sum_{n=1}^{N-1}(u_{n+1}-u_{n})^{2},$ (42)
with $M_{\rm tot}=NM$ and $\sum_{n}p_{n}=0$.
The electron is described by a nearest-neighbor tight-binding Hamiltonian,
$H_{e}=-\sum_{n=1}^{N-1}t_{n}\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right),$ (43)
with bond-length-dependent hopping
$t_{n}\equiv t(x_{n+1}-x_{n})\simeq t_{0}-\alpha(u_{n+1}-u_{n}).$ (44)
Thus the electron–lattice coupling depends only on internal coordinates $u_{n+1}-u_{n}$; the CM coordinate $R$ drops out exactly.
Phonon-diagonal backscattering via CM recoil
Because the full electron+lattice Hamiltonian is translationally invariant, the total momentum is exactly conserved. Since the band energy satisfies $\varepsilon(k)=\varepsilon(-k)$, electronic reversal $k\to-k$ does not change the electronic energy.
Consider an incoming wavepacket narrowly peaked at +k+k, an internal lattice eigenstate $|\nu\rangle$, and a CM momentum eigenstate $|P_{R}\rangle$. A phonon-diagonal backscattering channel has the form
$|k\rangle\otimes|\nu\rangle\otimes|P_{R}\rangle\longrightarrow|-k\rangle\otimes|\nu\rangle\otimes|P_{R}+2\hbar k\rangle,$ (45)
i.e. the internal state is unchanged while the CM absorbs the recoil $2\hbar k$.
The associated change in CM kinetic energy is
$\Delta E_{\rm CM}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{4\hbar k\,P_{R}+4\hbar^{2}k^{2}}{2M_{\rm tot}}.$ (46)
Exact energy conservation with strictly unchanged internal state requires $\Delta E_{\rm CM}=0$, i.e.
$P_{R}=-\hbar k.$ (47)
For generic $P_{R}$ there is a recoil-energy mismatch
$\Delta E_{\rm recoil}=\frac{(P_{R}+2\hbar k)^{2}-P_{R}^{2}}{2M_{\rm tot}}=\frac{2\hbar k\,(P_{R}+\hbar k)}{M_{\rm tot}}.$ (48)
However, for a macroscopic chain $M_{\rm tot}\propto N$ so $|\Delta E_{\rm recoil}|\sim 1/N$ becomes parametrically small. For a wavepacket of spatial extent $L$ with intrinsic energy width $\Delta E_{\rm pkt}\sim\hbar v_{g}/L$, or for scattering occurring over a finite time $\tau_{\rm scatt}$ with uncertainty $\hbar/\tau_{\rm scatt}$, the process is effectively elastic whenever $\Delta E_{\rm recoil}$ is below the available energy resolution. In a number-state phonon description, any residual mismatch may be accommodated by very soft internal modes, but the momentum reversal $k\to-k$ is carried by CM recoil; internal excitations, if present, serve primarily as an energetically low-cost bookkeeping channel.
Role of finite temperature
At finite temperature $T$, the CM momentum is thermally distributed,
$f(P_{R})\propto\exp\!\left(-\frac{P_{R}^{2}}{2M_{\rm tot}k_{B}T}\right),$ (49)
with width $\sigma_{P}=\sqrt{M_{\rm tot}k_{B}T}$. The resonant value $P_{R}=-\hbar k$ therefore occurs with nonzero probability density for any $T>0$, so strictly phonon-diagonal backscattering events exist at finite temperature in an isolated chain.
Bulk and boundary backscattering
The CM-recoil channel does not require the electron to reach the ends: it can occur in the bulk from the static component of thermal distortions (phonon-diagonal scattering). The SSH coupling
$H_{\rm int}=\alpha\sum_{n=1}^{N-1}(u_{n+1}-u_{n})\left(c_{n+1}^{\dagger}c_{n}+c_{n}^{\dagger}c_{n+1}\right)$
connects $|k\rangle$ and $|-k\rangle$ through the $2k$ Fourier component of the bond-distortion field. CM recoil is enforced by the CM translation factor. In particular, the CM shift operator produces
$\langle P_{R}^{\prime}|e^{-i(2k)R}|P_{R}\rangle\propto\delta(P_{R}^{\prime}-P_{R}-2\hbar k),$ (50)
so momentum conservation is implemented by a delta function (or Kronecker delta in a finite quantization volume), not by an inner product between distinct CM momentum eigenstates.
Free boundaries can also reflect a wavepacket once it reaches an end. Such reflections reverse $k$ and are likewise accompanied by CM recoil. Boundaries provide a location for reflection but do not constitute a momentum sink; both bulk (distortion-mediated) and boundary reflections obey the same conservation principle enforced by the CM zero mode.
Other scattering channels
In addition to the phonon-diagonal CM-recoil channel, inelastic and hybrid processes are allowed.
Single-phonon emission and absorption
If momentum is transferred to internal modes, the internal lattice state changes:
$|k\rangle\otimes|\{n_{q}\}\rangle\to|k^{\prime}\rangle\otimes|\{n_{q}^{\prime}\}\rangle,\qquad n_{q}^{\prime}=n_{q}\pm 1.$
Energy conservation requires
$\varepsilon(k^{\prime})=\varepsilon(k)\pm\hbar\omega_{q}+\Delta E_{\rm CM}.$
A Golden-Rule rate has the form
$W_{k\to k^{\prime};q}=\frac{2\pi}{\hbar}|M_{k,k^{\prime};q}|^{2}\,\delta\!\left(\varepsilon(k^{\prime})-\varepsilon(k)\mp\hbar\omega_{q}-\Delta E_{\rm CM}\right),$
with matrix elements scaling as
$M_{k,k^{\prime};q}\sim\alpha\sqrt{\frac{\hbar}{2M\omega_{q}}}\,F_{k,k^{\prime};q}.$
Multi-phonon and hybrid CM–phonon processes
Higher-order multi-phonon processes and hybrid processes involving both CM recoil and internal phonon exchange relax the strict resonance condition (47) at the cost of reduced matrix elements.
Kinematic summary
  • Phonon-diagonal channel: $k\to-k$ with unchanged internal state is allowed by translational invariance and proceeds via CM recoil; it is strictly elastic only on resonance (47) and effectively elastic for large $N$.
  • Inelastic channels: internal phonon emission/absorption become important when the phonon spectrum permits energy matching.
  • Finite size: discrete internal spectra constrain inelastic channels and can enhance the relative importance of CM-mediated phonon-diagonal scattering.
Practical consequences
  • In wavepacket simulations, phonon-diagonal backscattering is identified by momentum reversal with unchanged internal state and a compensating shift of $P_{R}$ by $2\hbar k$.
  • Bulk and boundary reflections differ only in spatial location; both satisfy the same momentum conservation enforced by the CM zero mode.
  • Neglecting the CM zero mode (implicit clamping) qualitatively changes the available scattering channels.
Appendix C Whither Umklapp?
The traditional role of Umklapp for finite resistivity in an infinite, perfect crystal arises from freezing the background and projecting out its pseudomomentum. Once the background is treated as dynamical, with an explicit translational degree of freedom, momentum relaxation can occur through elastic pseudomomentum exchange between electron and lattice. Phonons then represent one possible basis for describing background dynamics rather than the exclusive channel for momentum transfer.
This perspective clarifies long-standing ambiguities surrounding Umklapp and explains why momentum relaxation is observed experimentally even in regimes where Umklapp processes are expected to be suppressed. It also underscores that the association of Umklapp with dissipation and entropy production is not intrinsic, but follows from additional assumptions about the background.
The expectation that Umklapp processes are necessary for finite resistivity in a perfect crystal has deep roots in transport theory. Peierls first distinguished between normal and Umklapp processes in his analysis of phonon thermal conductivity, showing that only the latter can relax total crystal momentum in a closed system. This reasoning was later extended to electrons by Bloch, Peierls, Ziman, and others, and appears explicitly in standard texts (e.g. Peierls, Quantum Theory of Solids; Ziman, Electrons and Phonons). If the Fermi surface lies well inside the Brillouin zone, or at low temperature where large-wavevector phonons are thermally suppressed, Umklapp processes are expected to be exponentially rare. Under these conditions conventional theory predicts extremely long momentum lifetimes in a clean crystal.
The present formulation revisits this conclusion by restoring the dynamical translational degree of freedom of the background. Momentum exchange between subsystems does not require phonon creation, and the strict identification of Umklapp with momentum relaxation relies on additional assumptions about background clamping or environmental coupling.
Umklapp processes are best understood as a classification of crystal-momentum conservation rather than as the fundamental cause of momentum relaxation. Their elevated role in conventional transport theory stems from treating the lattice background as inert. Recognizing the background as a dynamical object carrying its own pseudomomentum restores a more symmetric and physically complete description in which elastic momentum exchange is allowed and Umklapp is no longer indispensable.