### v3

# Elastic and Quasielastic Electron Scattering in Perfect Crystals

Abstract
Momentum relaxation of electrons in perfect crystals is conventionally associated with phonon emission or absorption and is therefore regarded as intrinsically inelastic. We show that this is not a fundamental requirement. Starting from the Fröhlich Hamiltonian, but making fewer assumptions than usual—retaining total mechanical momentum and the full density-density electron-lattice interaction—we find a substantially larger class of allowed processes. When total momentum conservation is enforced together with all pseudomomenta, the lattice center-of-mass provides a continuous momentum reservoir, and phonons are no longer required to account for the electron’s momentum change. The conventional phonon-mediated processes emerge as a restricted subset of a higher-dimensional manifold of transitions. Within this expanded space, elastic and quasielastic momentum-transfer channels appear already at first order, even in a perfect crystal. Momentum relaxation can therefore proceed rapidly while energy relaxation remains slow, consistent with a wide range of experimental observations including hot-electron injection, quantum oscillations, weak localization, and Coulomb drag in low-dimensional conductors. In particular, scattering in a perfect one-dimensional wire is not kinematically constrained once the lattice recoil degree of freedom is retained.
preprint: APS/123-QED
I Introduction
Electron–lattice interactions play a central role in condensed matter physics. They govern both momentum relaxation and energy exchange in crystalline solids, and therefore underlie electrical resistivity and the loss of electronic coherence.
In the modern transport theory of metals (Bloch–Peierls–Ziman–Mott–Ashcroft–Mermin), the electron–lattice problem is formulated by imposing periodic boundary conditions and expanding lattice displacements using internal normal modes. In this construction, going back to Born and von K’arm’an, the lattice center-of-mass (zero) mode is removed at the outset, leaving only internal vibrational degrees of freedom. When an electron applies force to a lattice atom, it can only respond with finite energy phonons. The zero energy recoil, without which Mössbauer scattering would falter, has been removed. In first order, momentum transfer is forced to reside in a single phonon, leading to the familiar selection rules $k’=k\pm q$ and $\varepsilon_{k’}=\varepsilon_{k}\pm\hbar\omega_{q}$. In a perfect crystal, a change in electron momentum thus requires phonon emission or absorption and is therefore inelastic. Although this is true within the formulation, it is an unnatural result, not followed in the laboratory.
When the lattice center-of-mass degree of freedom is retained, the total pseudomomentum is conserved and can be exchanged between the electron and the background. If the zero mode is removed, this exchange channel is absent. The phonon sector alone does not form a closed system for pseudomomentum, and the familiar selection rule
$\Delta K_{e}+\Delta K_{\rm ph}=Gm$
need not hold for individual processes. Indeed, it is violated once the density-density interaction is treated beyond the linear Fröhlich approximation. Pseudomomentum is therefore not a conserved quantity within the electron-phonon subspace alone. The traditional clean pseudomomentum accounting falters if even the next terms are kept in the the density-density expansion, e.g.
Purely elastic processes are a rarity in the classical world, but are commonplace in the quantum world. Consider a linear atom-diatom collision, i.e. three translational degrees of freedom, separating to a center mass of the diatom, the atom, and the internal virbatipnal coordinate: if not too violent, upon measurement the diatom is likely to be in the state it started in, even though it has definitely been disturbed. Again, Mössbauer emission comes to mind. In fact for pass-through experiments, like fast electrons, both Bragg and a good portion of diffuse scattering is known to be elastic, with the crystal recoiling as a whole. Somehow an electron residing inside a crystal has found itself treated a different way, but it should not be.
Experimentally, momentum and energy relaxation often occur on markedly different timescales. In clean metals over broad temperature ranges, electrical resistivity and magnetotransport measurements indicate rapid momentum randomization, while hot-electron experiments, Johnson-noise thermometry, and related nonequilibrium probes show that energy relaxation to the lattice can remain comparatively slow. The resulting hierarchy
$\tau_{p}\ll\tau_{E},$ (1)
where $\tau_{p}$ is the momentum-relaxation time and $\tau_{E}$ the energy-relaxation time, indicates that electronic momentum can be efficiently scrambled without substantial net energy transfer to internal lattice excitations.
The experimentally observed separation between $\tau_{p}$ and $\tau_{E}$ does not seem to comport with the “always inelastic” narrative, and this reality check motivates a careful examination of how momentum conservation is implemented in a perfectly periodic crystal.
At finite temperature, lattice vibrations introduce perturbations that allow electrons to scatter between Bloch states. This mechanism was first analyzed by Bloch, who obtained the characteristic low-temperature $T^{5}$ resistivity associated with electron-phonon scattering in metals [8]. These ideas were incorporated into the Bloch-Grüneisen-Boltzmann transport framework [8, 13], in which the electronic distribution function evolves according to a Boltzmann equation for Bloch electrons. Within this formulation, scattering is described as transitions between Bloch states accompanied by phonon absorption or emission, enforced by energy-conserving delta functions. In the standard Fröhlich Hamiltonian [10], the interaction is expressed in terms of phonon creation and annihilation operators, so that momentum transfer is associated, at lowest order, with changes in phonon occupation and is therefore intrinsically inelastic, subject to pseudomomentum and energy conservation [39, 4]. Elastic momentum relaxation is accordingly attributed to impurities or other forms of explicit disorder. The subsequent development of Migdal-Eliashberg theory further reinforced this picture, embedding the identification of momentum relaxation with inelastic electron-phonon scattering into the standard theoretical framework.
The situation has been different for external, “pass through” probes such as X-rays or neutrons or even fast electrons. In such circumstances, elastic and even diffuse elastic recoil of the entire crystal is routinely included through Debye-Waller theory. Here, we will see that there is no solid basis for the asymmetry in the treatment of an interior electron.
As we will see, bringing back the center of mass dramatically changes the narrative of internal scattering processes. Somehow that seems implausible, which is one reason perhaps that the inelastic narrative has persisted for so long.
We revisit this issue determined to be rigorous about momentum conservation in a fully isolated electron-crystal system with all applicable momenta retained. (The isolation can later be broken once a full understanding is reached of the isolated system.) When the electron-lattice interaction is written in its full density-density form, elastic and quasi-elastic momentum exchange channels appear already at first order, without any defects present other than the transient ones implied by the thermal deformation potential.
A few preliminary, qualitative remarks are in order. First, removing the center of mass makes electrons in crystals is akin to the venerable particle in a box model. There too there is no center of mass of the box, and no box recoil. Momentum is not conserved. In the Appendix we show what happens when the box is isolated, given mass, and allowed to recoil. Second, if the center of mass is absent, the only degree of freedom that could carry the recoil resulting from an elastic interior electron deflection is also absent. Phonons will be forced to take on that role and indeed by construction there can be no elastic events of this sort.
If the crystal total mechanical momentum is retained, and total momentum conservation is again enforceable, a transfer of momentum $\mathbf{q}$ to an electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background. Those degrees of freedom include the several lattice zero frequency recoil modes, and internal vibrational modes.
An instructive analog is provided by Mössbauer scattering, in which the dominant elastic line, perhaps 97% of the total probability, implies recoil of the entire lattice without internal phonon excitation.
We hasten to point out that the question here involves real physics, and not mere representation: are phonons created or destroyed when electrons deflect in a pure crystal, or not? That is an essential, physical difference, with profound experimental consequences.
In a finite-temperature many-body system, strictly elastic processes do not exist in a provable, absolute sense. Even the dominant “elastic” line in Mössbauer scattering exhibits slight thermal broadening. We therefore use the term elastic to denote processes in which phonon occupation does not change, and quasielastic to denote processes involving small energy exchanges that do not balance electron momentum transfer; that task will be left to the crystal center of mass recoil. In fact there are so many ways that momentum conservation can be accommodated that the textbook phonon carrying away exactly the right momentum is a measure zero event, in spite of its utility as a model. This will become clearer below.
II A minimal example: the electron on a free atomic ring
Figure 1: Left: Ring lattice with a localized phonon excitation. All angular coordinates are retained, including the 0 mode (angular momentum). Right: The tight-binding electronic wavefunction is added. The color wheel represents the amplitude and phase of the electronic amplitude $e^{ikx}$, winding here by $4\pi$ around the ring. Only total angular momentum $L_{\rm tot}=L+R\hbar k$. and energy are conserved quantities.
Before turning to the general many-body problem, it is useful to display the dramatic effect of including the zero mode in the simplest finite crystal that has exact momentum bookkeeping: a ring of $N$ identical atoms of mass $m$, constrained to move on a circle of radius $R$ and coupled by nearest-neighbor springs (figure 1, left). Completing the picture, an electron moves on the same ring by dynamic tight binding nearest-neighbor hopping (figure 1, right). Dynamic means the atoms move along the ring, unconstrained except for their neighbors, and carry their orbitals with them. An early example of this approach applied to graphene is found in ref. [28]. The importance of this model is its simplicity, yet it retains essential electron-phonon interactions. The lattice zero mode is present explicitly, total angular momentum is conserved, and the distinction between internal phonons and rigid-body recoil is completely transparent. All lattice degrees of freedom are retained, including the global rotational (zero) mode, which carries the total angular momentum of the lattice.
The standard conclusion for such a 1D model, like a perfect wire, is that only elastic backscattering is allowed, and even then under restricted circumstances, due to the simultaneous constraints of energy and momentum conservation when only the electron and a phonon are present. Efficient relaxation requires an additional degree of freedom. In conventional treatments[26, 18] this is introduced in the form of a third electronic excitation, such as a deep hole drawn from the Fermi sea to satisfy the combined energy and momentum constraints that cannot be met within a strictly two-body electron-phonon process. Multiphonon processes may also be invoked, but at higher order. An especially clean case is provided by Coulomb drag[24] in wire pairs, where exponential suppression of drag with temperature is again predicted, but instead power law dependence is seen in experiments. We will find something completely different, with manifolds of transitions possible obeying all conserved quantities. The third, “new” degree of freedom turns out to be the long lost center of mass.
In drag, current in one wire creates fluctuating electric fields that transfer momentum to the other wire through the Coulomb interaction. For that transfer to occur, both wires must be able to absorb the same small energy $\hbar\omega$ and momentum $q$. The drag rate is therefore governed by the overlap of the two wires’ low-energy dynamical response functions. If each wire has gapless excitations near the Fermi points, then as $T\to 0$ there remains a shrinking but nonzero window of allowed $\omega$ and $q$, typically of size set by $k_{B}T$. Integrating over that thermal window gives a power of $T$.
II.1 Hamiltonian and collective coordinates
Let $\phi$ denote the angular coordinate of the electron, and let $\theta_{\ell}$ denote the instantaneous angular position of atom $\ell=1,\dots,N$ on the ring. The atomic coordinates are dynamical variables and include thermal displacements.
For each instantaneous lattice configuration $\{\theta_{\ell}\}$, define a set of localized electronic orbitals
$\chi_{\ell}(\phi;\theta_{\ell})=\chi(\phi-\theta_{\ell}),$ (2)
where $\chi(\phi-\theta_{\ell})$ is an electron wavefunction centered on the nucleus at $\theta_{\ell}$. A general single-electron state may then be written as
$\Psi(\phi;\{\theta_{\ell}\})=\sum_{\ell=1}^{N}\psi_{\ell}\,\chi(\phi-\theta_{\ell}),$ (3)
where the coefficients $\psi_{\ell}$ give the electronic amplitudes on the lattice-following orbitals.
The full Hamiltonian acts both on the lattice coordinates $\{\theta_{\ell}\}$ and on the electronic amplitudes $\{\psi_{\ell}\}$:
$H=H_{\rm lat}(\{\theta_{\ell}\})+H_{\rm el}(\{\theta_{\ell}\}),$ (4)
Here $H_{\rm el}(\{\theta_{\ell}\})$ denotes the electronic Hamiltonian in the lattice-following orbital basis. It encodes the electron-lattice interaction. Its matrix elements depend on the instantaneous atomic positions and therefore include both the electronic energies and the electron-lattice interaction.
$H_{\rm lat}$ $=-\frac{\hbar^{2}}{2I_{\rm ion}}\sum_{\ell=1}^{N}\frac{\partial^{2}}{\partial\theta_{\ell}^{2}}+\frac{K}{2}\sum_{\ell=1}^{N}\left(\theta_{\ell+1}-\theta_{\ell}-\tfrac{2\pi}{N}\right)^{2},$ (5)
$\qquad\theta_{N+1}\equiv\theta_{1}+2\pi.$
In the nearest-neighbor tight-binding approximation, the electronic Hamiltonian acts on the coefficients $\psi_{\ell}$ as
$(H_{e}\psi)_{\ell}=-\,t_{\ell-1}\,\psi_{\ell-1}-\,t_{\ell}\,\psi_{\ell+1},\qquad\psi_{\ell+N}\equiv\psi_{\ell},$ (6)
where the hopping amplitudes $t_{\ell}$ are determined by the overlap of neighboring orbitals and therefore depend on the instantaneous bond lengths, i.e. on the differences $\theta_{\ell+1}-\theta_{\ell}$.
II.2 Lattice: zero mode and internal phonons
Let $\theta_{\ell}$ be the angular position of atom $\ell=1,\dots,N$, with periodic indexing $\theta_{N+1}\equiv\theta_{1}$. Introduce the center angle
$\Theta=\frac{1}{N}\sum_{\ell=1}^{N}\theta_{\ell}$ (7)
and internal coordinates
$\varphi_{\ell}=\theta_{\ell}-\Theta,\qquad\sum_{\ell=1}^{N}\varphi_{\ell}=0.$ (8)
The angular coordinates $\theta_{\ell}$ denote the instantaneous positions of the atoms and include thermal displacements. We therefore write
$\theta_{\ell}=\Theta+\frac{2\pi\ell}{N}+\varphi_{\ell},$ (9)
where $\Theta$ describes the rigid rotation of the lattice and $\varphi_{\ell}$ the internal (thermal) displacements, with $\sum_{\ell}\varphi_{\ell}=0$.
The lattice Hamiltonian is
$H_{\rm lat}=\sum_{\ell=1}^{N}\frac{L_{\ell}^{2}}{2mR^{2}}+\frac{\kappa R^{2}}{2}\sum_{\ell=1}^{N}(\varphi_{\ell+1}-\varphi_{\ell})^{2}.$ (10)
In these variables the Hamiltonian separates exactly as
$H_{\rm lat}=\frac{L^{2}}{2I_{\rm tot}}+H_{\rm ph},\qquad I_{\rm tot}=NmR^{2},$ (11)
where
$L=\sum_{\ell=1}^{N}L_{\ell}=-i\hbar\frac{\partial}{\partial\Theta}$ (12)
is the angular momentum of rigid rotation of the entire ring. The term $H_{\rm ph}$ contains the N1N-1 internal normal modes built from the coordinates $\varphi_{\ell}$.
Thus the $q=0$ mode is not discarded: it appears explicitly as the rotor term $L^{2}/2I_{\rm tot}$, while the remaining modes are the usual phonons.
In the lattice-following description, the electron is represented in a basis of localized orbitals $\chi(\phi-\theta_{\ell})$ centered on the instantaneous atomic positions. A general state may be written as
$\Psi(\phi;\{\theta_{\ell}\})=\sum_{\ell}\psi_{\ell}\,\chi(\phi-\theta_{\ell}),$ (13)
where $\psi_{\ell}$ are the amplitudes on the orbitals.
In this basis the electronic Hamiltonian acts on the amplitudes through nearest-neighbor matrix elements,
$(H_{\rm el}\psi)_{\ell}=\varepsilon_{\ell}(\{\theta\})\,\psi_{\ell}-t_{\ell-1}(\{\theta\})\,\psi_{\ell-1}-t_{\ell}(\{\theta\})\,\psi_{\ell+1},$ (14)
with periodic boundary conditions $\psi_{\ell+N}\equiv\psi_{\ell}$. The matrix elements depend on the instantaneous lattice configuration and therefore incorporate both the electronic structure and the electron-lattice coupling.
For small distortions, the bond length depends only on internal differences,
$d_{\ell}\simeq a+R(\varphi_{\ell+1}-\varphi_{\ell}),\qquad a=\frac{2\pi R}{N},$ (15)
so that, to linear order,
$t_{\ell}\simeq t_{0}-\alpha R(\varphi_{\ell+1}-\varphi_{\ell}).$ (16)
Thus the electronic Hamiltonian depends only on the internal coordinates $\{\varphi_{\ell}\}$: the rigid rotation $\Theta$ drops out.
II.3 Relative coordinate and Bloch-like structure
Because the Hamiltonian depends on $\phi$ and $\Theta$ only through the relative coordinate $\chi=\phi-\Theta$, the eigenstates may be chosen in Bloch-like form,
$\Psi_{k}(\phi,\Theta,\{\varphi_{\ell}\})=e^{ik(\phi-\Theta)}\sum_{\ell}u_{\ell k}(\{\varphi_{\ell}\})\,\chi(\phi-\theta_{\ell}),$ (17)
where the coefficients $u_{\ell k}$ describe the internal structure of the state in the lattice-following basis.
Here the instantaneous atomic positions are
$\theta_{\ell}=\Theta+\frac{2\pi\ell}{N}+\varphi_{\ell},$ (18)
so that the orbitals $\chi(\phi-\theta_{\ell})$ move with the lattice atoms11.
In this form the exponential factor $e^{ik(\phi-\Theta)}$ carries the crystal pseudomomentum, while the coefficients $u_{\ell k}$ encode the internal, lattice-following structure of the electronic state. In the absence of distortions, $u_{\ell k}\propto e^{ik\frac{2\pi\ell}{N}}$, and Eq. (17) reduces to the standard tight-binding Bloch wave.
The factor $e^{ik(\phi-\Theta)}$ in equation 17 is revealing and promising. The overall zero mode rotation angle $\Theta$ now takes its place among the dynamical variables. It is not benign, because it is not quite the angle belonging to the the total angular momentum, since that includes the electron. The key point is that the evolution of $\Theta$ is nontrivial, depending on recoil from changes in the angular momentum of the electron. It is clear from this form that if the electron $\phi$ has pseudomomentum, so must $\Theta$, and both can change, always equally and oppositely. This will be developed in the next section. The sum of the two pseudomomenta, call them $k_{\phi}=k$ and $k_{\Theta}=-k$, vanishes, and remains zero under changes of $k$, assuring the the conservation of total pseudomomentum.
Because zero mode rotation it is now included, a world of elastic and quasielastic processess will open up.
II.4 Total angular momentum balance
Rotational invariance of the full Hamiltonian implies conservation of the total angular momentum operator
$L_{\rm tot}=-i\hbar\frac{\partial}{\partial\phi}-i\hbar\frac{\partial}{\partial\Theta}.$ (19)
Acting on the Bloch-like state gives the one-line conservation rule
$L_{\rm tot}=L+R\hbar k,$ (20)
where $L$ is the angular momentum conjugate to $\Theta$.
Equation (20) implies that any change in the electronic pseudomomentum $k$ must be accompanied by a compensating change in $L$, corresponding to recoil of the ring as a whole. This is of course necessary to keep total angular momentum constant (lattice plus electron). Interestingly, there is no role yet for phonon momentum, and we can leave it this way if we want. As Peierls pointed out, phonon momentum can be assigned to the lattice as a whole[29], Chpts 1-2. Here the momentum appeared as belonging to the lattice, and indeed in must, since $L=\sum_{\ell}L_{\ell},$ which is a sum over all the motion of every atom. Any supposed phonon momentum would be double counting.
We have already remarked that the form $\exp[ik(\psi-\Theta)]$ already implies the $L$ (not $L_{tot}$) must be regarded as a pseudomomentum. This will br defended on symmetry grounds in the next section. Both the electron and its host have pseudomomentum, which is actually a very comfortable circumstance. We should indeed be uncomfortable if we suppose that a pseudomomentum and a mechanical momentum could combine.
II.5 Phonon-diagonal (elastic) backscattering
The simplest illustration of the dramatic effect of keeping the zero recoil mode is provided by phonon-diagonal backscattering. In the absence of lattice distortions, the electronic eigenstates are plane waves
$\left|k\right\rangle=\frac{1}{\sqrt{N}}\sum_{\ell}e^{ika\ell}c_{\ell}^{\dagger}\left|0\right\rangle,$ (21)
with even dispersion $\varepsilon(k)=\varepsilon(-k).$ For an initial state
$\left|\Psi_{i}\right\rangle=\left|k\right\rangle\left|\nu\right\rangle\left|L\right\rangle,$ (22)
one allowed elastic transition is
$\left|k\right\rangle\left|\nu\right\rangle\left|L\right\rangle\;\rightarrow\;\left|-k\right\rangle\left|\nu\right\rangle\left|L+2\hbar kR\right\rangle,$ (23)
in which the electron reverses its direction while the internal phonon state remains unchanged. The compensating angular momentum is absorbed by rigid-body recoil of the lattice.
The elastic backscattering channel
$\left|k\right\rangle\left|\nu\right\rangle\to\left|-k\right\rangle\left|\nu\right\rangle$
is kinematically forbidden in the conventional phonon-only representation unless Umklapp is invoked. In this process $\Delta K_{e}=-2k$ while $\Delta K_{\rm ph}=0$, so the standard selection rule cannot be satisfied by phonons alone. More fundamentally, the process has no carrier of the corresponding mechanical momentum once the lattice center-of-mass degree of freedom is removed. The absence of this channel is therefore not a dynamical result, but a consequence of having eliminated the recoil degree of freedom that would otherwise carry the compensating momentum.With the pseudoangularmomentum intact, it is kinematically robust, but to verify that this channel is dynamically allowed, we evaluate the corresponding matrix element of the electron-lattice interaction. In the density-density (bond-length dependent) form, the hopping amplitude depends on the relative displacements of neighboring sites, $t_{\ell}=t_{0}+\alpha(u_{\ell+1}-u_{\ell}).$ Because the interaction depends only on internal coordinate differences, it is independent of the global rotational coordinate, and therefore does not couple directly to $L$.
Nonvanishing elastic recoil matrix element (moving basis).
We evaluate the matrix element in an instantaneous lattice basis, in which localized electronic orbitals follow the atomic positions. Let
$R_{\ell}(\Theta,\{\xi\})=R\Theta+\ell a+\xi_{\ell},$ (24)
where $\Theta$ is the global rotational coordinate and $\{\xi_{\ell}\}$ are internal displacements. The electronic states are
$\left|k;\Theta,\{\xi\}\right\rangle=\frac{1}{\sqrt{N}}\sum_{\ell}e^{ik\ell a}\,c_{\ell}^{\dagger}(\Theta,\{\xi\})\left|0\right\rangle.$ (25)
For fixed configuration, the Hamiltonian is
$H_{e}=-\sum_{\ell}t_{\ell}(\{\xi\})\left(c_{\ell+1}^{\dagger}c_{\ell}+\mathrm{h.c.}\right),$ (26)
with
$t_{\ell}(\{\xi\})=t\!\left(a+\xi_{\ell+1}-\xi_{\ell}\right).$ (27)
The global coordinate $\Theta$ drops out, as $t_{\ell}$ depends only on internal differences.
For the elastic recoil channel
$\left|k,\nu,L\right\rangle\rightarrow\left|-k,\nu,L+2\hbar kR\right\rangle,$ (28)
the electronic matrix element at fixed configuration is
$\langle-k|H_{e}|k\rangle=-\frac{2\cos ka}{N}\sum_{\ell}t_{\ell}(\{\xi\})\,e^{2ika\ell}.$ (29)
For a uniform lattice this vanishes, but for a generic instantaneous configuration $t_{\ell}$ is nonuniform and the $2k$ component is nonzero. To leading order,
$t_{\ell}\simeq t_{0}+\alpha(\xi_{\ell+1}-\xi_{\ell}),$ (30)
so that
$\langle-k|H_{e}|k\rangle\propto\sum_{\ell}e^{2ika\ell}(\xi_{\ell+1}-\xi_{\ell}).$ (31)
This is a satisfying outcome: the backscattering amplitude is controlled by the $2k$ Fourier component of the instantaneous deformation field and is generically nonvanishing. The operator acts only on internal coordinates, while the required compensating change is carried by the global mode, $L\to L+2\hbar kR$. The process therefore remains elastic in the phonon sector..
The associated rotational energy shift
$\Delta E_{\rm rot}=\frac{(L+2\hbar kR)^{2}-L^{2}}{2I_{\rm tot}}$ (32)
is negligible for macroscopic $I_{\rm tot}=NmR^{2}$, so the process is effectively elastic.
This example shows explicitly that momentum transfer need not be tied to a change in the internal phonon state: it may instead be carried by the lattice zero mode.
It is worth emphasizing that the phonon-diagonal backscattering process in Eq. (23) is absent in standard treatments in which the lattice center-of-mass (zero mode) is eliminated. In that representation, momentum conservation is enforced solely through the internal phonon modes, implying $k^{\prime}=k\pm q$ and excluding the channel $\nu^{\prime}=\nu$ for $k\to-k$. The absence of this process therefore reflects the removal of the recoil degree of freedom rather than a dynamical constraint.
II.6 Elastic and quasi-elastic scattering
The general kinematics of electron-lattice scattering in the ring geometry follow directly from the conservation law Eq. (20). If the electron changes crystal momentum
$k\rightarrow k^{\prime},$ (33)
the lattice angular momentum must adjust according to
$L^{\prime}=L+\hbar R(k-k^{\prime}).$ (34)
Thus the exact momentum-conserving transitions take the form
$\left|k,\nu,L\right\rangle\;\rightarrow\;\left|k^{\prime},\nu^{\prime},L^{\prime}\right\rangle.$ (35)
The corresponding energy balance is
$\varepsilon(k)-\varepsilon(k^{\prime})=\sum_{q}\hbar\omega_{q}(n_{q}^{\prime}-n_{q})+\frac{(L^{\prime})^{2}-L^{2}}{2I_{\rm tot}},$ (36)
where the final term represents the recoil energy of the lattice. For a macroscopic ring this contribution is typically negligible, so that
$\varepsilon(k)-\varepsilon(k^{\prime})\simeq\sum_{q}\hbar\omega_{q}(n_{q}^{\prime}-n_{q}).$ (37)
Because the lattice angular momentum $L$ is retained as a dynamical degree of freedom, each electronic transition $k\to k^{\prime}$ is compatible with a broad family of vibrational responses:
  1. 1.
    Elastic recoil
    $\left|k,\nu,L\right\rangle\rightarrow\left|k^{\prime},\nu,L+\hbar R(k-k^{\prime})\right\rangle,$ (38)
    in which the phonon configuration is unchanged and the rigid rotation of the lattice carries the momentum transfer.
  2. 2.
    Quasi-elastic single-phonon processes
    $\left|k,\nu,L\right\rangle\rightarrow\left|k^{\prime},\nu\pm 1_{q},L+\hbar R(k-k^{\prime})\right\rangle,$ (39)
    in which the zero mode absorbs the angular momentum while a phonon adjusts the energy balance.
  3. 3.
    Multiphonon processes
    $\left|k,\nu,L\right\rangle\rightarrow\left|k^{\prime},\nu+\{\Delta n_{q}\},L+\hbar R(k-k^{\prime})\right\rangle.$ (40)
  4. 4.
    Hybrid recoil-phonon processes in which the momentum transfer is shared between the rigid lattice rotation and the internal phonon manifold.
Restoring the zero mode therefore enlarges the accessible final states from $\left|k^{\prime},\nu^{\prime}\right\rangle$ to $\left|k^{\prime},\nu^{\prime},L^{\prime}\right\rangle$. The additional angular momentum degree of freedom opens a wide manifold of momentum-conserving elastic and quasi-elastic scattering channels. The interaction matrix elements
$\langle k^{\prime},\nu^{\prime},L^{\prime}|H|k,\nu,L\rangle$ (41)
are therefore generically nonzero across this enlarged space of states.
II.7 Consequences for one-dimensional transport and ring geometries
In strictly one-dimensional electronic systems, two-particle collisions are strongly constrained by the simultaneous conservation of energy and momentum,
$k_{1}+k_{2}=k_{1}^{\prime}+k_{2}^{\prime},\qquad\varepsilon(k_{1})+\varepsilon(k_{2})=\varepsilon(k_{1}^{\prime})+\varepsilon(k_{2}^{\prime}).$ (42)
For a monotone dispersion these relations admit only trivial rearrangements, so that two-body processes cannot efficiently relax a near-Fermi distribution. In particular, within conventional electron-phonon treatments, even backscattering is highly constrained, since momentum transfer must be accompanied by phonon creation or annihilation, and the combined conservation laws severely restrict the available phase space.
Standard approaches therefore invoke an additional dynamical participant. This is typically taken to be a third electronic excitation,
$k_{1}+k_{2}+k_{3}=k_{1}^{\prime}+k_{2}^{\prime}+k_{3}^{\prime},$ (43)
with $k_{3}$ a deep hole drawn from the Fermi sea [26, 18, 24]. Because such excitations are thermally rare, the resulting relaxation rates are predicted to be exponentially suppressed at low temperature. Multiphonon processes can also relax the constraints, but only at higher order.
A particularly clear manifestation of this reasoning appears in Coulomb drag between parallel quantum wires, where theory predicts exponential suppression of drag at low temperature, in contrast with the power-law behavior observed experimentally.
Once the lattice center-of-mass mode is retained, however, an additional dynamical participant is always present. The conserved momentum becomes
$P_{\rm tot}=P_{\rm CM}+\sum_{i}\hbar k_{i},$ (44)
or in the ring geometry,
$L_{\rm tot}=L+\sum_{i}\hbar k_{i}R.$ (45)
An electronic momentum change $k\rightarrow k^{\prime}$ can then be balanced by recoil of the lattice,
$\Delta P_{\rm CM}=-\hbar(k^{\prime}-k),$ (46)
with associated energy
$\Delta E_{\rm CM}=\frac{(\Delta P_{\rm CM})^{2}}{2M_{\rm tot}},$ (47)
which is negligible for macroscopic $M_{\rm tot}$.
The additional degree of freedom required by one-dimensional kinematics is therefore always available and need not be supplied by a thermally activated deep electronic excitation. The “third particle” is the collective zero mode of the lattice.
As a result, large manifolds of elastic and quasi-elastic transitions exist that satisfy all conservation laws. Energy redistribution occurs within the phonon manifold rather than through rare electronic excitations, leading naturally to power-law, rather than exponentially suppressed, low-temperature behavior.
Such power-law relaxation is widely observed in clean quantum wires and related one-dimensional conductors [6, 1]. In this picture the lattice zero mode provides the natural third participant in the scattering process, enabling momentum relaxation through a broad class of elastic and quasi-elastic channels.
The same kinematics appears in the ring geometry familiar from the theory of persistent currents. Once the lattice center-of-mass motion is retained, the electronic phase winding can transfer angular momentum to the lattice through recoil, with the zero mode acting as the collective degree of freedom that accommodates the momentum balance.
Finally, we emphasize that the appearance of this additional degree of freedom is not an assumption but a consequence of enforcing the exact symmetries of the underlying Hamiltonian. No new interaction has been introduced, and no approximation has been invoked beyond those already implicit in the microscopic model. Retaining the lattice center-of-mass simply restores the full momentum conservation law of the isolated system.
Any objection that the zero mode should not act as the required third participant must therefore confront the fact that, once this conservation law is imposed, a large manifold of elastic and quasi-elastic scattering channels follows directly. These channels satisfy all conservation laws and arise without fine tuning or higher-order processes. The resulting kinematics is thus not a hypothesis but an unavoidable consequence of the complete Hamiltonian description.
III Foreground, Background, and Pseudomomentum
III.1 Defining Pseudomomentum
It is important that we clearly distinguish component parts We distinguish foreground and background degrees of freedom. The foreground is the object of interest, here a single electron, or it could be a collection of them, or a defect, or a lattice phonoon. The background consists of all remaining degrees of freedom, including the lattice atoms and any other electrons not treated explicitly. This distinction is essential for identifying the relevant symmetries and conserved quantities.
In a crystalline environment, the electronic Hamiltonian is invariant under discrete translations of the electron by lattice vectors, with the background held fixed. The associated conserved quantity is pseudomomentum (crystal momentum); for an electron in a periodic potential this is the familiar Bloch momentum $\hbar\mathbf{k}$.
The same discrete translational symmetry applies to the background. Translating the entire lattice while holding the electron fixed leaves the Hamiltonian invariant, and therefore assigns a pseudomomentum to the background as well. (see Fig. 2). If the electron (foreground A) coordinate is held fixed and the entire background B is translated by a lattice vector, the Hamiltonian is again unchanged. A consistent application of translational symmetry therefore assigns pseudomomentum not only to the electron but also to the background, even though only their sum corresponds to a mechanical momentum.
One might object that the problem can always be formulated in the rest frame of the background, with only the electron moving. However, this choice of frame obscures the conservation of total momentum of the combined system and suppresses the dynamical role of background recoil, which is essential for understanding local scattering events in the interior of a perfect crystal.
Figure 2: Sliding Blocks. Untethered, two blocks attract one another and slide frictionlessly along their common long axis. At the top, both blocks carry conserved mechanical momenta, and clearly $P_{\rm tot}=p_{A}+p_{B}$. At the bottom, the same relation holds, but $p_{A}$ and $p_{B}$ are now not independently conserved; if the boxes are sliding over one another,their sum nevertheless equals the conserved mechanical momentum. The two blocks serve as surrogates for a foreground electron (A) in a perfect background crystal (B), illustrating that both the foreground and the background acquire pseudomomentum in each other’s presence.
Let $\hat{U}(\mathbf{h})$ translate all coordinates, foreground and background, by an arbitrary vector $\mathbf{h}\in\mathbb{R}^{3}$. For an isolated electron+background system,
[…]
so the corresponding symmetry is the conserved total mechanical momentum, denoted by $\hat{\mathbf{Q}}$.
In the foreground/background formulation $\hat{\mathbf{Q}}$ is obtained as the sum of two subsystem pseudomomenta. Because $\hat{U}(\mathbf{h})$ factorizes into a translation of the foreground coordinate and an equal translation of the background configuration,
$\hat{U}(\mathbf{h})=\hat{U}_{\mathrm{f}}(\mathbf{h})\,\hat{U}_{\mathrm{b}}(\mathbf{h}),$ (49)
we have
$\hat{\mathbf{Q}}=\hat{\mathbf{P}}+\hat{\mathbf{p}}$ (50)
Here $\hat{\mathbf{p}}$ is the pseudomomentum carried by the foreground electron (the generator of its translation), and $\hat{\mathbf{P}}$ is the pseudomomentum carried by the background. The total background pseudomomentum decomposes into
$\hat{\mathbf{P}}=\hat{\mathbf{P}}_{\bf R}+\hat{\mathbf{P}}_{\rm ph}$ (51)
the center-of-mass (zero-mode) contribution and the internal phonon contribution. (the generator of translating the background as a whole). Their sum $\hat{\mathbf{Q}}=\hat{\mathbf{P}}+\hat{\mathbf{p}}$ is the conserved mechanical momentum of the isolated system. For convenience, we can set it to 0; it never changes and we can now ignore it.
Even under the usual rule stating that defects, such as a crystal edge or impurity, scatter electrons elastically, there is a subtlety within the prior framework: the electron changes its pseudomomentum, but as is conventionally understood, the lattice cannot respond or recoil in kind. A pseudomomentum cannot be combined with a mechanical momentum. We have now seen, however, that both foreground and background have pseudomomentum, and can exchange it freely. Total momentum is conserved for example in the very real deflection of an electron at an edge by equal and opposite pseudomomentum changes.
Historically, there has been a trend to treat electron scattering in a crystal as if it were Compton scattering in free space. There is no ether, no center of mass of a substrate. This is elegant, but also it misses something important. Here, we restore for a periodic lattice a full set of rules that apply in free space, as one would wish. There is a total psuedomomentum conserved even if an electron encounters a defect. However, unlike in free space, in crystals there is an aether, so to speak, a silent partner: the lattice. It is more than its phonons, it has a center of mass.
Figure 3 reflects the revised understanding and diagram including the participation of the lattice. This collision takes place in an “ether.”
Figure 3: (top) Traditional diagram of electron deflection by phonon emission or absorption. (bottom) Momentum transfer in a perfect crystal when the lattice zero mode is retained. The collision occurs on a dynamical lattice: quasielastic phonons may be produced, while the change in electron momentum is balanced jointly by the phonon momentum and the center-of-mass recoil of the lattice. Note: When momentum $\Delta p=\hbar q$ is transferred to the lattice center of mass, the associated de Broglie wavelength is $\lambda=2\pi/q$, identical to that of the scattering process. The large lattice mass suppresses the recoil energy but not the wavelength or the momentum transfer.
The total momentum $\hat{\mathbf{Q}}$ is conserved, but the foreground and background momenta, $\hat{\mathbf{p}}$ and $\hat{\mathbf{P}}=\hat{\mathbf{P}}_{R}+\hat{\mathbf{P}}_{\rm int}$, undergo equal and opposite changes, in exact analogy with collisions in free space. In particular, any change in $\hat{\mathbf{p}}$ necessarily induces a compensating change in the global (zero-mode) contribution $\hat{\mathbf{P}}_{R}$, while changes in the internal degrees of freedom are optional This exchange does not require the creation or annihilation of internal lattice excitations and occurs immediately at the collision event (see Fig. 4). The very real recoil of the atom labeled in red in Fig. 4 may very well prove to have excited no vibrational quanta when measured in a number state basis. This is just the elementary fact that a slightly displaced ground state of a harmonic oscillator is still mostly in the ground state, if measured.
In what follows, we formalize a symmetric foreground-background structure, derive the elastic scattering amplitude within first-order perturbation theory, and compute the elastic fraction explicitly. We show that this fraction is close to unity for copper at room temperature. We also clarify how the present formulation extends the Fröhlich framework without contradicting its successful predictions, and discuss the implications for the interpretation of transport phenomena in crystalline solids.
Figure 4: The total momentum (black) remains constant as the foreground (electron, blue) and background (lattice, red) momenta jump abruptly at the collision time $t^{*}$. There is no delay in the total lattice momentum change, even though only a single atom initially carries the momentum. While this is elementary, we emphasize it because there is a common tendency to associate momentum flow directly with energy flow in the medium. Indeed there is momentum flow among particles, but the total is fixed immediately. The total appears in the wavefunctions. As with Mössbauer scattering, the red atom has gained momentum and with it the whole lattice, but when measured, the lattice may still be in the same vibrational state as before the collision (elastic scattering).
III.2 Finite-window symmetry
In any physically realizable crystal the interaction is periodic only over a finite spatial extent $L$. Strict translational invariance is therefore absent at the boundaries, and crystal momentum is not an exact conserved quantity for the full finite system. This situation is directly analogous to the textbook particle in a box: the exact eigenstates are standing waves that do not carry definite momentum, yet traveling wave packets with well-defined average momentum can be constructed as superpositions of these eigenstates. Over times short compared with the traversal time to the boundary, such wave packets propagate as if momentum were conserved.
The same reasoning applies to a large but finite crystal. Although global translational symmetry is broken by the boundaries, the Hamiltonian is locally periodic in the bulk. Wave packets localized far from the boundaries evolve, for long but finite times, according to the translationally invariant bulk Hamiltonian. In this regime one may assign a well-defined crystal pseudomomentum to the packet, up to corrections that vanish as $L\to\infty$.
This type of approximate symmetry is standard in discussions of the thermodynamic limit, bulk transport theory, and wave-packet dynamics in finite systems. For clarity, we refer to it here as finite-window symmetry: translational invariance holds within a spatial and temporal window set by the system size. No new symmetry is implied; rather, this terminology emphasizes that the use of bulk momentum quantum numbers in a finite crystal is an approximation controlled by the ratio of the observation time and length scales to the system size.
In a macroscopic crystal this approximation is exceedingly accurate. Localized wave packets (“corpuscles”) can be constructed that propagate nearly as Bloch waves with well-defined pseudomomentum and experience long travel times before interacting with boundaries. This construction permits the use of the notation and results associated with Born–von Kármán boundary conditions as a calculational device, without assuming literal periodicity of the finite sample.
In a finite crystal, the exact normal modes are standing waves. As exact eigenmodes they have zero net mechanical momentum. Any “phonon carrying momentum $\hbar q$” is then not an eigenstate; it’s a wavepacket (a superposition of standing modes) that only behaves like a traveling object transiently until it reflects, mode-mixes, etc.
IV Lattice degrees of freedom and zero modes
We have already encountered a zero mode in the minimal model of an electron and an atom chain. section II. The mode there is the angular momentum.
Now we venture out more generally, and find the same large phase space becomes available for elastic and quasielastic scattering processes, a phase space that is simply absent if the center of mass is absent in the Hamiltonian.
We consider a finite crystal consisting of atoms with coordinates $\hat{\mathbf{R}}_{j}$ and masses $m_{j}$, with a total lattice mass $M=\sum_{j}m_{j}$, minus the foreground electron. We define the center-of-mass coordinate of the lattice background as
$\hat{\mathbf{R}}\equiv\frac{1}{M}\sum_{j}m_{j}\,\hat{\mathbf{R}}_{j}.$ (52)
It is convenient to decompose the atomic coordinates as
$\hat{\mathbf{R}}_{j}$ $=\hat{\mathbf{R}}+\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j},$ (53)
$\sum_{j}m_{j}\bm{\xi}^{(0)}_{j}$ $=\mathbf{0},\qquad\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$
where $\hat{\mathbf{R}}$ is the exact lattice center-of-mass coordinate, $\bm{\xi}^{(0)}_{j}$ denote fixed reference positions measured relative to the center of mass, and $\hat{\mathbf{u}}_{j}$ represent internal, dynamical lattice displacements. The mass-weighted constraints ensure that all internal degrees of freedom carry zero net momentum, so that $\hat{\mathbf{R}}$ describes the unique translational zero mode.
The classical and quantum theory of lattice vibrations explicitly separates rigid translation from internal vibrations. Born and Huang showed that the equations of motion of a finite crystal admit three zero-frequency normal modes corresponding to rigid translation of the crystal as a whole, while the internal normal modes satisfy the mass-weighted constraint $\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$ so that internal vibrations carry no net mechanical momentum [9]. All mechanical momentum of the crystal resides in the translational zero mode.
V Electron-Lattice interaction
V.1 Foreground and background Hamiltonians
. We write the total Hamiltonian as
$H=H_{\rm fg}+H_{\rm bg}+H_{\rm int},$ (54)
where $H_{\rm fg}$ describes the foreground electron, $H_{\rm bg}$ the background degrees of freedom, and $H_{\rm int}$ their coupling.
For a single foreground electron we take
$H_{\rm fg}=\frac{\hat{\mathbf{p}}^{2}}{2m},$ (55)
where $\hat{\mathbf{p}}$ denotes the generator of translations of the foreground coordinate $\hat{\mathbf{r}}$. The use of a bare kinetic-energy operator reflects the fact that the electron has not been pre-diagonalized into Bloch states. All effects of the periodic crystal potential, including band structure, pseudomomentum exchange, and scattering, are generated through the interaction Hamiltonian $H_{\rm int}$. This is the same starting point adopted by Fröhlich [10].
The remaining electrons of the metal are treated as part of the background. Together with the ions they establish the effective lattice potential, provide screening, and ensure mechanical stability. Their degrees of freedom are therefore implicit in $H_{\rm bg}$.
V.2 Density operator and exact factorization
The lattice density operator admits an exact factorization into a center-of-mass translation and an internal part:
$\hat{\rho}_{\mathbf{q}}$ $=\sum_{j}e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}_{j}}=e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{\mathbf{q}},$ (56)
$\hat{\rho}^{\rm int}_{\mathbf{q}}$ $=\sum_{j}e^{-i\mathbf{q}\cdot(\bm{\xi}^{(0)}_{j}+\hat{\mathbf{u}}_{j})}.$
Here $\hat{\mathbf{R}}$ is the lattice center-of-mass operator, $\bm{\xi}^{(0)}_{j}$ are equilibrium positions, and $\hat{\mathbf{u}}_{j}$ are internal displacements. Equation (56) is an exact operator identity, independent of any choice of basis (plane waves, Bloch waves, or standing waves).
This identity expresses a purely kinematic fact: a rigid translation of the lattice multiplies every density Fourier component by the phase $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, while all internal lattice dynamics are contained entirely in $\hat{\rho}^{\rm int}_{\mathbf{q}}$.
V.3 Density–density interaction and pseudomomentum conservation
The electron–lattice interaction may be written in density–density form as
$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,\rho_{\rm fg}(\mathbf{q})\,\rho_{\rm bg}(-\mathbf{q}),$ (57)
following the formalism of Van Hove [34]. No assumption about Bloch-wave eigenstates is required; $\mathbf{q}$ labels Fourier components of the interaction kernel. This equation is the essence of the term: density-density interaction.
For a tagged foreground electron, we write the aperiodic part of the Bloch waves as
$\rho_{\rm fg}(\mathbf{q})=e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}$. (58)
Substituting into Eq. (57) gives
$\boxed{H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}}$ (59)
Equation (59) shows that the interaction depends only on the relative coordinate $\hat{\bm{\rho}}=\hat{\mathbf{r}}-\hat{\mathbf{R}}$. The symmetry of the phase factor $e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}$ is the central structural feature: pseudomomentum $\mathbf{q}$ is exchanged between foreground and background with change‘s always equal and opposite.
When the electron momentum changes from $k$ to $k^{\prime}$, the lattice background absorbs the compensating momentum
$\Delta P_{\rm bg}=-\hbar(k^{\prime}-k).$ (60)
This momentum is carried entirely by the center-of-mass (zero mode) of the background. Internal lattice excitations describe only relative motion and carry no net mechanical momentum.
The associated energy change separates into internal and recoil contributions,
$\Delta E_{\rm lat}=\Delta E_{\rm int}+\frac{\Delta P_{R}^{\,2}}{2M_{\rm tot}},$ (61)
where $\Delta P_{R}=-\hbar(k^{\prime}-k)$ is the center-of-mass recoil momentum. For macroscopic $M_{\rm tot}$, the recoil energy is negligible, so large momentum transfer can occur with little or no change in internal lattice energy.
Translational invariance implies that the background degrees of freedom contain a rigid (center-of-mass) coordinate $\mathbf{R}$ with conjugate momentum
$\hat{\mathbf{P}}_{\bf R}=-\,i\hbar\frac{\partial}{\partial\mathbf{R}}.$
The remaining background variables describe internal excitations (phonons or distortions) defined relative to the center of mass. Accordingly, the background Hilbert space factorizes as
$\mathcal{H}_{\rm bg}=\mathcal{H}_{R}\otimes\mathcal{H}_{\rm int}.$
All mechanical momentum of the background resides in the zero mode,
$\hat{\mathbf{P}}=\hat{\mathbf{P}}_{\bf R},$ (62)
while the internal degrees of freedom carry no net mechanical momentum.
The interaction operator
$e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}$
transfers momentum $\hbar\mathbf{q}$ from the electron to the background as a whole. This momentum is carried entirely by the center-of-mass degree of freedom. Lattice recoil is therefore not an additional dynamical channel but a direct consequence of translational symmetry.
A nonvanishing matrix element of $H_{\rm int}$ requires
$\Delta\mathbf{p}=+\hbar\mathbf{q},\qquad\Delta\mathbf{P}=-\hbar\mathbf{q},$ (63)
so that the total mechanical momentum $\hat{\mathbf{Q}}=\hat{\mathbf{p}}+\hat{\mathbf{P}}$ is conserved exactly.
Crucially, this momentum exchange is independent of phonon occupation. Elastic matrix elements of $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ are generically nonzero, so the lattice can absorb finite momentum through center-of-mass recoil without any change in its internal state.
VI Elastic internal scattering and the Debye-Waller fraction
We now show that, even within a strictly perturbative treatment based on Bloch electrons and phonon number states, robust elastic scattering channels exist already at first order in the electron–lattice coupling. These channels become explicit once the full lattice density operator is retained and the translational zero mode is treated as a dynamical degree of freedom. The resulting processes are elastic with respect to internal lattice excitations, despite involving momentum exchange with the lattice as a whole.
VI.1 Emergence of the Bloch potential
Separating the background density operator into its equilibrium expectation value (taken in the lattice center-of-mass frame) and fluctuations,
$\rho_{\rm bg}(-\mathbf{q})=\langle\rho_{\rm bg}(-\mathbf{q})\rangle+\delta\rho_{\rm bg}(-\mathbf{q}),$ (64)
the static component produces an effective periodic potential
$U_{\rm Bloch}(\hat{\mathbf{r}})=\sum_{\mathbf{G}}V(\mathbf{G})\,e^{i\mathbf{G}\cdot\hat{\mathbf{r}}}\,\langle\rho_{\rm bg}(-\mathbf{G})\rangle,$ (65)
where $\mathbf{G}$ are reciprocal lattice vectors.
The corresponding foreground Hamiltonian becomes
[…]
which is the usual Bloch Hamiltonian. In this way, the periodic potential is not imposed a priori, but emerges as the mean-field component of the translationally invariant density-density interaction in the background center-of-mass frame.
The remaining part of $H_{\rm int}$ describes coupling to fluctuations of the background density. Crucially, it retains the explicit center-of-mass translation factor, ensuring that momentum transfer is accounted for by recoil of the background as a whole, rather than being assigned solely to internal phonon modes.
VI.2 Initial and final states
For a perturbative treatment, we consider product states
$\Psi_{i}=\Psi_{n\mathbf{k}}(\mathbf{r})\,\Psi^{\rm bg}_{i},\qquad\Psi_{f}=\Psi_{n^{\prime}\mathbf{k}^{\prime}}(\mathbf{r})\,\Psi^{\rm bg}_{f},$ (67)
where $\Psi_{n\mathbf{k}}$ are extended-zone Bloch eigenstates of the static lattice Hamiltonian and $\Psi^{\rm bg}_{i,f}$ are arbitrary background states in the full lattice Hilbert space (not restricted to phonon number eigenstates).
VI.3 Interaction and matrix element
From Eq. (59), the interaction is
$H_{\rm int}=\sum_{\mathbf{q}}V(\mathbf{q})\,e^{i\mathbf{q}\cdot(\hat{\mathbf{r}}-\hat{\mathbf{R}})}\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (68)
where $\hat{\mathbf{R}}$ is the lattice center-of-mass operator and $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ acts on internal background degrees of freedom.
Between product states the interaction matrix element is
$\mathcal{M}_{fi}$ $=\big\langle n^{\prime}\mathbf{k}^{\prime},\Psi_{f}^{\rm bg}\big|H_{\rm int}\big|n\mathbf{k},\Psi_{i}^{\rm bg}\big\rangle$
$=V(\mathbf{q})\,\langle n^{\prime}\mathbf{k}^{\prime}|e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}|n\mathbf{k}\rangle\,\langle\Psi_{f}^{\rm bg}|e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\hat{\rho}^{\rm int}_{-\mathbf{q}}|\Psi_{i}^{\rm bg}\rangle.$ (69)
VI.4 Total pseudomomentum conservation
Translational invariance of the full electron-lattice system implies exact conservation of total pseudomomentum,
$\mathbf{k}^{\prime}-\mathbf{k}=-\big(\mathbf{K}^{\rm bg}_{f}-\mathbf{K}^{\rm bg}_{i}\big),$ (70)
where $\mathbf{K}^{\rm bg}_{i,f}$ are eigenvalues of the total background pseudomomentum operator, including the zero (center-of-mass) mode.
Equation (70) is the only kinematic constraint imposed by translational symmetry. No additional reciprocal lattice vector is required when electronic states are treated in the extended-zone scheme.
In mechanical form this reads
$\Delta\mathbf{p}=-\Delta\mathbf{P},$ (71)
so that the total mechanical momentum $\hat{\mathbf{Q}}=\hat{\mathbf{p}}+\hat{\mathbf{P}}$ is conserved exactly.
VII Partition of pseudomomentum transfer
We now examine how the transferred pseudomomentum $\hbar(\mathbf{k}^{\prime}-\mathbf{k})$ is stored within the background.
The background matrix element contains two distinct structures,
$e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}\qquad\text{(center-of-mass recoil)},$ (72)
and
$\hat{\rho}^{\rm int}_{-\mathbf{q}}\qquad\text{(internal lattice operator)}.$ (73)
The conservation law (70) constrains only the total background pseudomomentum change $\Delta\mathbf{K}^{\rm bg}.$ It places no requirement on how that momentum is distributed between internal modes and the center-of-mass degree of freedom.
VII.1 Conventional phonon bookkeeping
In the clamped-lattice approximation one neglects the recoil operator and expands background states in phonon number eigenstates. In that restricted description one identifies
$\Delta\mathbf{K}^{\rm bg}=\sum_{\mathbf{q}}\Delta n_{\mathbf{q}}\,\mathbf{q},$ (74)
so that electronic momentum change is compensated entirely by phonon creation or annihilation.
Energy conservation then requires
$\epsilon_{n^{\prime}\mathbf{k}^{\prime}}=\epsilon_{n\mathbf{k}}\pm\hbar\omega_{\mathbf{q}\lambda}.$ (75)
This phonon-counting condition is therefore a consequence of the harmonic number-state basis together with suppression of the center-of-mass degree of freedom. It restricts the dynamics to a low-codimension subset of the full many-body phase space, in which momentum and energy must be matched by discrete phonon quanta.
VII.2 Full Hilbert-space description
In the complete electron-lattice Hilbert space, Eq. (70) remains the only exact constraint. The compensating background momentum is carried entirely by the center-of-mass (zero mode) of the lattice. Internal degrees of freedom describe only relative motion and carry no net mechanical momentum.
Thus, the association of background momentum transfer with specific phonon occupation changes is not fundamental, but a consequence of discarding the zero mode. Phonons may accompany a scattering event, but they do not provide the required momentum balance. The zero mode is the unique carrier of mechanical momentum.
In the clamped-lattice formulation, an electronic transition $k\to k^{\prime}$ is restricted to those phonon modes whose wavevector and frequency simultaneously satisfy both momentum and energy matching conditions. In low dimensions or at low temperature, such solutions are sparse and may be absent altogether except through Umklapp. When it is said that Umklapp momentum is “absorbed by the lattice,” the role of the zero mode is implicitly reintroduced for the special case of a reciprocal lattice vector $G$, even though the lattice can in fact absorb arbitrary momentum.
By contrast, when the translational zero mode is retained, momentum conservation is enforced by recoil of the background as a whole, while energy conservation is governed independently by the internal spectrum. The internal lattice states form an exponentially dense many-body manifold, and all configurations within an energy window $|E_{f}-E_{i}|\lesssim\Gamma$ contribute. The number of quasi-elastic channels therefore scales with the many-body density of states and becomes macroscopic in the thermodynamic limit. This vast phase space is absent when the zero mode is omitted.
VII.3 Recoil-dominated channel
When the zero recoil mode is retained, the required pseudomomentum transfer is taken up entirely by the background center-of-mass degree of freedom:
$\langle\mathbf{K}^{\prime}|e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}|\mathbf{K}\rangle=\delta_{\mathbf{K}^{\prime},\,\mathbf{K}-\mathbf{q}},$ (76)
so that total pseudomomentum conservation $\mathbf{k}^{\prime}-\mathbf{k}=-(\mathbf{K}^{\rm bg}_{f}-\mathbf{K}^{\rm bg}_{i})$ is satisfied without reference to phonon occupation changes.
The associated recoil energy is
$\Delta E_{\rm recoil}=\frac{\hbar^{2}}{2M_{\rm tot}}\left[(\mathbf{K}-\mathbf{q})^{2}-\mathbf{K}^{2}\right]=\frac{(\hbar q)^{2}}{2M_{\rm tot}}-\frac{\hbar^{2}}{M_{\rm tot}}\mathbf{K}\!\cdot\!\mathbf{q},$ (77)
which scales as $1/M_{\rm tot}$ and therefore vanishes in the thermodynamic limit. To leading order in $1/N$, the scattering is thus elastic with respect to internal lattice excitations.
The internal factor
$\langle\Psi_{f}^{\rm bg}|\hat{\rho}^{\rm int}_{-\mathbf{q}}|\Psi_{i}^{\rm bg}\rangle$ (78)
remains finite even when no phonon quanta are created or annihilated. For a thermally fluctuating background its diagonal contribution reduces, upon averaging, to the Debye-Waller factor $e^{-W(\mathbf{q},T)}.$
VII.4 Quasielastic and superelastic processes
More generally, $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ contains all orders in internal lattice operators. Transitions therefore exist in which the center-of-mass absorbs the entire pseudomomentum transfer, while the internal modes undergo small changes that adjust the energy balance. Such processes are quasielastic:
$\epsilon_{n^{\prime}\mathbf{k}^{\prime}}\approx\epsilon_{n\mathbf{k}},\qquad|\Delta E_{\rm int}|\ll\epsilon_{n\mathbf{k}},$ (79)
with the exact kinematic constraint always enforced by center-of-mass recoil.
At finite temperature, superelastic channels occur when pre-existing internal excitations are partially annihilated, transferring energy to the electron while the net pseudomomentum remains balanced by the center-of-mass degree of freedom.
VII.5 Elastic and quasi-elastic processes beyond the Fröhlich term
Consider the electron-lattice interaction in its full density-density form
$H_{\rm int}=\sum_{q}V(q)\,\rho_{e}(-q)\,\rho_{\rm bg}(q),\qquad\rho_{e}(-q)=\sum_{k}c^{\dagger}_{k+q}c_{k}.$ (80)
The background density may be written as
$\rho_{\rm bg}(q)$ $=\sum_{j}e^{-iq(R_{j}+u_{j})}$ (81)
$=\sum_{j}e^{-iqR_{j}}\left(1-iq\,u_{j}-\frac{q^{2}}{2}u_{j}^{2}+\cdots\right).$
where $R_{j}$ are equilibrium positions and $u_{j}$ displacements.
The zeroth-order term gives
$H^{(0)}=\sum_{k,q}V(q)\,c^{\dagger}_{k+q}c_{k}\sum_{j}e^{-iqR_{j}},$ (82)
which contains no phonon operators and is therefore purely elastic. In a perfect periodic crystal this reduces to reciprocal lattice scattering ($q=G$), i.e. Bragg processes. It is responsible for the static band structure. As Kohn noted[20], there is plenty of elastic deflection going into the construction of bands, which gets masked in making Bloch waves. Figure 5 makes this clear.
Figure 5: Interrupted Bloch wave Both panels feature a circular absorbing zone seen clearly, as the real part of a plain wave (left) and of a Bloch wave (right) enter from above. A shadow persists and slowly fills by diffraction on the left, plane wave case, but strong scattering by the triangular lattice present on the right quickly fills in the void. This scattering, which is obscured by periodic coherence in the unobstructed Bloch waves, is plainly revealed as the void fills. The point is that the strong unprotected scattering in the shadow region cannot be strongly inelastic, or else passive barriers in periodic lattices would cause heating downstream.
The linear term
$H^{(1)}=-i\sum_{k,q}q\,V(q)\,c^{\dagger}_{k+q}c_{k}\sum_{j}e^{-iqR_{j}}u_{j},$ (83)
yields, upon expanding $u_{j}$ in phonon operators, the familiar Fröhlich interaction in which each term carries a definite phonon pseudomomentum.
At quadratic order one finds
$H^{(2)}=-\frac{1}{2}\sum_{k,q}q^{2}V(q)\,c^{\dagger}_{k+q}c_{k}\sum_{j}e^{-iqR_{j}}u_{j}^{2},$ (84)
and writing
$u_{j}=\frac{1}{\sqrt{N}}\sum_{Q}U_{Q}\,e^{iQR_{j}}(a_{Q}+a^{\dagger}_{-Q}),$ (85)
one obtains terms of the form
$\begin{split}H^{(2)}\sim\sum_{k,q}\sum_{Q,Q^{\prime}}c^{\dagger}_{k+q}c_{k}\,(a_{Q}+a^{\dagger}_{-Q})(a_{Q^{\prime}}+a^{\dagger}_{-Q^{\prime}})\\ \times\sum_{j}e^{-i(q-Q-Q^{\prime})R_{j}}.\end{split}$ (86)
Among these are phonon-diagonal contributions with $Q^{\prime}=-Q$,
$H^{(2)}_{\rm el}\sim\sum_{k,q,Q}c^{\dagger}_{k+q}c_{k}\,\bigl(2a_{Q}^{\dagger}a_{Q}+1\bigr)\sum_{j}e^{-iqR_{j}},$ (87)
which have nonzero matrix elements
$\langle\nu|H^{(2)}_{\rm el}|\nu\rangle\neq 0.$ (88)
These describe elastic electron scattering with $\nu\to\nu$, and more generally quasi-elastic processes when $\nu\to\nu^{\prime}$ with small energy change.
Thus, once the interaction is treated beyond the linear Fröhlich approximation, elastic and quasi-elastic channels appear already at first order in the density-density interaction. In such processes $\Delta K_{\rm ph}=0$, while in general $\Delta K_{e}\neq Gm$, so the standard selection rule
$\Delta K_{e}+\Delta K_{\rm ph}=Gm$ (89)
is not satisfied within the zero-mode-removed theory.
This is consistent with the observation, emphasized by Kohn, that the electronic band structure itself arises from coherent elastic scattering from the periodic lattice potential. Elastic scattering is therefore not an exceptional process but the dominant one that defines the electronic states. The restriction to inelastic, one-phonon processes is a feature of the linearized phonon representation, not of the underlying dynamics.
The appearance of elastic and quasi-elastic channels thus exposes the limitations of the reduced description. Once the lattice center-of-mass degree of freedom is omitted, neither mechanical momentum nor pseudomomentum is properly accounted for within the phonon subspace alone. The full problem requires a dynamical background capable of carrying the compensating momentum.
VII.6 Thermal averaging and the Debye-Waller fraction
In thermal equilibrium the internal lattice degrees of freedom are described by the density matrix
$\hat{\varrho}=\frac{e^{-\beta\hat{H}_{\mathrm{lat}}}}{Z}=\sum_{n}p_{n}\,|n\rangle\langle n|,\qquad p_{n}=\frac{e^{-\beta E_{n}}}{Z},$ (90)
where $|n\rangle$ denote exact many-body eigenstates of the lattice Hamiltonian $\hat{H}_{\mathrm{lat}}$ with energies $E_{n}$. For elastic internal scattering the relevant quantity is the thermal expectation value of the internal density operator,
$\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}=\sum_{n}p_{n}\,\langle n|\rho^{\rm int}_{\mathbf{q}}|n\rangle.$ (91)
The coherent elastic scattering intensity is
$I_{\mathrm{el}}(\mathbf{q})=\big|\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}\big|^{2}.$ (92)
The total scattering intensity summed over all final internal states is
$I(\mathbf{q})=\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}.$ (93)
The elastic (coherent) branching fraction is therefore
$f_{\mathrm{el}}(\mathbf{q},T)=\frac{\big|\langle\rho^{\rm int}_{\mathbf{q}}\rangle_{T}\big|^{2}}{\left\langle\rho^{\rm int}_{\mathbf{q}}\rho^{\rm int}_{-\mathbf{q}}\right\rangle_{T}}.$ (94)
For a harmonic lattice one finds[21, 25]
$f_{\mathrm{el}}(\mathbf{q},T)\simeq e^{-2W(\mathbf{q},T)},\qquad W(\mathbf{q},T)=\frac{1}{2}\left\langle(\mathbf{q}\cdot\hat{\mathbf{u}})^{2}\right\rangle_{T},$ (95)
the familiar Debye-Waller result. Even at finite temperature a substantial fraction of momentum-transfer events leave the lattice in the same internal state.
VII.7 Estimate of the elastic internal fraction at representative momentum transfers
For momentum relaxation in a metal it is natural to consider momentum transfers of order the Fermi momentum, since large-angle scattering dominates the transport relaxation rate through the usual $(1-\cos\theta)$ weighting. For copper, a free-electron estimate yields
$k_{F}\simeq 1.36\ \text{\AA }^{-1},\qquad 2k_{F}\simeq 2.7\ \text{\AA }^{-1}.$ (96)
We therefore evaluate the elastic internal (phonon-diagonal) fraction $f_{\mathrm{el}}(\mathbf{q},T)$ at representative values $q\ll k_{F}$, $q\sim k_{F}$, and $q\sim 2k_{F}$.
Within the isotropic Debye-Waller estimate used above,
$2W(\mathbf{q},T)=\frac{q^{2}\langle u^{2}\rangle_{T}}{3},\qquad f_{\mathrm{el}}(\mathbf{q},T)\simeq e^{-2W(\mathbf{q},T)}.$ (97)
Using the room-temperature mean-square displacement for copper, $\langle u^{2}\rangle_{300\mathrm{K}}\simeq 1.9\times 10^{-2}\ \text{\AA }^{2}$, one finds the values summarized in Table 1.
qq-1) $2W$ $f_{\mathrm{el}}$
$0.3$ $5.7\times 10^{-4}$ $0.9994$
$k_{F}\simeq 1.36$ $1.2\times 10^{-2}$ $0.988$
$2k_{F}\simeq 2.7$ $4.6\times 10^{-2}$ $0.955$
Table 1: Estimated elastic internal fraction $f_{\mathrm{el}}$ for copper at room temperature ($T=300$ K) for representative momentum transfers. Even for large-angle scattering with $q\sim 2k_{F}$, the majority of momentum-relaxing events remain phonon-diagonal, with momentum absorbed predominantly by the lattice background rather than by internal phonon excitation.
These estimates show that, for momentum transfers relevant to transport, the overwhelming fraction of scattering events are internally elastic, in the sense that they leave the lattice in the same vibrational state. Phonon creation constitutes only a minor branching channel at room temperature, even for large-angle scattering. Momentum exchange with the lattice therefore occurs predominantly through elastic or quasi-elastic coupling to the background degrees of freedom, with internal phonon excitation representing a secondary correction.
We can motivate this result as follows: For a Cu atom modeled as an Einstein oscillator, a sudden momentum transfer $\Delta p=\hbar q$ leaves the oscillator in its ground state with probability
$P_{0\to 0}=\left|\langle 0|e^{iqx}|0\rangle\right|^{2}=\exp\!\left(-\frac{E_{R}}{\hbar\omega_{E}}\right),\qquad E_{R}=\frac{(\hbar q)^{2}}{2M}.$ (98)
Taking $q=2k_{F}$ for copper, with $k_{F}\approx 1.36\,\mathrm{\AA }^{-1}$ and $M\approx 63.5\,u$, gives $E_{R}\approx 0.24\,\mathrm{meV}$. For a typical local vibrational energy $\hbar\omega_{E}\sim 20$-$30\,\mathrm{meV}$, one finds $P_{0\to 0}\approx 0.99$.
VIII Experiments suggesting large elastic contributions
We have already touched on pure wire resistivity and persistent currents above, as allowing a more satisfactory explanation of the experiments through inclusion of the zero modes and the expansion to elastic and quasielastic channels that they enable.
Other experimental evidence suggesting predominantly elastic momentum relaxation in metals at low temperature has accumulated for decades, particularly in the mesoscopic-transport and quantum-interference literature. In this section we show that the framework developed here can resolve several longstanding disparities between theory and experiment regarding electronic transport. We review a range of experiments that demonstrate a clear separation between momentum relaxation and energy relaxation, consistent with elastic or quasi-elastic scattering dominating momentum randomization.
VIII.1 Hierarchy of momentum and energy relaxation times
Electrical resistivity and optical conductivity probe momentum relaxation, quantum-interference phenomena such as weak localization probe phase coherence, and nonequilibrium measurements including shot noise, Johnson-noise thermometry, and hot-electron relaxation directly access energy exchange. Strikingly, these probes consistently reveal a strong hierarchy of timescales,
$\tau_{\mathrm{p}}\ll\tau_{E},$ (99)
established using independent experimental techniques and across a wide range of materials. Here $\tau_{\mathrm{p}}$ is the timescale on which electronic momentum (or pseudomomentum) is randomized by scattering, while $\tau_{E}$ is the timescale over which electrons exchange energy with the background and relax toward thermal equilibrium.
The consistent appearance of this hierarchy demonstrates that momentum relaxation and energy relaxation are distinct physical processes. These experiments show that electronic momentum can be efficiently scrambled by a time-dependent environment without significant energy exchange. In other words, the lattice need not act as an energy sink in order to act as a momentum sink. This separation is precisely what is expected when electrons scatter quasi-elastically from a dynamically fluctuating lattice background.
VIII.2 Quantum interference and dephasing
The suppression of quantum-interference effects with increasing temperature is commonly described in terms of inelastic scattering processes that exchange energy with environmental degrees of freedom [23, 2]. It is useful, however, to distinguish between dephasing as operationally defined in transport experiments and irreversible decoherence in the strict quantum-mechanical sense.
A broad class of experiments indicates that substantial suppression of interference can occur even when direct energy relaxation to the lattice remains weak. Mesoscopic interference measurements, for example, show a pronounced reduction— and in some cases an apparent saturation—of the phase-coherence time $\tau_{\phi}$ at low temperatures, while independent probes suggest that electron–phonon energy exchange rates continue to decrease in the same regime [27, 17, 35]. This empirical separation of timescales illustrates that phase and momentum randomization need not scale directly with the rate of substantial energy transfer to internal lattice excitations.
Weak-localization is robust in experiments, implying that the backscattering is coherent and therefore elastic. [23, 27].Elastic scattering amplitudes, including diffuse non-Bragg amplitudes, are reduced by the well known Debye–Waller factor.
VIII.3 Hot-electron relaxation
A particularly direct separation of momentum and energy relaxation is provided by hot-electron experiments in metals [36, 30, 11, 12]. In these measurements, nonequilibrium electron distributions are injected, and the subsequent momentum randomization and energy loss are probed independently. Over broad temperature ranges, the momentum-relaxation time inferred from transport or magnetotransport is found to be substantially shorter than the energy-relaxation time associated with thermalization. Electrons therefore undergo many momentum-deflecting collisions while remaining close to isoenergetic on the same timescale.
This pronounced separation of timescales indicates that efficient momentum randomization does not require comparably rapid energy transfer to lattice excitations. While inelastic electron–phonon scattering ultimately governs thermalization, the observed hierarchy suggests that additional scattering channels, with weak net energy exchange on electronic scales, can play an important role in momentum relaxation within the metallic state.
Johnson-noise thermometry provides a direct probe of electron–lattice energy exchange. Experiments consistently show that the electron temperature relaxes on timescales much longer than those associated with momentum relaxation inferred from transport[11]. This separation, $\tau_{p}\ll\tau_{E}$, is difficult to reconcile with a picture in which momentum relaxation is intrinsically inelastic. It is, however, naturally explained if momentum is transferred predominantly through elastic or quasielastic processes, while energy relaxation proceeds through slower coupling to internal lattice degrees of freedom.
VIII.4 Shubnikov-de Haas oscillations
Quantum oscillation phenomena such as the Shubnikov-de Haas and de Haas-van Alphen effects provide a precise probe of electronic coherence and scattering in metals at low temperature [33, 3]. These oscillations arise from Landau quantization and require that electrons execute many cyclotron orbits without loss of phase coherence.
Within the conventional Lifshitz-Kosevich (LK) framework, the oscillation amplitude may be written as
$A(T,B)\propto\frac{X}{\sinh X}\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right),\qquad X=\frac{2\pi^{2}k_{B}T}{\hbar\omega_{c}},$ (100)
where the factor $X/\sinh X$ describes thermal smearing of the Fermi surface, while the Dingle factor
$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right)$ (101)
encodes Landau-level broadening through the quantum lifetime $\tau_{q}$. In standard treatments, $\tau_{q}$ is associated with inelastic or quasi-inelastic processes, and is often assumed to become weakly temperature dependent at low $T$.
From the present viewpoint, this interpretation is unnecessarily restrictive. When the lattice center-of-mass degree of freedom is retained, momentum relaxation need not proceed through phonon creation or annihilation. Instead, the electron can exchange momentum elastically with the lattice background via recoil of the zero mode, while total mechanical momentum is conserved. These processes are elastic in energy, but nonetheless randomize the electronic phase.
As a result, the Dingle factor should be understood more generally as encoding phase randomization rather than strictly inelastic scattering. Elastic recoil processes provide a continuous channel for Landau-level broadening, even in the absence of phonon excitation. The temperature dependence then arises not from phonon occupation factors, but from the growth of available recoil phase space in the thermally fluctuating background.
This leads naturally to a temperature-dependent quantum lifetime of the form
$\frac{1}{\tau_{q}(T)}\sim T^{\alpha},$ (102)
with $\alpha\sim 1$ in a Planckian regime, implying
$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}(T)}\right)\;\sim\;\exp\!\left(-\mathrm{const}\,\frac{T}{B}\right).$ (103)
Thus, the damping of quantum oscillations can remain strongly temperature dependent down to low $T$ without invoking inelastic scattering.
The persistence of Shubnikov-de Haas oscillations over wide temperature ranges therefore indicates that substantial momentum randomization can occur through predominantly elastic processes, while phase coherence is degraded only gradually. This resolves the apparent tension between strong scattering and long-lived cyclotron coherence, and places quantum oscillation damping on the same footing as weak localization: both probe phase randomization in a dynamically fluctuating, but not necessarily energy-relaxing, environment.
This interpretation is consistent with experimental observations of a marked temperature dependence of the Dingle factor, often extending to low $T$ with approximately power-law or linear-in-$T$ behavior, rather than saturation. In contrast to conventional theory, which attributes Landau-level broadening primarily to inelastic processes and therefore expects weak temperature dependence at low $T$, the present framework predicts a continued $T$-dependent damping arising from elastic recoil-driven phase randomization.
IX Discussion
Here we examine the consistency of the present framework with well-established experimental results, including the Wiedemann–Franz law, weak localization, and quantum oscillations.
Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott–Ioffe–Regel crossover. The latter discussion is exploratory and is presented as a coherent physical interpretation rather than as a definitive claim.
IX.1 Implications for the Lorenz ratio
Elastic scattering randomizes momentum as effectively as inelastic scattering within standard transport formalisms. Accordingly, a predominantly elastic microscopic channel does not conflict with the established success of conventional transport theory.
In a degenerate metal, the Wiedemann–Franz (WF) law states that the ratio of the electronic thermal conductivity $\kappa$ to the electrical conductivity $\sigma$ times temperature approaches the Sommerfeld value
$L\equiv\frac{\kappa}{\sigma T}\;\xrightarrow[T\to 0]{}\;L_{0}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}.$ (104)
Although some inelastic processes must ultimately relax the electronic energy current and establish global thermal equilibrium, it has long been understood that these processes need not determine the leading transport coefficients themselves [39, 22, 31].
If the dominant momentum-relaxing channel is elastic and only weakly energy dependent over the thermal window $\sim k_{B}T$ about the Fermi level, then both electrical and thermal currents are governed by the same relaxation time $\tau_{\mathrm{el}}$,
$\sigma\propto\tau_{\mathrm{el}},\qquad\kappa\propto\tau_{\mathrm{el}},$ (105)
and the Sommerfeld value $L_{0}$ follows in the usual way. Inelastic processes with rate $1/\tau_{\mathrm{in}}\ll 1/\tau_{\mathrm{el}}$ transfer energy between electrons and lattice degrees of freedom but do not enter the leading expressions for $\sigma$ or $\kappa$. A transport regime dominated by elastic momentum randomization is therefore fully compatible with the WF law, even when true energy relaxation is parametrically slower.
IX.2 Weak localization
Weak localization (WL) provides a clear illustration of the central role of elastic backscattering in low-temperature metallic transport. The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that revisit the same spatial region after multiple scattering events [7, 23]. It is often the case that the WL correction is important only with defects causing elastic backscattering. However the electrons still have to arrive back intact, so to speak (i.e. in phase), so other inelastic events cannot have decohered the electron on the way out and back. This is a tall order if any “natural” deflections are inelastic. elcrtherefore requires that a substantial fraction of backscattering processes preserve phase coherence over the relevant timescale.
If scattering were strongly inelastic on the transport timescale, interference between time-reversed paths would be suppressed before a WL correction could develop. Experimentally, however, a well-defined suppression of conductivity and characteristic negative magnetoresistance are observed at low temperature. As temperature increases, the WL correction is gradually reduced. Although this reduction is commonly described in terms of increasing “inelastic scattering,” what is operationally measured is the loss of phase coherence rather than a rapid increase in electronic energy relaxation.
Independent probes, including hot-electron relaxation and Johnson-noise thermometry, indicate that electron–phonon energy exchange can remain comparatively slow in the same temperature range. The suppression of WL is therefore consistent with the onset of a time-dependent scattering environment that randomizes phase while involving only modest energy transfer on electronic scales.
Within the present framework this behavior fits naturally into the hierarchy of timescales
$\tau_{p}\ll\tau_{\phi}\ll\tau_{E},$
where elastic momentum randomization produces diffusive transport, while slower temporal evolution of the lattice background reduces phase coherence without requiring substantial phonon excitation or rapid energy dissipation.
IX.3 Planckian diffusion and linear-in-$T$ resistivity
Much recent literature discusses Planckian behavior in terms of a microscopic relaxation time
$\tau_{\mathrm{Pl}}\sim\frac{\hbar}{k_{B}T},$
often interpreted as setting a characteristic scale for scattering or energy dissipation [37, 15, 14, 32]. Within this perspective, transport coefficients are commonly estimated by inserting $\tau_{\mathrm{Pl}}$ into kinetic or hydrodynamic expressions.
The framework developed in Ref. [5, 38] and extended here suggests a complementary interpretation. The primary microscopic result is the emergence of real-space diffusion with diffusion constant
$D\sim\frac{\hbar}{m^{\ast}},$ (106)
arising from elastic momentum scrambling in a time-dependent lattice background. In this formulation, diffusion emerges from quantum dynamics in a fluctuating potential and does not rely explicitly on strong inelastic scattering or rapid energy dissipation. Wave-on-wave simulations of the coupled electron–lattice problem demonstrate this mechanism explicitly [5, 19, 38].
Once diffusion is established, Einstein relations connect $D$ to transport coefficients,
$\sigma=\chi D,\qquad D=\mu\frac{k_{B}T}{e},$ (107)
so that expressing $D$ in terms of an effective relaxation time yields [5]
$\tau_{\mathrm{eff}}\sim\frac{\hbar}{k_{B}T},$ (108)
up to factors of order unity. From this viewpoint, the Planckian timescale may be understood not as a fundamental bound on inelastic scattering, but as an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics. It characterizes the rate at which charge spreads spatially rather than the rate at which energy must be irreversibly transferred to internal degrees of freedom, and thus fits nicely into the main elastic and quasielastic theme of this paper.
IX.4 Coulomb drag
We began discussing Coulomb drag in section II. It is an ideal case in which the zero recoil mode, missing in the literature on coulomb drag, is likely the agent of the power law seen in experiments. Theoretical analyses of one-dimensional transport consistently emphasize the strong kinematic constraints imposed by simultaneous energy and momentum conservation, leading to the conclusion that relaxation processes should be strongly suppressed at low temperature [26, 18] . Experimentally, however, relaxation and drag in quantum wires are often observed to exhibit power-law temperature dependence rather than strong suppression. This discrepancy is typically attributed to additional mechanisms such as disorder, band curvature, or coupling to external degrees of freedom that relax the kinematic constraints. This will be the topic of a future publication.
IX.5 Phonon drag
In conventional transport theory, the phonon-drag contribution to the thermopower is attributed to momentum transfer from a nonequilibrium phonon population to the electrons, and is therefore tied to phonon creation and annihilation. This reflects a formulation in which the lattice center-of-mass (zero mode) is removed and momentum exchange is forced into the phonon sector.
When the zero mode is retained, momentum transfer is no longer tied to phonon number. The electron couples to the lattice deformation field, and momentum can be exchanged with the lattice as a whole through recoil, without requiring phonon excitation. The conventional phonon-drag contribution therefore represents only a subset of a more general lattice-drag response.
Accordingly, the thermopower should be written as
$S=S_{\rm diff}+S_{\rm lat},\qquad S_{\rm lat}=S_{\rm el}+S_{\rm inel},$ (109)
where $S_{\rm inel}$ is the usual phonon-drag term and $S_{\rm el}$ arises from elastic or quasielastic momentum transfer via lattice recoil. In regimes where scattering is predominantly elastic, $S_{\rm el}$ need not be small and may dominate the drag response.
X Summary
We have revisited electron-lattice scattering starting from the standard microscopic Hamiltonian, but retaining a degree of freedom that is nearly always removed: the translational center-of-mass (zero mode) of the lattice. In conventional formulations, this mode is eliminated (e.g., through Born-von Kármán boundary conditions), leaving momentum transfer to be carried entirely by internal phonon excitations. Together with the form of the first order Fröhlich expansion, changes in electronic momentum are tied to phonon creation or annihilation, and scattering is treated as intrinsically inelastic.
When the lattice center of mass is retained explicitly, and the density-density form of the interaction is retained and not expanded, this picture changes qualitatively. Total mechanical momentum is conserved by recoil of the lattice as a whole, while pseudomomentum is redistributed between electron and background. A broad class of elastic and quasielastic momentum-transfer processes then becomes available, in which electronic momentum changes are balanced by the zero mode without requiring phonon excitation.
This enlargement of the allowed phase space follows directly from the exact symmetries of the Hamiltonian. No new interaction has been introduced. Retaining the full density-density interaction reveals that elastic matrix elements arise already at the level of the state-to-state transition amplitude $\langle f|H_{\rm int}|i\rangle$, without invoking higher-order or multiphonon processes. The familiar phonon-mediated processes remain, but represent a restricted subset of a much larger set of allowed transitions.
The resulting kinematics has broad implications. In one-dimensional systems, the need for thermally activated three-particle processes is removed, and power-law relaxation follows naturally. More generally, momentum relaxation need not be tied to energy dissipation: momentum can randomize rapidly while energy relaxes on much longer timescales. This resolves a long-standing tension between theory and experiment in clean metals, including observations from weak localization, quantum oscillations, hot-electron transport, and Coulomb drag.
We emphasize that these conclusions are not model-dependent. Once the full momentum conservation law of the isolated electron-lattice system is enforced, the existence of elastic and quasielastic scattering channels is unavoidable. The restriction to phonon-mediated inelastic scattering is therefore not a fundamental feature of the physics, but a consequence of having discarded the lattice center-of-mass degree of freedom.
XI Acknowledgments
E. Heller thanks B. Halperin for many discussions that provided a thorough understanding of the traditional theory. He also thanks Prof. Martina Hentschel for insightful comments that have improved the manuscript. This work was supported by the U.S. Department of Energy, Office of Science, under Grant No. DE-SC0025489.
Figure 6: Isolated systems of mass $M+m$ containing a foreground particle of mass $m$ and a background of mass $M$. The particle is attracted to the background and moves frictionlessly, except in C and D it may excite vibrations of the chain. In A, the particle is free except for walls. In B, the particle is in a fixed periodic potential within its host, the box. In C we replace the fixed corrugation with a chain of masses connected by springs. In D, we dispense with the box, coming close to the situation that matters here. Of course the box could be sitting on a table, which could be fixed to a floating laboratory, which could be upon something so massive that it has significant gravitational attraction, yet that too is floating in space…
Appendix A Bloch waves form from elastic scattering of electrons in a lattice
By way or reinforcing the conclusion that elastic internal electron scattering must be very common in pure metals, we raise a point about the rather violent elastic scattering that is taking place, cl oacked by the symmetry of Bloch waves. The formation of Bloch waves in a periodic crystal may be viewed as the result of coherent multiple elastic scattering from the atoms of the lattice[20]. Band formation itself is therefore an elastic scattering phenomenon. One might then suppose that Bloch waves are somehow protected from further elastic deflection by phase coherence. However, this inference is incorrect, as illustrated in Fig. 5.
We extend this insight slighty and suppose some atom were temporarily slightly out of place due to thermal acoustic waves. The atoms would individually scatter as before, but now the phase locking is imperfect, and the electron can deflect. There is no new “phonon needed” physics that sets in due to truly minuscule thermal displacements. If the electron thus deflects, momentum is exchanged with the lattice without phonon excitation or energy dissipation.
Appendix B Tracking Momentum, and the Effects of Zero Mode Removal
Figure 6 helps illuminate the choices that can be made in treating electron-phonon interaction. It helps to underline that there is a great difference between treating the background as a dynamical body (with a translational zero mode) or as frozen, without recoil. Frozen is an impossibility in the real world, total mometum is always conserved, although this a self-consistent model. The question is, does that model lead to good real-world predictions, or does it mislead? Retaining the zero mode total momentum is not a matter of interpretation; it fixes the form of the wavefunction and alters the physical predictions in important ways.
Translation invariance of the whole requires that the Hamiltonian depends only on the center of mass and coordinate differences, so eigenstates separate into a plane wave for the center of mass coordinate and an internal state that depends only on relative coordinates.
B.1 Uplifting the particle in a box to the real world.
We therefore free the textbook particle in a box from its traditional and physically impossible bondage in case A.: considering a“free floating box” of mass $M$ containing a particle of mass $m$, with particle coordinate $r$ and box center-of-mass coordinate $R$, introduce
$\mathcal{R}=\frac{mr+MR}{m+M},\qquad$
Then a complete set of eigenstates (delta-normalized in $P$) may be written as
$\Psi_{P,n}(r,R)=\frac{1}{\sqrt{2\pi\hbar}}\,e^{iP\mathcal{R}/\hbar}\,\phi_{n}(r-R).$ (110)
Equation (110) is the operative content of keeping the zero mode: the internal structure is literally a function of rRr-R. The conserved total mechanical momentum is
$\hat{P}_{\rm tot}=-i\hbar\frac{\partial}{\partial\mathcal{R}},\qquad[H,\hat{P}_{\rm tot}]=0.$
Any internal state may be expanded in Fourier components on |rR|<L/2|r-R|<L/2,
$\phi_{n}(r-R)=\sum_{q}a_{nq}\,e^{iq(r-R)},$
and each component carries the shared phase
$e^{iq(r-R)}=e^{iqr}e^{-iqR}.$ (111)
Thus, in the plane-wave decomposition of the relative coordinate, momentum exchange is enforced at the level of the wavefunction: each Fourier component correlates $\hbar q$ on the particle with $\hbar q$ on the background. Phonons and other internal excitations may redistribute momentum within the background.
B.2 Box is corrugated
Suppose the box is as above but now contains a rigid, corrugated periodic potential (Fig. 6, B). Before the corrugation, the particle and the box each carry mechanical momentum, and only the total $P_{\rm tot}=p+P$ is conserved. In collisions, momentum is exchanged between the two in equal and opposite amounts, exactly as in free space.
The introduction of a periodic potential does not eliminate this structure, but reorganizes it. Continuous translational symmetry is reduced to discrete translational symmetry, and the corresponding conserved quantities become pseudomomenta defined modulo reciprocal lattice vectors. Importantly, the two-body structure of momentum exchange is preserved.
The particle now carries a pseudomomentum $\hbar k$, while the background box carries a corresponding pseudomomentum. These two pseudomomenta are exchanged in equal and opposite amounts during scattering, just as the mechanical momenta were exchanged before the corrugation was introduced. Their sum remains fixed and corresponds to the conserved total momentum of the isolated system.
Thus the corrugation does not replace two mechanical momenta with a single pseudomomentum; rather, it produces two pseudomomenta, one for the particle and one for the background. The center-of-mass degree of freedom of the box continues to carry the total mechanical momentum, while the periodic structure introduces a new, discrete labeling of the same underlying exchange process.
The situation depicted in Fig. 6, B therefore retains the essential physics of momentum exchange: the lattice remains a dynamical participant, even though its role is partially obscured by the use of pseudomomentum.
B.3 Elastic Chain in a Box, Free CM (phonons plus the zero mode)
Now take the background to be a dynamical chain (pink atoms) of total mass $M_{\rm lat}$ inside the box, with atomic coordinates $R_{j}$ and a center-of-mass coordinate $R_{\rm lat}$ (Fig. 6, C). A minimal model is
$H=\frac{p_{r}^{2}}{2m}+\sum_{j}\frac{p_{j}^{2}}{2M}+\frac{K}{2}\sum_{j}(R_{j+1}-R_{j}-a)^{2}+H_{\rm int}.$
With density-density coupling
$H_{\rm int}=\int dr\,dr^{\prime}\,\rho_{p}(r)V(r-r^{\prime})\rho_{\rm lat}(r^{\prime}),$
the Hamiltonian depends only on coordinate differences and is invariant under rigid translation
$r\rightarrow r+a_{0},\qquad R_{j}\rightarrow R_{j}+a_{0}.$
Consequently the conserved quantity is the true total mechanical momentum
$P_{\rm tot}=P_{e}+P_{\rm lat},\qquad[H,P_{\rm tot}]=0.$
Momentum exchange is enforced directly through the relative coordinate. Fourier transforming the lattice density yields
$\rho_{\rm lat}(q)=e^{-iqR_{\rm lat}}\rho^{\rm int}_{q},\qquad H_{\rm int}\propto e^{iq(r-R_{\rm lat})}.$
The phase factor $e^{iq(r-R_{\rm lat})}$ encodes the fundamental constraint: momentum $\hbar q$ transferred to the electron is accompanied by $\hbar q$ carried by the lattice center of mass. Internal degrees of freedom describe only relative motion and carry no net mechanical momentum.
The translational zero mode $R_{\rm lat}$ is therefore the unique carrier of the compensating momentum. Internal modes (phonons) may change the energy of the system, but they are not required to enforce momentum conservation.
Deleting the zero mode is not an innocuous simplification. It removes the only degree of freedom capable of carrying recoil and eliminates the shared-coordinate structure $\phi(r-R)$ from the Hilbert space. The resulting theory no longer represents true mechanical momentum conservation; instead it replaces it with pseudomomentum bookkeeping within a reduced space. Momentum exchange is then no longer enforced through the relative coordinate rRr-R, but appears only as a selection rule requiring phonon pseudomomentum to balance electron pseudomomentum (up to Umklapp).
Thus the apparent kinematic necessity of phonon production does not arise from fundamental principles. It is imposed by amputating the translational degree of freedom that would otherwise carry the compensating momentum.
B.4 Free Chain (no box), Free CM (translation invariance made manifest)
Removing the box entirely leaves a floating chain. This is simply Fig. 6, C with boundary constraints eliminated. The lesson is unchanged but becomes explicit: the conserved momentum is the generator of rigid translation of all coordinates.
Any description that retains translational invariance necessarily contains the shared phase $e^{iq(r-R_{\rm lat})}$, since it is the Fourier representation of the relative coordinate. This phase structure is the direct expression of momentum conservation in the full system.
B.5 Removing the Zero Mode
The conventional electron-phonon kinematics is obtained by a construction that is the Hilbert-space analog of “nailing the box down,” but without introducing an external dynamical reservoir. The translational zero mode is removed at the outset. One expands the lattice displacement field in normal modes and omits the $q=0$ component. The resulting description retains only internal degrees of freedom.
With the zero mode removed, continuous translation symmetry is no longer present, and the conserved quantity is not the true mechanical momentum. What remains is discrete lattice translation symmetry, with conserved crystal pseudomomentum
$K_{\rm tot}=K_{e}+K_{\rm ph}\quad(\mathrm{mod}\;G),\qquad K_{\rm ph}=\sum_{q}\hbar q\,a_{q}^{\dagger}a_{q},$
and the corresponding selection rule
$\Delta K_{e}+\Delta K_{\rm ph}=Gm.$
In this representation, momentum transfer is confined to phonons: changes in the electron pseudomomentum must be balanced entirely by phonon pseudomomentum, up to Umklapp. The lattice center-of-mass degree of freedom, which would otherwise provide a direct channel for momentum exchange through rigid recoil, is absent.
The apparent kinematic necessity of phonon emission or absorption is therefore not a physical result. It is a consequence of having eliminated the translational degree of freedom that would otherwise carry the compensating momentum. It is not a choice one normally would wish to make, because it changes the predicted physical processes. It is surprising perhaps, but it does. The “thermodynamic limit” does not protect us here, and if it seems that it does, the limit $N\to\infty$ has not been taken properly. Jumping to $N=\infty$ is not taking the limit of course.
Appendix C Whither Umklapp?
The thermodynamic limit and Umklapp processes are often invoked to account for momentum relaxation in perfect crystals. These mechanisms are valid within the conventional formulation, but their apparent necessity reflects the prior removal of the lattice center-of-mass (zero-mode) degree of freedom.
In the standard phonon description, momentum conservation is imposed only at the level of crystal pseudomomentum,
$K_{\rm tot}=K_{e}+K_{\rm ph}\quad(\mathrm{mod}\;G),$
and electronic momentum can relax only through phonon processes, supplemented by Umklapp when reciprocal lattice vectors are invoked. This framework is kinematically restrictive: the phonon system tracks momentum only modulo $G$, and the available channels are limited to discrete reciprocal lattice transfers.
When the zero mode is retained, this restriction is lifted. The lattice background can recoil and absorb arbitrary mechanical momentum, and the scattering problem is no longer confined to Umklapp processes. Instead, momentum transfer occurs within a continuous manifold of elastic and quasielastic processes involving the full dynamical background.
To make this explicit, we separate a single electron (the foreground) from the rest of the system (the background), which includes all ions and all other electrons. The total mechanical momentum is
$\hat{P}_{\rm tot}=\hat{p}+\hat{P}_{\rm bg},\qquad\hat{P}_{\rm bg}=\sum_{\alpha\in\mathrm{bg}}\hat{P}_{\alpha}.$ (112)
The operator $\hat{P}_{\rm bg}$ generates rigid translation of the background and constitutes its center-of-mass (zero-mode) momentum. Internal phonon modes describe only relative motion and do not furnish an independent contribution to $\hat{P}_{\rm bg}$. Thus any change in the electron’s mechanical momentum must be balanced by the background zero mode.
This makes clear what is implicit in Umklapp. When a reciprocal lattice vector $G$ appears in a scattering process, the associated momentum $\hbar G$ cannot be attributed to the electron alone, nor is it carried by internal phonon modes. It must be carried by the lattice as a whole. In the conventional formulation, however, the zero mode that would carry this momentum has been removed, and the role of recoil is replaced by bookkeeping modulo $G$.
The traditional role assigned to Umklapp in producing finite resistivity therefore arises within a restricted description in which the background is effectively clamped. Once the background is treated dynamically, electronic momentum need not decay through Umklapp. Instead, it can be transferred elastically to the lattice as a whole.
Umklapp remains a useful classification of pseudomomentum conservation modulo reciprocal lattice vectors. It is not, however, the unique or even generic mechanism for momentum relaxation in a real crystal, where the background is free to recoil.
Appendix D The landscape seen by an electron
At temperatures of order $T\sim 100\,$K, thermally excited acoustic distortions in a typical metal generate substantial deformation-potential landscapes on nanometer length scales. Using standard values for the deformation potential ($g\sim 5$$15\,\mathrm{eV}$), sound velocity ($v_{s}\sim 3$$5\times 10^{3}\,\mathrm{m/s}$), and the characteristic thermal phonon wavevector $q_{T}\sim k_{B}T/\hbar v_{s}$, one finds order-of-magnitude estimates for the resulting deformation-potential field gradients of
$|\nabla V_{\rm def}|\sim 10^{4}\text{–}10^{6}\ \mathrm{V/cm}\qquad(T\sim 100\,\mathrm{K}).$ (113)
Over a representative lateral scale of $L\sim 10\,$nm, this corresponds to peak-to-peak variations of the screened scalar potential of order
$\Delta V_{\rm pp}\sim|\nabla V_{\rm def}|\,L\sim 10\text{–}300\ \mathrm{mV}.$ (114)
[…]